### v4

# Spectrum-Aware Debiasing: A Modern Inference Framework with Applications to Principal Components Regression

Abstract
Debiasing is a fundamental concept in high-dimensional statistics. While degrees-of-freedom adjustment is the state-of-the-art debiasing technique in high-dimensional linear regression, it largely remains limited to independent, identically distributed samples and sub-Gaussian covariates. These limitations hinder its wider practical use. In this paper, we break this barrier and introduce Spectrum-Aware Debiasing—-a novel inference method that applies to challenging high-dimensional regression problems with structured row-column dependencies, heavy tails, asymmetric properties, and latent low-rank structures. Our method achieves debiasing through a rescaled gradient descent step, where the rescaling factor is derived from the spectral properties of the sample covariance matrix. This spectrum-based approach enables accurate debiasing in much broader contexts. We study the common modern regime where the number of features and samples scale proportionally. We establish asymptotic normality of our proposed estimator (suitably centered and scaled) under various convergence notions when the covariates are right-rotationally invariant. Such designs have garnered significant recent attention due to their crucial role in compressed sensing. Furthermore, we devise a consistent estimator for its asymptotic variance.
Our work has two notable by-products: first, Spectrum-Aware Debiasing rectifies the bias in principal components regression (PCR), providing the first debiased PCR estimator in high dimensions. Second, we introduce a principled test for checking the presence of alignment between the signal and the eigenvectors of the sample covariance matrix. This test is independently valuable for statistical methods developed using approximate message passing, leave-one-out, or convex Gaussian min-max theorems. We demonstrate the utility of our method through diverse simulated and real data experiments. On the technical front, we connect approximate message passing algorithms with debiasing and provide the first proof of the Cauchy property of vector approximate message passing (V-AMP).
62E20,
62F12,
High-dimensional inference,
Debiasing,
Spectral Properties,
Principal Components Regression,
Debiased PCR,
Right-rotationally invariant designs,
Vector approximate message passing,
keywords:
[class=MSC]
keywords:
1 Introduction
Regularized estimators constitute a basic staple of high-dimensional regression. These estimators incur a regularization bias, and characterizing this bias is imperative for accurate uncertainty quantification. This motivated debiased versions of these estimators [105, 55, 93] that remain unbiased asymptotically around the signal of interest. To describe debiasing, consider the setting of a canonical linear model where one observes a sample of size $n$ satisfying
$\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\bm{\varepsilon}.$
Here $\mathbf{y}\in\mathbb{R}^{n}$ denotes the vector of outcomes, $\mathbf{X}\in\mathbb{R}^{n\times p}$ the design matrix, $\bm{\beta}^{\star}\in\mathbb{R}^{p}$ the unknown coefficient vector, and $\bm{\varepsilon}$ the unknown noise vector. Suppose $\bm{\hat{\bm{\beta}}}$ denotes the estimator obtained by minimizing $\mathcal{L}(\bm{\cdot}\;;\mathbf{X},\mathbf{y}):\mathbb{R}^{p}\mapsto\mathbb{R}_{+}$ given by
$\mathcal{L}(\bm{\beta};\mathbf{X},\mathbf{y}):=\frac{1}{2}\|\mathbf{y}-\mathbf{X}\bm{\beta}\|^{2}+\sum_{i=1}^{p}h\left(\beta_{i}\right),\qquad\bm{\beta}\in\mathbb{R}^{p},$ (1)
where $h:\mathbb{R}\mapsto[0,+\infty)$ is some convex penalty function. Commonly used penalties include the ridge $h(b)=\lambda b^{2},\lambda>0$, the Lasso $h(b)=\lambda|b|,\lambda>0$, the Elastic Net $h(b)=\lambda_{1}|b|+\lambda_{2}b^{2},\lambda_{1},\lambda_{2}>0$, etc. The debiased version of $\bm{\hat{\bm{\beta}}}$ takes the form
$\hat{\bm{\beta}}^{u}=\hat{\bm{\beta}}+\frac{1}{\widehat{\mathsf{adj}}}\bm{M}\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}}),$ (2)
[…]
Classical statistics textbooks tell us that when the dimension $p$ is fixed and the sample size $n$ approaches infinity, the debiased estimator $\hat{\bm{\beta}}^{u}$ reduces to the well-known one-step estimator. In this case, Gaussianity of $\hat{\bm{\beta}}^{u}-\bm{\beta}^{\star}$ follows from [94, Theorem 5.45] by choosing $\mathbf{M}=(\mathbf{X}^{\top}\mathbf{X})^{-1}$, the inverse of the sample covariance matrix, and $\widehat{\mathsf{adj}}=1$, requiring no adjustment. Early work on ultra high-dimensional problems ($p\gg n$) [93, 105, 55, 16] established that when the signal $\bm{\beta}^{\star}$ is sufficiently sparse, the Lasso can be debiased by taking $\mathbf{M}$ as suitable “high-dimensional” substitutes of $(\mathbf{X}^{\top}\mathbf{X})^{-1}$ and setting $\widehat{\mathsf{adj}}=1$. However, later work uncovered that an adjustment of $\widehat{\mathsf{adj}}<1$ is necessary to relax sparsity assumptions on $\bm{\beta}^{\star}$ or to debias general regularized estimators beyond the Lasso. For instance, [53, 11] established that when the signal is not sufficiently sparse, the adjustment for the Lasso should be $\widehat{\mathsf{adj}}=1-\hat{s}/n$ with $\mathbf{M}=\bm{\Sigma}^{-1}$, where $\hat{s}$ denotes the number of non-zero entries in $\hat{\bm{\beta}}$. This correction term was named the “degrees-of-freedom adjustment” since $\hat{s}$ corresponds to the degrees-of-freedom of the estimator $\hat{\bm{\beta}}$ [110].
Figure 1: Histograms of empirical distribution of $(\hat{\tau}_{*}^{-1/2}(\hat{{\beta}}^{u}_{i}-{\beta}^{\star}_{i}))_{i=1}^{p}$ comparing Degrees-of-Freedom Debiasing [12] with our Spectrum-Aware Debiasing, where $\hat{\bm{\beta}}^{u}$ is the debiased Elastic-Net estimator with tuning parameters $\lambda_{1}=1,\lambda_{2}=0.1$. The first row uses the Degrees-of-Freedom Debiasing from [12] with $\mathbf{M}=\mathbf{I}_{p}$ (denoted DF). The second row uses our Spectrum-Aware Debiasing (denoted SA) as in Definition 3.4. Entries of the signal $\bm{\beta}^{\star}$ are i.i.d. draws from $0.24\cdot N(-20,1)+0.06\cdot N(10,1)+0.7\cdot\delta_{0}$ where $\delta_{0}$ is Dirac-delta function at 00. Thereafter, the signal is fixed and we generate the responses using $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\bm{\varepsilon}$ where $\epsilon_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N(0,1)$. The solid black curve indicates a normal density fitted to the blue histograms whereas the dotted black line indicates the empirical mean corresponding to the histogram. See the corresponding QQ plot in Figure 5 from Appendix. All designs are re-scaled so that the average of eigenvalues of $\mathbf{X}^{\top}\mathbf{X}$ is 1. The design matrices are of shape $n=500,p=1000$.
Degrees-of-Freedom Debiasing introduced a novel perspective. However, it relied on some strict assumptions, namely independent and identically distributed (i.i.d.) data with Gaussian covariates that follow $\mathbf{X}_{i}\sim\mathcal{N}(\bm{0},\bm{\Sigma})$. Furthermore, Degrees-of-Freedom Debiasing used $\mathbf{M}=\bm{\Sigma}^{-1}$; thus, even when the i.i.d. assumption holds, implementing this estimator required knowledge of the true covariance matrix $\bm{\Sigma}$ or an accurate estimate. In summary, Degrees-of-Freedom Debiasing suffers three key limitations: (i) it is restricted to Gaussian type distributions (see [43] for extension to sub-Gaussians), thus failing to capture heavy-tailed or asymmetrically distributed covariates; (ii) it is ineffective in scenarios with heterogeneity or dependency among samples, and (iii) it faces challenges in choosing $\mathbf{M}$ when precise estimates of $\bm{\Sigma}$ are unavailable. These limitations restrict the broader applicability of Degrees-of-Freedom Debiasing to real-world settings that violate these strict assumptions.
We exemplify this issue in Figure 1, where we consider the following design distributions: (i) $\mathsf{MatrixNormal}$: $\mathbf{X}$ drawn from a matrix normal distribution with row and column correlations, i.e. $\mathbf{X}\sim N(\rm{0},\mathbf{\Sigma}^{\mathrm{(col)}}\otimes\mathbf{\Sigma}^{\mathrm{(row)}}),$ where $\mathbf{\Sigma}^{\mathrm{(col)}}_{ij}=0.5^{|i-j|}$ and $\bm{\Sigma}^{\mathrm{(row)}}$ follows an inverse-Wishart distribution with identity scale and degrees-of-freedom $1.1p$, which is chosen to be close to $p$ to encourage heavy-tails in the covariates; (ii) $\mathsf{Spiked}$: $\mathbf{X}$ contains latent structure, i.e. $\mathbf{X}=\alpha\cdot\mathbf{V}\mathbf{W}^{\top}+n^{-1}N(\rm{0},\mathbf{I}_{n}\otimes\mathbf{I}_{p})$ where $\alpha=10$ and $\mathbf{V}\in\mathbb{R}^{n\times m},\mathbf{W}\in\mathbb{R}^{p\times m}$ are drawn randomly from Haar matrices of dimensions $n,p$, and then we retain $m=50$ columns; (iii) $\mathsf{LNN}$: $\mathbf{X}$ formed by product of multiple random matrices (see [44, 45] for connections to linear neural networks), i.e. $\mathbf{X}=\mathbf{X}_{1}\cdot\mathbf{X}_{2}\cdot\mathbf{X}_{3}\cdot\mathbf{X}_{4}$ where $\mathbf{X}_{i}$’s have i.i.d. entries from $N(\rm{0},1)$; (iv) $\mathsf{VAR}$: rows of $\mathbf{X}$ drawn from a vector time series with the $i$-th row given by $\mathbf{X}_{i,\bullet}=\sum_{k=1}^{\tau\vee i}\alpha_{k}\mathbf{X}_{i-k,\bullet}+\bm{\varepsilon}_{i}$ where $\tau=3,\alpha=\quantity(0.4,0.08,0.04)$ and $\bm{\varepsilon}_{i}\sim N(\mathbf{0},\mathbf{\Sigma})$ with $\mathbf{\Sigma}$ drawn from an inverse-Wishart distribution with the same parameters as in (i); (v) $\mathsf{Mult}$-$\mathsf{t}$: rows of $\mathbf{X}$ drawn independently from a multivariate t-distribution with identity scale and degrees-of-freedom $3$. Figure 1 plot histograms of the empirical distribution of $\hat{\bm{\beta}}^{u}-\bm{\beta}^{\star}$ scaled by an estimate $\hat{\tau}_{*}$ of its standard deviation. The topmost panel uses the Degrees-of-Freedom Debiasing formula for $\hat{\bm{\beta}}^{u}$. We observe that the histograms in this panel deviate substantially from the overlaid standard Gaussian density. Degrees-of-Freedom Debiasing thus fails in these challenging settings. To underscore the difficulties posed by these examples, note that cases (i)-(iv) involve non-i.i.d. designs and (i),(iv),(v) involve heavy-tailed covariates.
In this paper, we propose a new debiasing formula that addresses the shortcomings of previous techniques and enables accurate debiasing in the aforementioned settings. To develop our method, we leverage the insight that a debiasing procedure effective for a wide range of scenarios must thoughtfully utilize the spectral characteristics of the data. To accomplish this, we explore an alternative path for modeling the randomness in the design. Instead of assuming that the rows of the design are i.i.d. Gaussian vectors, we require that the singular value decomposition of $\mathbf{X}$ satisfies certain natural structure that allows dependence among samples and potentially heavy-tailed distributions. Specifically, we assume that $\mathbf{X}$ is right-rotationally invariant (formally defined in Definition 2.1).
This class of designs has been widely studied in signal processing, information theory, and statistical physics [85, 64, 74, 86, 26, 31, 40, 87, 88, 95, 41, 62, 59, 97]. They serve as important prototypes for understanding fundamental high-dimensional phenomena in compressed sensing. Furthermore, recent advancements indicate that a wide variety of covariate distributions fall within the same universality class as these designs, provided that the eigenvectors of the sample covariance matrix are adequately “generic” [23, 27, 96]. Thus, the range of distributions covered by these designs is much larger than those considered in prior debiasing literature. In particular, this includes designs (i)-(v) from Figure 1.
We discover that for right-rotationally invariant designs, the accurate debiasing formula is given by
$\hat{\bm{\beta}}^{u}=\hat{\bm{\beta}}+\widehat{\mathsf{adj}}^{-1}\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}}),$ (3)
where $\widehat{\mathsf{adj}}$ solves the equation
$\frac{1}{p}{{\sum}}_{i=1}^{p}\frac{1}{\left(d_{i}^{2}-\widehat{\mathsf{adj}}\right)\left(\frac{1}{p}\sum_{j=1}^{p}\quantity(\widehat{\mathsf{adj}}+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right))^{-1}\right)+1}=1.$ (4)
Here, $\{d_{i}^{2}\}_{1\leq i\leq p}$ represents the eigenvalues of the sample covariance matrix $\mathbf{X}^{\top}\mathbf{X}$, and $h^{\prime\prime}$ denotes the second derivative of the penalty function $h$ used in calculating the regularized estimator $\hat{\bm{\beta}}^{u}$. At points of non-differentiability, we extend $h^{\prime\prime}$ by +$\infty$+ ∞ (cf. Section 2.2). The solution $\widehat{\mathsf{adj}}$ of (4) is unique for any $p\geq 1$ under mild assumptions (cf. Section 3.1). We refer to $\widehat{\mathsf{adj}}$ as “Spectrum-Aware adjustment” and the debiasing approach in (3) as “Spectrum-Aware Debiasing” since $\widehat{\mathsf{adj}}$ depends on the eigenvalues of $\mathbf{X}^{\top}\mathbf{X}$. Figure 1 illustrates the efficacy of our method. The second panel shows the empirical distribution of $\hat{\bm{\beta}}^{u}-\bm{\beta}^{\star}$, scaled by an appropriate estimate of its standard deviation, when $\hat{\bm{\beta}}^{u}$ is given by our Spectrum-Aware formula (3). Note the remarkable agreement with the overlaid standard Gaussian density. We should mention that, unlike Degrees-of-Freedom Debiasing, we do not need an estimate of $\bm{\Sigma}$, yet we can handle various correlation structures among the features, as demonstrated in Figure 1.
Despite the strengths of Spectrum-Aware Debiasing, we observe that it falls short when $\mathbf{X}$ contains outlier eigenvalues and/or the signal aligns with some eigenvectors of $\mathbf{X}$. To address these issues, we introduce an enhanced procedure that integrates classical Principal Components Regression (PCR) ideas with Spectrum-Aware Debiasing. In this approach, we employ PCR to handle the outlier eigenvalues while using a combination of PCR and Spectrum-Aware Debiasing to estimate the parts of the signal that do not align with an eigenvector. We observe that this hybrid PCR-Spectrum-Aware approach works exceptionally well in challenging settings where these issues are present.
We next summarize our main contributions below.
  • (i)
    We establish that our proposed debiasing formula is well-defined, that is, (4) admits a unique solution (Proposition 3.1). Then we establish that $\hat{\bm{\beta}}^{u}-\bm{\beta}^{\star}$, with this choice of $\widehat{\mathsf{adj}}$, converges to a mean-zero Gaussian with some variance $\tau_{\star}$ in a Wasserstein-2 sense (Theorem 3.1; Wasserstein-2 convergence notion introduced in Definition 2.2). Under an exchangeability assumption on $\bm{\beta}^{\star}$, we strengthen this result to convergence guarantees on finite-dimensional marginals of $\hat{\bm{\beta}}^{u}-\bm{\beta}^{\star}$ (Corollary 3.8).
  • (ii)
    We develop a consistent estimator for $\tau_{*}$ (Theorem 3.1) by developing new algorithmic insights and new proof techniques that can be of independent interest in the context of vector approximate message passing algorithms [74, 77, 37] (details in Section 5.2).
  • (iii)
    To establish the aforementioned points, we imposed two strong assumptions: (a) the signal $\bm{\beta}^{\star}$ is independent of $\mathbf{X}$ and cannot align with any subspace spanned by a small number of eigenvectors of $\mathbf{X}^{\top}\mathbf{X}$; (b) $\mathbf{X}^{\top}\mathbf{X}$ does not contain outlier eigenvalues. To mitigate these, we develop a PCR-Spectrum-Aware Debiasing approach (Section 4) that applies when these assumptions are violated. We prove asymptotic normality for this approach in Theorem 4.1.
  • (iv)
    We demonstrate the utility of our debiasing formula in the context of hypothesis testing and confidence interval construction with explicit guarantees on quantities such as the false positive rate, false coverage proportion, etc. (Sections 3.4 and 4.5).
  • (v)
    As a by-product, our PCR-Spectrum-Aware approach introduces the first methodology for debiasing the classical PCR estimator (Theorem 4.1), which would otherwise exhibit a shrinkage bias due to omission of low-variance principal components. We view this as a contribution in and of itself to the PCR literature since inference followed by PCR is under-explored despite the widespread usage of PCR.
  • (vi)
    As a further byproduct, we introduce a hypothesis test to identify alignment between principal components of the design matrix and the unknown regression coefficient $\bm{\beta}^{\star}$. This may be of independent interest in the context of statistical methods developed based on approximate message passing/leave-one-out/convex Gaussian min-max theorems.
  • (vii)
    Theoretically, we establish rigorously the risk of regularized estimators under right-rotationally invariant designs, a class much larger than Gaussian designs (Theorem 5.1). One should compare this theorem to [7] that developed a risk characterization of the Lasso in high dimensions under Gaussian design matrices.
  • (viii)
    Finally, we demonstrate the applicability of our Spectrum-Aware approach across a wide variety of covariate distributions, ranging from settings with heightened levels of correlation or heterogeneity among the rows or a combination thereof (Figure 3, top-left experiment), to diverse real data designs (Figure 3, bottom-left experiment). We observe that PCR-Spectrum-Aware Debiasing demonstrates superior performance across the board.
In the remaining Introduction, we walk the readers through some important discussion points, before we delve into our main results. In Section 1.1, we provide some intuition for our Spectrum-Aware construction using the example of the ridge estimator, since it admits a closed form and is simple to study. In Section 1.2, we discuss right-rotationally invariant designs and related literature, with the goal of motivating why this assumption is natural for debiasing, and emphasizing its degree of generality compared to prior Gaussian assumptions. In Section 1.3, we describe how the debiasing methods tend to fail when the design $\mathbf{X}$ contains outlier eigenvalues and/or the signal aligns with some eigenvectors of $\mathbf{X}$. In Section 1.4, we discuss a novel PCR-Spectrum-Aware Debiasing approach which addresses the aforementioned two issues and an associated hypothesis test for alignment between signal and principal components.
1.1 Intuition via ridge estimator
To motivate Spectrum-Aware Debiasing, let us focus on the simple instance of a ridge estimator that admits the closed-form
$\hat{\bm{\beta}}=\left(\mathbf{X}^{\top}\mathbf{X}+\lambda_{2}\mathbf{I}_{p}\right)^{-1}\mathbf{X}^{\top}\mathbf{y},\quad\lambda_{2}>0.$ (5)
Recall that we seek a debiased estimator of the form $\hat{\bm{\beta}}^{u}=\hat{\bm{\beta}}+\widehat{\mathsf{adj}}^{-1}\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})$. Suppose we plug in (5), leaving $\widehat{\mathsf{adj}}$ unspecified for the moment. If we denote the singular value decomposition of $\mathbf{X}$ to be $\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}$, we obtain that
$\mathbb{E}[\hat{\bm{\beta}}^{u}\mid\mathbf{X},\bm{\beta}^{\star}]=\underbrace{\left[\left(1+\frac{\lambda_{2}}{\widehat{\mathsf{adj}}}\right)\sum_{i=1}^{p}\left(\frac{d_{i}^{2}}{d_{i}^{2}+\lambda_{2}}\right)\mathbf{o}_{i}\mathbf{o}_{i}^{\top}\right]}_{=:\mathbf{V}}\bm{\beta}^{\star},$ (6)
where $\mathbf{o}_{i}^{\top}\in\mathbb{R}^{p}$ denotes the $i$-th row of $\mathbf{O}$ and recall that $d_{i}^{2}$’s denote the eigenvalues of $\mathbf{X}^{\top}\mathbf{X}$.
For $\hat{\bm{\beta}}^{u}$ to be unbiased, it appears necessary to choose $\widehat{\mathsf{adj}}$ so that it centers $\mathbf{V}$ around the identity matrix $\mathbf{I}_{p}$. We thus choose $\widehat{\mathsf{adj}}$ to be solution to the equation
$\left(1+\frac{\lambda_{2}}{\widehat{\mathsf{adj}}}\right)\frac{1}{p}\sum_{i=1}^{p}\frac{d_{i}^{2}}{d_{i}^{2}+\lambda_{2}}=1.$ (7)
This choice guarantees that the average of the eigenvalues of $\mathbf{V}$ equals 1. Solving for $\widehat{\mathsf{adj}}$, we obtain
$\widehat{\mathsf{adj}}=\left(\quantity(\frac{1}{p}\sum_{i=1}^{p}\frac{\lambda_{2}d_{i}^{2}}{d_{i}^{2}+\lambda_{2}})^{-1}-\frac{1}{\lambda_{2}}\right)^{-1}.$ (8)
This is precisely our Spectrum-Aware adjustment formula for the ridge estimator! However, it is not hard to see that centering $\mathbf{V}$ does not guarantee debiasing in general: for instance, $\hat{\bm{\beta}}^{u}$ would have an inflation bias if $\bm{\beta}^{\star}$ completely aligns with the top eigenvector $\mathbf{o}_{1}$. To ensure suitable debiasing, one requires $\mathbf{X}$ and $\bm{\beta}^{\star}$ to satisfy additional structure. To this end, if we further assume that $\mathbf{O}$ is random, independent of $\bm{\beta}^{\star}$, and satisfies $\mathbb{E}\left(\mathbf{o}_{i}\mathbf{o}_{i}^{\top}\right)=\frac{1}{p}\cdot\mathbf{I}_{p}.$ we would obtain, after choosing $\widehat{\mathsf{adj}}$ following (8), that
$\mathbb{E}\left[\hat{\bm{\beta}}^{u}\mid\bm{\beta}^{\star}\right]$ $=\mathbb{E}\underbrace{\left[\left(1+\frac{\lambda_{2}}{\widehat{\mathsf{adj}}}\right)\sum_{i=1}^{p}\left(\frac{d_{i}^{2}}{d_{i}^{2}+\lambda_{2}}\right)\mathbf{o}_{i}\mathbf{o}_{i}^{\top}\right]}_{=:\mathbf{V}}\bm{\beta}^{\star}\stackrel{{\scriptstyle(\star)}}{{=}}\bm{\beta}^{\star},$ (9)
This motivates us to impose the following assumption on $\mathbf{O}$.
Assumption 1.
$\mathbf{O}$ is drawn uniformly at random from the set of all orthogonal matrices of dimension $p$, independent of $\bm{\beta}^{\star}$ (this is the orthogonal group of dimension $p$ that we denote as $\mathbb{O}(p)$), in other words, $\mathbf{O}$ is drawn from the Haar measure on $\mathbb{O}(p)$.
We operate under this assumption since it ensures ()($\star$)( ⋆ ) holds and our Spectrum-Aware adjustment turns out to be the correct debiasing strategy in this setting. Meanwhile, the degrees-of-freedom adjustment [12] yields the correction factor
$\breve{\mathsf{adj}}=1-n^{-1}\operatorname{Tr}\bigg{(}\mathbf{X}\left(\mathbf{X}^{\top}\mathbf{X}+\lambda_{2}\mathbf{I}_{p}\right)^{-1}\mathbf{X}^{\top}\bigg{)}=1-\frac{1}{n}\sum_{i=1}^{p}\frac{d_{i}^{2}}{d_{i}^{2}+\lambda_{2}}.$
Notably, $\widehat{\mathsf{adj}}$ and $\breve{\mathsf{adj}}$ may be quite different. Unlike $\widehat{\mathsf{adj}}$, $\breve{\mathsf{adj}}$ may not center the spectrum of $\mathbf{V}$, and does not yield $\mathbb{E}(\hat{\bm{\beta}}^{u}\mid\bm{\beta}^{\star})=\bm{\beta}^{\star}$ in general. However, it is important to note that they coincide asymptotically and $\breve{\mathsf{adj}}$ would provide accurate debiasing if one assumes that the empirical distribution of $\left(d_{i}^{2}\right)_{i=1}^{p}$ converges weakly to the Marchenko-Pastur law (cf. Section A.4 from Appendix), a property that many design matrices do not satisfy. In other words, Degrees-of-Freedom Debiasing is sub-optimal in the sense that it implicitly makes the assumption that the spectrum of $\mathbf{X}^{\top}\mathbf{X}$ converges to the Marchenko-Pastur law, rather than using the actual spectrum. We provide examples of designs where Degrees-of-Freedom Debiasing fails in Figure 1. In contrast, $\widehat{\mathsf{adj}}$ is applicable under much broader settings as it accounts for the actual spectrum of $\mathbf{X}^{\top}\mathbf{X}$. Figure 1 shows the clear strengths of our approach over Degrees-of-Freedom Debiasing.
1.2 Right-rotationally invariant designs
Roughly speaking, Assumption 1 lands $\mathbf{X}$ in the class of right-rotationally invariant designs (Definition 2.1). Our arguments in the preceding section indicate that right-rotational invariance is a more fundamental assumption for debiasing than prior Gaussian/sub-Gaussian assumptions in the literature. Since it preserves the spectral information of $\mathbf{X}^{\top}\mathbf{X}$, we expect methods developed under this assumption to exhibit improved robustness when applied to designs that may not satisfy the right-rotational invariance property or designs observed in real data. We demonstrate this via Figure 3, bottom-left, where we conduct experiments with our PCR-Spectrum-Aware Debiasing on designs arising from six real datasets spanning image data, financial data, socio-economic data, and so forth.
Indeed varied research communities realized the strength of such designs, as demonstrated by the recent surge in literature in this space [85, 92, 86, 26, 27, 74, 108, 77, 6, 71, 72, 64, 73, 88, 62, 97, 34, 59, 76]. In particular, [27, 96] established that properties of high-dimensional systems proven under such designs continue to hold for a broad class of designs (including nearly deterministic designs as observed in compressed sensing [23]) as long as they satisfy certain spectral properties. In fact, the universality class for such designs is far broader than that for Gaussians, suggesting that these may serve as a solid prototype for modeling high-dimensional phenomena arising in non-Gaussian data. Despite such exciting developments, there are hardly any results when it comes to debiasing or inference under such designs (with the exception of [84] which we discuss later). This paper develops this important theory and methodology.
Despite the generality of right-rotationally invariant designs, studying these presents new challenges. For starters, analogs of the leave-one-out approach [67, 66, 9, 29, 104, 28, 82, 81, 21, 56] and Stein’s method [79, 20, 12, 10, 11, 2], both of which form fundamental proof techniques for Gaussian designs, are nonexistent or under-developed for this more general class. To mitigate this issue, we resort to an algorithmic proof strategy that the senior authors’ earlier work and that of others have used in the context of Gaussian designs. To study $\hat{\bm{\beta}}^{u}$, we observe that it depends on the regularized estimator $\bm{\hat{\beta}}$. However, $\bm{\hat{\beta}}$ does not admit a closed form in general, thus studying these turns out difficult. To circumvent this, we introduce a surrogate estimator that approximates $\bm{\hat{\beta}}$ in a suitable high-dimensional sense. We establish refined properties of these surrogates and use their closeness to our estimator of interest to infer properties about the latter. Prior literature has invoked this algorithmic proof strategy for Gaussian designs under the name of approximate message passing theory [24, 7, 82, 81, 107]. In case of right-rotationally invariant designs, we create the surrogate estimators using vector approximate message passing (VAMP) algorithms [74] (see details in Section A.6 and B.1 from Appendix). However, unlike the Gaussian case, proving that these surrogates approximate $\bm{\hat{\beta}}$ presents deep challenges. We overcome this via developing novel properties of VAMP algorithms that can be of independent interest to the signal processing [85], probability [92], statistical physics [86], information and coding theory [74, 73, 97] communities that seek to study models with right-rotationally invariant designs in the context of myriad other problems.
Among the literature related to right-rotationally invariant designs, two prior works are the most relevant for us. Of these, [41] initiated a study of the risk of $\bm{\hat{\beta}}$ under right-rotationally invariant designs using the VAMP machinery. However, their characterization is partially heuristic, meaning that they assume certain critical exchange of limits is allowed and that limits of certain fundamental quantities exist. The former assumption may often not hold, and the latter is unverifiable without proof (see Remark 5.1 for further details). As a by-product of our work on debiasing, we provide a complete rigorous characterization of the risk of regularized estimators under right-rotationally invariant designs (Theorem 5.1) without these unverifiable assumptions. The second relevant work is [84], which conjectures a population version of a debiasing formula for the Lasso using non-rigorous statistical physics tools. To be specific, they conjecture a debiasing formula that involves unknown parameters related to the underlying limiting spectral distribution of the sample covariance matrix. This formula does not provide an estimator that can be calculated from the observed data. In contrast, we develop a complete data-driven pipeline for debiasing and develop a consistent estimator for its asymptotic variance.
1.3 Practical issues and PCR-Spectrum-Aware Debiasing
Our discussion in Section 1.1 precludes two crucial settings that could occur in practice. Continuing our discussion on ridge regression, recall that $\mathbb{E}[\hat{\bm{\beta}}^{u}\mid\bm{\beta}^{\star}]=\mathbb{E}[\mathbf{V}\bm{\beta}^{\star}]$ for $\mathbf{V}$ defined in (6), and we chose $\widehat{\mathsf{adj}}$ to center the spectrum of $\mathbf{V}$ at 1 so that $\mathbb{E}[\mathbf{V}]=\mathbf{I}_{p}$ under our assumptions. Thus our choice of $\widehat{\mathsf{adj}}$ leads to the following,
$\mathbf{V}\approx\mathbf{I}_{p}+\mathsf{unbiased\;component}.$ (10)
This ensures that $\hat{\bm{\beta}}^{u}$ remains centered around $\bm{\beta}^{\star}$. However, to achieve this, we implicitly assumed that $\bm{\beta}^{\star}$ does not align with any of the $\mathbf{o}_{i}$’s. Potential issues may arise when this assumption is violated. For instance, if $\bm{\beta}^{\star}$ perfectly aligns with the top eigenvector $\mathbf{o}_{1}$, we would obtain
$\mathbb{E}[\hat{\bm{\beta}}^{u}\mid\mathbf{X},\bm{\beta}^{\star}]=\left(\frac{1}{p}\sum_{i=1}^{p}\frac{d_{i}^{2}}{d_{i}^{2}+\lambda_{2}}\right)^{-1}\frac{d_{1}^{2}}{d_{1}^{2}+\lambda_{2}}\bm{\beta}^{\star}.$
This results in an inflation bias since $\frac{d_{1}^{2}}{d_{1}^{2}+\lambda_{2}}>\frac{1}{p}\sum_{i=1}^{p}\frac{d_{i}^{2}}{d_{i}^{2}+\lambda_{2}}.$ Similar problems arise if $\bm{\beta}^{\star}$ aligns with other eigenvectors, and the resulting bias could lead to inflation or shrinkage depending on the set of aligned eigenvectors. We refer to this as the alignment issue.
Another common issue arises when the top few eigenvalues of the sample covariance matrix $\mathbf{X}^{\top}\mathbf{X}$ are significantly separated from the bulk of the spectrum. In this case, after centering the spectrum of $\mathbf{V}$, the variance of the “$\mathsf{unbiased\;component}$” in (10) will be large, making the debiasing procedure unstable. We refer to these eigenvalues as outlier eigenvalues.
In practice, these issues often arise simultaneously due to a small number of dominant principal components (PCs) that align with the signal. These PCs tend to distort desirable statistical properties that underlie Spectrum-Aware Debiasing. To address this, we propose a PCR-Spectrum-Aware Debiasing framework that integrates ideas from Principal Components Regression (PCR) with Spectrum-Aware Debiasing. In this enhanced method, we employ PCR to handle the outlier eigenvalues and the aligned eigenvectors, and then use Spectrum-Aware Debiasing on a transformed version of the original data to correct for shrinkage bias incurred from discarding low-variance PCs. We observe that this hybrid PCR-Spectrum-Aware approach works exceptionally well in challenging settings where alignment and outlier eigenvalue issues may both occur. In Figure 3, we demonstrate the efficacy of our PCR-Spectrum-Aware approach in situations with extremely strong correlations, heterogeneities, and heavy tails in the design matrix.
1.4 Notable outcomes: Alignment Testing and Debiased PCR
Our theory for Spectrum-Aware Debiasing has two significant by-products. In modern high-dimensional inference, calculating the precise asymptotic risk of regularized estimators has emerged as a prominent research area. Technical tools such as approximate message passing [24, 8, 7, 104, 52, 81, 5, 36], the convex Gaussian min-max theorem [91, 80], and the cavity or leave-one-out method [67, 89, 29, 28, 9, 82, 81, 21, 56] have proven invaluable for this purpose. These tools have facilitated the discovery of novel high-dimensional phenomena that other mathematical techniques simply fail to capture [22, 29, 9, 28, 82, 81, 17, 107, 61, 60, 56, 109]. Consequently, they have inspired new high-dimensional estimators that outperform traditional ones by a margin [81, 19]. Despite such remarkable progress, these technical tools suffer a crucial limitation. They typically assume that the design matrices are random and independent of the true signal, implying that the PCs are random vectors in generic position relative to the true signal. To the best of our knowledge, a principled test to validate this assumption has so far eluded the literature. In this paper, we introduce the first formal hypothesis test for PC-signal alignment, utilizing our PCR-Spectrum-Aware approach (see Corollary 4.3 and the subsequent discussion). We hope this serves as a foundation for more systematic investigations into this issue, thereby enhancing the applicability of statistical methods developed based on approximate message passing algorithms/leave-one-out/convex Gaussian min-max theorems.
As a second outcome, our work contributes to an extensive and growing body of work on PCR methodologies [57, 49, 4, 47, 30, 1, 46, 106, 50, 78, 83, 15]. Similar to the ridge and Lasso estimators, the traditional PCR estimator exhibits shrinkage bias due to the discarding of low-variance PCs [35, 38, 39, 14, 25, 58]. To the best of our knowledge, no previous work has investigated how this bias can be eliminated in high dimensions and its implications for inference. We develop the first approach for debiasing the classical PCR estimator, complete with formal high-dimensional guarantees. We will next formally introduce Spectrum-Aware Debiasing and discuss its properties.
1.5 Organization
We organize the rest of the paper as follows. In Section 2, we introduce our assumptions and preliminaries. In Sections 3 and 4, we introduce our Spectrum-Aware and PCR-Spectrum-Aware methods with formal guarantees. In Section 5, we present our proof outline and technical novelties. Finally in Section 6, we conclude with potential directions for future work.
2 Assumptions and Preliminaries
In this section, we introduce our assumptions and preliminaries that we require for the sequel.
2.1 Design matrix, signal and noise
We first formally define right-rotationally invariant designs.
Definition 2.1 (Right-rotationally invariant designs).
Consider the singular value decomposition $\mathbf{X}=\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}$ where $\mathbf{Q}\in\mathbb{R}^{n\times n}$ and $\mathbf{O}\in\mathbb{R}^{p\times p}$ are orthogonal and $\mathbf{D}\neq 0\in\mathbb{R}^{n\times p}$ is diagonal. We say a design matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$ is right-rotationally invariant if $\mathbf{Q},\mathbf{D}$ are deterministic, and $\mathbf{O}$ is uniformly distributed on the orthogonal group.
We work in a high-dimensional regime where $p$ and $n(p)$ both diverge and $n(p)/p\to\delta\in(0,+\infty)$. Known as proportional asymptotics, this regime has gained increasing popularity in recent times owing to the fact that asymptotic results derived under this assumption demonstrate remarkable finite sample performance (cf. extensive experiments in [82, 81, 17, 107, 61, 56] and the references cited therein). In this setting, we consider a sequence of problem instances $\quantity{\mathbf{y}(p),\mathbf{X}(p),\bm{\beta}^{\star}(p),\bm{\varepsilon}(p)}_{p\geq 1}$ such that $\mathbf{y}(p),\bm{\varepsilon}(p)\in\mathbb{R}^{n(p)},\mathbf{X}(p)\in\mathbb{R}^{n(p)\times p},\bm{\beta}^{\star}(p)\in\mathbb{R}^{p}$ and $\mathbf{y}(p)=\mathbf{X}(p)\bm{\beta}^{\star}(p)+\bm{\varepsilon}(p)$. In the sequel, we drop the dependence on $p$ whenever it is clear from context.
For a vector $\bm{v}\in\mathbb{R}^{p}$, we call its empirical distribution to be the probability distribution that puts equal mass $1/p$ to each coordinate of the vector. Some of our convergence results will be in terms of empirical distributions of sequences of random vectors. Specifically, we will use the notion of Wasserstein-2 convergence frequently so we introduce this next.
Definition 2.2 (Convergence of empirical distribution under Wasserstein-2 distance).
For a matrix $\left(\mathbf{v}_{1},\ldots,\mathbf{v}_{k}\right)=$ $\left(v_{i,1},\ldots,v_{i,k}\right)_{i=1}^{n}\in\mathbb{R}^{n\times k}$ and a random vector $\left(\mathsf{V}_{1},\ldots,\mathsf{V}_{k}\right)$, we write
$\left(\mathbf{v}_{1},\ldots,\mathbf{v}_{k}\right)\stackrel{{\scriptstyle W_{2}}}{{\rightarrow}}\left(\mathsf{V}_{1},\ldots,\mathsf{V}_{k}\right)$
to mean that the empirical distribution of the columns of $\left(\mathbf{v}_{1},\ldots,\mathbf{v}_{k}\right)$ converge to $(\mathsf{V}_{1},\ldots,\mathsf{V}_{k})$ in Wasserstein-$2$ distance. This means that for any continuous function $f:\mathbb{R}^{k}\rightarrow\mathbb{R}$ satisfying
$\left|f\left(\mathbf{v}_{1},\ldots,\mathbf{v}_{k}\right)\right|\leq C\left(1+\left\|\left(\mathbf{v}_{1},\ldots,\mathbf{v}_{k}\right)\right\|^{2}\right)$ (11)
for some $C>0$, we have
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}f\left(v_{i,1},\ldots,v_{i,k}\right)=\mathbb{E}\left[f\left(\mathsf{V}_{1},\ldots,\mathsf{V}_{k}\right)\right],$
where $\mathbb{E}\left[\left\|\left(\mathsf{V}_{1},\ldots,\mathsf{V}_{k}\right)\right\|^{2}\right]<\infty$. See in Section A.1 from Appendix for a review of the properties of the Wasserstein-2 convergence.
Assumption 2 (Measurement matrix).
We assume that $\mathbf{X}\in\mathbb{R}^{n\times p}$ is right-rotationally invariant (Definition 2.1) and independent of $\bm{\varepsilon}$. For the eigenvalues, we assume that as $n,p\rightarrow\infty$,
$\mathbf{d}:=\mathbf{D}^{\top}\bm{1}_{n\times 1}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{D},$ (12)
where $\mathsf{D}^{2}$ has non-zero mean with compact support $\operatorname{supp}(\mathsf{D}^{2})\subseteq[0,\infty)$. We denote $d_{-}:=\min(x:x\in\operatorname{supp}(\mathsf{D}^{2}))$. Furthermore, we assume that as $p\to\infty$,
$d_{+}:=\lim_{p\to\infty}\max_{i\in[p]}d_{i}^{2}<+\infty.$ (13)
Remark 2.3.
The constraint (13) states that $\mathbf{X}^{\top}\mathbf{X}$ has bounded operator norm. It has important practical implications. It prevents the occurrence of outlier eigenvalues, where a few prominent eigenvalues of $\mathbf{X}^{\top}\mathbf{X}$ deviate significantly from the main bulk of the spectrum.
We work with Assumption 2 for part of the sequel, in particular, Section 3. But later in Section 4, we relax restriction (13).
Since our debiasing procedure relies on the spectrum of $\mathbf{X}^{\top}\mathbf{X}$, analyzing its properties requires a thorough understanding of the properties of $\mathsf{D}$ (from (12)), the limit of the empirical spectral distribution of $\mathbf{X}^{\top}\mathbf{X}$. Often these properties can be expressed using two important quantities—the Cauchy and the R-transform. We define these next. For technical reasons, we will define these transforms corresponding to the law of $-\mathsf{D}^{2}$.
Definition 2.4 (Cauchy- and R-transform).
Under Assumption 2, let $G:(-d_{-},\infty)\to(0,\infty)$ and $R:(0,G(-d_{-}))\to(-\infty,0)$ be the Cauchy- and R-transforms of the law of $-\mathsf{D}^{2}$, defined as
$G(z)=\mathbb{E}\left[\frac{1}{z+\mathsf{D}^{2}}\right],\qquad R(z)=G^{-1}(z)-\frac{1}{z},$ (14)
where $G^{-1}(\cdot)$ is the inverse function of $G(\cdot)$. See properties and well-definedness of these in Lemma A.8 from Appendix. We set $G(-d_{-})=\lim_{z\to-d_{-}}G(z)$.
We next move to discussing our assumptions on the signal.
Assumption 3 (Signal and noise).
We assume throughout that $\bm{\varepsilon}\sim N(0,\sigma^{2}\cdot\mathbf{I}_{p})$ for potentially unknown noise level $\sigma^{2}>0$. We require that $\bm{\beta}^{\star}$ is either deterministic or independent of $\mathbf{O},\bm{\varepsilon}$. In the former case, we assume that $\bm{\beta}^{\star}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{B}^{\star}$ where $\mathsf{B}^{\star}$ is a random variable with finite variance. In the latter case, we assume the same convergence holds almost surely.
Remark 2.5.
The independence condition between $\bm{\beta}^{\star}$ and $\mathbf{O}$, along with the condition that $\mathbf{O}$ is uniformly drawn from the orthogonal group enforces that $\bm{\beta}^{\star}$ cannot align with a small number of these eigenvectors. Once again, we require these assumptions in Section 3 but we relax these later in Section 4.
Remark 2.6.
We believe the assumption on the noise can be relaxed in many settings. For instance, if we assume $\mathbf{Q}$ (Definition 2.1) to be uniformly distributed on the orthogonal group independent of $\mathbf{O}$ and $\bm{\beta}^{\star}$, one may work with the relaxed assumption that $\bm{\varepsilon}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{E}$ for any random variable $\mathsf{E}$ with mean 0 and variance $\sigma^{2}$. This encompasses many noise distributions beyond Gaussians. Even without such an assumption on $\mathbf{Q}$, allowing for sub-Gaussian noise distributions should be feasible invoking universality results. However, in this paper, we prefer to focus on fundamentally breaking the i.i.d. Gaussian assumptions on $\mathbf{X}$ in prior works. In this light, we work with the simpler Gaussian assumption on the noise.
In the next segment, we describe the penalty functions that we work with.
2.2 Penalty function
As observed in the vast majority of literature on high-dimensional regularized regression, the proximal map of the penalty function plays a crucial role in understanding properties of $\bm{\hat{\beta}}$. We introduce this function next.
Let the proximal map associated to $h$ be
$\forall v>0,x,y\in\mathbb{R},\quad\operatorname{Prox}_{vh}(x)\equiv\underset{y\in\mathbb{R}}{\arg\min}\left\{h(y)+\frac{1}{2v}(y-{x})^{2}\right\}.$
Assumption 4 (Penalty function).
We assume that $h:\mathbb{R}\mapsto[0,+\infty)$ is proper, closed and satisfies that for some $c_{0}\geq 0,\forall x,y\in\mathbb{R},t\in[0,1]$,
$h(t\cdot x+(1-t)\cdot y)\leq t\cdot h(x)+(1-t)\cdot h(y)-\frac{1}{2}c_{0}\cdot t(1-t)\cdot(x-y)^{2}.$ (15)
Here, $c_{0}=0$ is equivalent to assuming $h$ is convex and $c_{0}>0$ to assuming $h$ is $c_{0}$-strongly convex. Furthermore, we assume that $h(x)$ is twice continuously differentiable except for a finite set $\mathfrak{D}$ of points, and that $h^{\prime\prime}(x)$ and $\operatorname{Prox}_{vh}^{\prime}(x)$ have been extended at their respective undefined points using Lemma 2.7 below.
Lemma 2.7 (Extension at non-differentiable points).
Fix any $v>0$. Under Assumption 4, $x\mapsto\operatorname{Prox}_{vh}(x)$ is continuously differentiable at all but a finite set $\mathcal{C}$ of points. Extending functions $x\mapsto h^{\prime\prime}(x)$ and $x\mapsto\operatorname{Prox}_{vh}^{\prime}(x)$ on $\mathfrak{D}$ and $\mathcal{C}$ by +$\infty$+ ∞ and 00 respectively, we have that for all $x\in\mathbb{R}$,
$\operatorname{Prox}_{vh}^{\prime}(x)=\frac{1}{1+vh^{\prime\prime}\left(\operatorname{Prox}_{vh}(x)\right)}\in\quantity[0,\frac{1}{1+vc_{0}}],\quad h^{\prime\prime}(x)\in[c_{0},+\infty].$ (16)
After the extension, for any $w>0$, $x\mapsto\frac{1}{w+h^{\prime\prime}\left(\operatorname{Prox}_{vh}(x)\right)}$ is piecewise continuous with finitely many discontinuity points on which it takes value 00.
We defer the proof to Section A.2 in Appendix. We considered performing this extension since our debiasing formula involves the second derivative of $h(\cdot)$. The extension allows us to handle cases where the second derivative may not exist everywhere. As an example, we compute the extension for the elastic net penalty and demonstrate the form our debiasing formula takes after plugging in this extended version of $h(\cdot).$
Example 2.8 (Elastic Net penalty).
Consider the elastic-net penalty
$h(x)=\lambda_{1}|x|+\frac{\lambda_{2}}{2}x^{2},\lambda_{1}\geq 0,\lambda_{2}\geq 0.$ (17)
This is twice continuously differentiable except at $x=0$ (i.e. $\mathfrak{D}=\quantity{0}$). Fix any $v>0$. Its $\operatorname{Prox}_{vh}(x)=\frac{1}{1+\lambda_{2}v}\operatorname{ST}_{\lambda_{1}v}\left(x\right)$ is continuously differentiable except at $x=$$\pm\lambda_{1}v$. Here, $\mathrm{ST}_{\lambda v}(x):=\operatorname{sgn}(x)(|x|-\lambda v)_{+}$ is the soft-thresholding function. Per Lemma 2.7, the extended $h^{\prime\prime},\operatorname{Prox}_{vh}^{\prime}$ are
$h^{\prime\prime}(x)=\left\{\begin{array}[]{c}+\infty,\text{ if }x=0\\ \lambda_{2},\text{ otherwise }\end{array}\right.,\quad\operatorname{Prox}_{vh}^{\prime}(x)=\frac{1}{1+\lambda_{2}v}\mathbb{I}\left(|x|>\lambda_{1}v\right)$
respectively, so that (16) holds for all $x\in\mathbb{R}$. Note also that for any $w>0,x\mapsto\frac{1}{1+wh^{\prime\prime}\left(\operatorname{Prox}_{vh}(x)\right)}=\frac{1}{1+\lambda_{2}w}\mathbb{I}(|x|>\lambda_{1}v)$ is piecewise continuous and takes value 0 on both of its discontinuity points. It follows that our adjustment (4) can be written as
$\frac{1}{p}{{\sum}}_{i=1}^{p}\frac{1}{\left(d_{i}^{2}\widehat{\mathsf{adj}}^{-1}-1\right)\left(\frac{\hat{s}}{p}\quantity(1+\widehat{\mathsf{adj}}^{-1}\lambda_{2})^{-1}\right)+1}=1,$ (18)
where $\hat{s}=\left|\left\{j:\hat{{\beta}}_{j}\neq 0\right\}\right|$.
As a sanity check, if one sets $\lambda_{2}=0$ and solves the population version of the above equation
$\mathbb{E}\frac{1}{\left(\mathsf{D}^{2}\widehat{\mathsf{adj}}^{-1}-1\right)\cdot\frac{\hat{s}}{p}+1}=1$ (19)
with $\mathsf{D}^{2}$ drawn from the Marchenko-Pastur law, then one recovers the well-known degrees-of-freedom adjustment for the Lasso: $\widehat{\mathsf{adj}}=1-\hat{s}/n.$
The following assumption is analogous to [12, Assumption 3.1] for the Gaussian design: we require either $h$ to be strongly convex or $\mathbf{X}^{\top}\mathbf{X}$ to be non-singular with smallest eigenvalues bounded away from 00.
Assumption 5.
Either $c_{0}>0$ or $\lim_{p\to\infty}\min_{i\in p}(d_{i}^{2})\geq c_{1}$ for some constant $c_{1}>0$.
2.3 Fixed-point equation
Recall our discussion from Section 1.2. We study the regularized estimator $\bm{\hat{\beta}}$ by introducing a more tractable surrogate $\bm{\hat{\beta}}^{t}$. Later we will see that we construct this surrogate using an iterative algorithmic scheme known as Vector Approximate Message Passing algorithm (VAMP) [74]. Thus to study the surrogate, one needs to study the VAMP algorithm carefully. One can describe the properties of this algorithm using a system of fixed point equations in four variables. We use $\gamma_{*},\eta_{*},\tau_{*},\tau_{**},\in(0,+\infty)$ to denote these variables, and define the system here:
$\frac{\gamma_{*}}{\eta_{*}}=\mathbb{E}\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}\left(\mathsf{B}^{\star}+\sqrt{\tau_{*}}\mathsf{Z}\right)$ (20a)
$\tau_{**}=\frac{\eta_{*}^{2}}{\left(\eta_{*}-\gamma_{*}\right)^{2}}\left[\mathbb{E}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\mathsf{B}^{\star}+\sqrt{\tau_{*}}\mathsf{Z}\right)-\mathsf{B}^{\star}\right)^{2}-\left(\frac{\gamma_{*}}{\eta_{*}}\right)^{2}\tau_{*}\right]$ (20b)
$\gamma_{*}=-R\left(\eta_{*}^{-1}\right)$ (20c)
$\tau_{*}=\left(\frac{\eta_{*}}{\gamma_{*}}\right)^{2}\left[\mathbb{E}\left[\frac{\sigma^{2}\mathsf{D}^{2}+\tau_{**}\left(\eta_{*}-\gamma_{*}\right)^{2}}{\left(\mathsf{D}^{2}+\eta_{*}-\gamma_{*}\right)^{2}}\right]-\left(\frac{\eta_{*}-\gamma_{*}}{\eta_{*}}\right)^{2}\tau_{**}\right],$ (20d)
where $\mathsf{Z}\sim N(0,1)$ is independent of $\mathsf{B}^{\star}$. We remind the reader that $x\mapsto\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}(x)$ is well-defined on $\mathbb{R}$ by the extension described in Lemma 2.7.
The following assumption ensures that at least one solution exists.
Assumption 6 (Existence of fixed points).
There exists a solution $\gamma_{*},\eta_{*},\tau_{*},\tau_{**}\in(0,+\infty)$ such that (20) holds.
Remark 2.9 (Existence implies uniqueness).
Under Assumptions 25, the existence of a solution implies uniqueness, as we show in Section A.5.2 from Appendix.
The following Proposition shows that Assumption 6 holds for any Elastic Net penalties (that are not Lasso). See proof in Section A.5.3 from Appendix.
Proposition 2.10 (Assumption 6 holds for Elastic Net).
Suppose that $\mathsf{D}^{2}$ and $\mathsf{B}^{\star}$ are as in Assumption 2 and 3. For the Elastic Net penalty $h$ defined in Example 2.8 with $\lambda_{2}>0$, Assumption 6 holds for any $\sigma^{2}>0$.
Remark 2.11 (Verifying Assumption 6).
Proving Assumption 6 is difficult for general penalties and $\mathsf{D}^{2}$. One therefore hopes to verify it on a case-by-case basis. For instance, when $\mathsf{D}^{2}$ follows the Marchenko-Pastur law and $h$ is the Lasso, [7, Proposition 1.3, 1.4] proved that (20) admits a unique solution. Similarly, [82] proved existence and uniqueness of the solution to an analogous fixed point equation in the context of logistic and probit regression. For general $\mathsf{D}^{2}$ in the context of linear regression, there are no rigorous results establishing that the system admits a solution. In Proposition 2.10, we take the first step and show for the first time that Assumption 6 indeed holds for the Elastic Net for a general $\mathsf{D}^{2}$ satisfying our assumptions. This result may be of independent interest: we anticipate that our proof can be adapted to other penalties beyond the Elastic Net; however, since our arguments require an explicit expression for the proximal operator, we limit our presentation to the Elastic Net. That said, one expects this to always hold true under Assumptions 25 (see e.g. [40, 41]).
Assumption 7 (Feasibility of noise-level estimation).
When the noise-level $\sigma^{2}$ is unknown, we require that $\gamma_{*},\eta_{*}$ defined in (20) and $\mathsf{D}^{2}$ defined in Assumption 2 satisfy
$\delta\cdot\frac{\mathbb{E}\mathsf{D}^{2}\cdot\left(1-\left(\frac{\eta_{*}-\gamma_{*}}{\mathsf{D}^{2}+\left(\eta_{*}-\gamma_{*}\right)}\right)^{2}\right)}{\mathbb{E}\mathsf{D}^{2}\cdot\mathbb{E}\left(1-\left(\frac{\eta_{*}-\gamma_{*}}{\mathsf{D}^{2}+\left(\eta_{*}-\gamma_{*}\right)}\right)^{2}\right)}\neq 1.$ (21)
Remark 2.12.
Assumption 7 serves as a technical condition to rule out degenerate scenarios where estimating $\sigma^{2}$ is impossible. For example, this condition is not satisfied when $n=p$ and $\mathbf{X}=\mathbf{I}_{p}$: in this case, our sole observation is $\mathbf{y}=\bm{\beta}^{\star}+\bm{\varepsilon}$ and it is indeed impossible to estimate $\sigma^{2}$. We provide a consistent estimator ifor the left-hand side of (21) in (39), facilitating the verification of Assumption 7.
3 Spectrum-Aware Debiasing
Recall that our debiasing formula involved $\widehat{\mathsf{adj}}$ obtained by solving (4). To ensure our estimator is well-defined, we need to establish that this equation has a unique solution. In this section, we address this issue, establish asymptotic normality of our debiased estimator (suitably centered and scaled), and present a consistent estimator for its asymptotic variance.
3.1 Well-definedness of our debiasing formula
To show that (4) admits a unique solution, we define the function $g_{p}:(0,+\infty)\mapsto\mathbb{R}$ as
$g_{p}(\gamma)=\frac{1}{p}{{\sum}}_{i=1}^{p}\frac{1}{\left(d_{i}^{2}-\gamma\right)\left(\frac{1}{p}\sum_{j=1}^{p}\frac{1}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}\right)+1}.$ (22)
Here $h^{\prime\prime}(\cdot)$ refers to the extended version we defined using Lemma 2.7.
The following Proposition is restated from Proposition B.10 in Appendix.
Proposition 3.1.
Fix $p\geq 1$ and suppose that Assumption 4 holds. Then, the function $\gamma\mapsto g_{p}(\gamma)$ is well-defined, strictly increasing for any $\gamma>0$, and
$g_{p}(\gamma)=1$ (23)
admits a unique solution in $(0,+\infty)$ if and only if there exists some $i\in[p]$ such that $h^{\prime\prime}(\hat{{\beta}}_{i})\neq+\infty$ and at least one of the following holds: (i) $\left\|h^{\prime\prime}(\hat{\bm{\beta}})\right\|_{0}=p$; (ii) $\mathbf{X}^{\top}\mathbf{X}$ is non-singular; (iii) $\norm{d}_{0}+\norm{h^{\prime\prime}(\hat{\bm{\beta}})}_{0}>p$
Example 3.2.
As a concrete example, the assumptions of Proposition 3.1 hold for the Elastic Net with $\lambda_{2}>0$.
Remark 3.3.
To find the unique solution of $g_{p}(\gamma)=1$, we recommend using Newton’s method initialized at $\gamma=\frac{1}{p}\sum_{i=1}^{p}d_{i}^{2}$. In rare cases where Newton’s method fails to converge, we suggest using a bisection-based method, such as the Brent’s method, to solve (4) on the interval $\left[0,\max_{i\in[p]}d_{i}^{2}\right]$, where convergence is guaranteed (by Jensen’s inequality, the solution must be upper bounded by $\max_{i\in[p]}d_{i}^{2}$). For numerical stability, we suggest re-scaling the design matrix $\mathbf{X}$ such that average of its eigenvalues equals 1, i.e. $\mathbf{X}_{\mathsf{rescaled}}\leftarrow\quantity(\frac{1}{p}\sum_{i=1}^{p}d_{i}^{2})^{-1/2}\cdot\mathbf{X}$.
3.2 The procedure
In this section, we introduce our Spectrum-Aware Debiasing procedure (Definition 3.4).
Definition 3.4 (Spectrum-Aware Debiasing).
Given observed data $(\mathbf{X},\mathbf{y})$ and a penalty function $h$, our procedure for Spectrum-Aware Debiasing computes the regularized estimator $\hat{\bm{\beta}}$ and eigenvalues $(d_{i}^{2})_{i=1}^{p}$ of the sample covariance matrix $\mathbf{X}^{\top}\mathbf{X}$. Subsequently, it solves for $\widehat{\mathsf{adj}}$ from
$\frac{1}{p}{{\sum}}_{i=1}^{p}\frac{1}{\left(d_{i}^{2}-\widehat{\mathsf{adj}}\right)\left(\frac{1}{p}\sum_{j=1}^{p}\quantity(\widehat{\mathsf{adj}}+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right))^{-1}\right)+1}=1.$ (24)
where $h^{\prime\prime}(\cdot)$ refers to the extended version we defined using Lemma 2.7 (see also Proposition 3.1 and Remark 3.3). Finally, we generate the debiased estimator as follows
$\hat{\bm{\beta}}^{u}=\hat{\bm{\beta}}+\widehat{\mathsf{adj}}^{-1}\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}}).$ (25)
3.3 Asymptotic normality
Theorem 3.1 below states that the empirical distribution of $(\hat{\tau}_{*}^{-1/2}(\hat{{\beta}}^{u}_{i}-{\beta}^{\star}_{i}))_{i=1}^{p}$ converges to a standard Gaussian.
Theorem 3.1 (Asymptotic normality of $\hat{\bm{\beta}}^{u}$).
Suppose that Assumption 27 hold. Then, we have that almost surely as $p\to\infty$,
$\hat{\tau}_{*}^{-1/2}(\hat{\bm{\beta}}^{u}-\bm{\beta}^{\star})\stackrel{{\scriptstyle W_{2}}}{{\to}}N(0,1).$
Next, we describe a consistent estimator for the asymptotic variance $\tau_{*}$. We require some intermediate quantities that depend on the observed data and the choice of the penalty. We define these next. Later in Section 5, we will provide intuition as to why these intermediate quantities are important and how we construct the variance estimator.
Definition 3.5 (Scalar statistics).
Let $\widehat{\mathsf{adj}}(\mathbf{X},\mathbf{y},h)\in(0,+\infty)$ be the unique solution to (4). We define the following scalar statistics
$\hat{\eta}_{*}(\mathbf{X},\mathbf{y},h)\leftarrow\left(\frac{1}{p}{{\sum}}_{j=1}^{p}\frac{1}{\widehat{\mathsf{adj}}+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}\right)^{-1}$ (26)
$\hat{\tau}_{**}(\mathbf{X},\mathbf{y},h)\leftarrow\frac{\left\|\quantity(\mathbf{I}_{n}+\frac{1}{\hat{\eta}_{*}-\widehat{\mathsf{adj}}}\mathbf{X}\mathbf{X}^{\top})\quantity(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})\right\|^{2}-n\hat{\sigma}^{2}}{\sum_{i=1}^{p}d_{i}^{2}}$
$\hat{\tau}_{*}(\mathbf{X},\mathbf{y},h)\leftarrow\frac{1}{p}{{\sum}}_{i=1}^{p}\frac{\hat{\eta}_{*}^{2}d_{i}^{2}\hat{\sigma}^{2}+\quantity(d_{i}^{2}-\widehat{\mathsf{adj}}+2\hat{\eta}_{*})\quantity(\widehat{\mathsf{adj}}-d_{i}^{2})\quantity(\hat{\eta}_{*}-\widehat{\mathsf{adj}})^{2}\hat{\tau}_{**}}{\quantity(d_{i}^{2}-\widehat{\mathsf{adj}}+\hat{\eta}_{*})^{2}\quantity(\widehat{\mathsf{adj}})^{2}}$
where $\hat{\sigma}^{2}$ is an estimator for the noise level $\sigma^{2}$ (see Remark 3.6 below). Note that the quantities in (26) are well-defined for any $p$ (i.e. no zero-valued denominators) if there exists some $i\in[p]$ such that $h^{\prime\prime}(\hat{\beta}_{i})\neq+\infty$ and there exists some $j\in[p]$ such that $h^{\prime\prime}(\hat{\beta}_{j})\neq 0$. Going forward, we suppress the dependence on $\mathbf{X},\mathbf{y},h$ for convenience.
Remark 3.6.
The computation of $\hat{\tau}_{*}$ and $\hat{\tau}_{**}$ in (26) requires an estimator $\hat{\sigma}^{2}$ for the noise level $\sigma^{2}$ when it is not already known. We provide a consistent estimator in (38) that we use in all our numerical experiments.
We illustrate Theorem 3.1 in Figure 1 under five different right-rotationally-invariant designs (cf. Remark E.2 in Appendix) with non-trivial correlation structures, and compare with Degrees-of-Freedom Debiasing with $\mathbf{M}=\mathbf{I}_{p}$. The corresponding QQ-plot can be found in Figure 5 in Appendix. We observe that our method outperforms Degrees-of-Freedom Debiasing by a margin.
We next develop a different result that characterizes the asymptotic behavior of finite-dimensional marginals of $\hat{\bm{\beta}}^{u}$. Corollary 3.8 below establishes this under an additional exchangeability assumption on $\bm{\beta}^{\star}$. To state the corollary, we recall to readers the standard definition of exchangeability for a sequence of random variables.
Definition 3.7 (Exchangeability).
We call a sequence of random variables $\left(\mathsf{V}_{i}\right)_{i=1}^{p}$ exchangeable if for any permutation $\pi$ of the indices $1,...,p$, the joint distribution of the permuted sequence $\left(\mathsf{V}_{\pi(i)}\right)_{i=1}^{p}$ is the same as the original sequence.
Corollary 3.8 below is a consequence of Theorem 3.1. We defer its proof to Section B.5 in Appendix. A numerical demonstration is included in Section E.3 from Appendix.
Corollary 3.8.
Fix any finite index set $\mathcal{I}\subset[p]$. Suppose that Assumption 27 hold, and $\left(\bm{\beta}^{\star}\right)_{j=1}^{p}$ is exchangeable independent of $\mathbf{X},\bm{\varepsilon}$. Then as $p\to\infty$, we have
$\frac{\hat{\bm{\beta}}^{u}_{\mathcal{I}}-\bm{\beta}^{\star}_{\mathcal{I}}}{\sqrt{\hat{\tau}_{*}}}\Rightarrow N(\rm{0},\mathbf{I}_{|\mathcal{I}|})$ (27)
where $\Rightarrow$ denotes weak convergence.
Corollary 3.8 is naturally useful for constructing confidence intervals for finite-dimensional marginals of $\bm{\beta}^{\star}$ with associated false coverage proportion guarantees.
3.4 Inference
Figure 2: The above plots the TPR and FPR of the hypothesis testing procedure defined in (28) with significance level $\alpha$ and FCP of the constructed confidence intervals (29) with $b=\Phi^{-1}(1-\alpha/2),a=\Phi^{-1}(\alpha/2)$ as $\alpha$ on the x-axis varies from 00 to $1$, for both degrees-of-freedom ($\mathsf{DF}$, blue) adjustment and Spectrum-Aware ($\mathsf{SA}$, red) adjustment. The setting here is the same as in Figure 1.
In this section, we discuss applications of our Spectrum-Aware Debiasing approach to hypothesis testing and construction of confidence intervals. Consider the null hypotheses $H_{i,0}:{\beta}^{\star}_{i}=0$ for all $i\in[p]$. We define $\mathrm{p}$-values $P_{i}$ and decision rule $T_{i}$ ($T_{i}=1$ means rejecting $H_{0,i}$) for the test $H_{0,i}$ based on the definitions
$P_{i}\quantity(\hat{{\beta}}^{u}_{i},\hat{\tau}_{*})=2\left(1-\Phi\left(\left|\frac{\hat{{\beta}}^{u}_{i}}{\sqrt{\hat{\tau}_{*}}}\right|\right)\right),\quad T_{i}(\hat{{\beta}}^{u}_{i},\hat{\tau}_{*})=\left\{\begin{array}[]{cc}1,&\text{if }P_{i}\quantity(\hat{{\beta}}^{u}_{i},\hat{\tau}_{*})\leq\alpha\\ 0,&\text{ if }P_{i}\quantity(\hat{{\beta}}^{u}_{i},\hat{\tau}_{*})>\alpha\end{array},\right.$ (28)
where $\Phi$ denotes the standard Gaussian CDF and $\alpha\in[0,1]$ is the significance level. We define the false positive rate (FPR) and true positive rate (TPR) below
$\mathsf{FPR}(p):=\frac{\sum_{j=1}^{p}\mathbb{I}\left(P_{j}\leq\alpha,\beta_{j}^{\star}=0\right)}{\sum_{j=1}^{p}\mathbb{I}\left(\beta_{j}^{\star}=0\right)},\quad\mathsf{TPR}(p):=\frac{\sum_{j=1}^{p}\mathbb{I}\left(P_{j}\leq\alpha,\left|\beta_{j}^{\star}\right|>0\right)}{\sum_{j=1}^{p}\mathbb{I}\left(\beta_{j}^{\star}>0\right)}$
when their respective denominators are non-zero. Fix $\alpha\in[0,1]$. We can construct confidence intervals
$\mathsf{CI}_{i}(\hat{{\beta}}^{u}_{i},\hat{\tau}_{*})=\left(\hat{{\beta}}^{u}_{i}+a\sqrt{\hat{\tau}_{*}},\hat{{\beta}}^{u}_{i}+b\sqrt{\hat{\tau}_{*}}\right),\qquad\forall i\in[p]$ (29)
for any $a,b\in\mathbb{R}$ such that $\Phi(b)-\Phi(a)=1-\alpha$. One can define the associated false coverage proportion (FCP)
$\mathsf{FCP}(p):=\frac{1}{p}\sum_{i=1}^{p}\mathbb{I}\left({\beta}^{\star}_{i}\notin\mathsf{CI}_{i}\right).$
for any $p\geq 1$. Theorem 3.1 directly yield guarantees on the FPR, TPR and FCP as shown in Corollary 3.9 below. We defer the proof to Section B.6 in Appendix.
Corollary 3.9.
Suppose that Assumption 26 hold. We have the following.
  • (a)
    Suppose that $\mathbb{P}\left(\mathsf{B}^{\star}=0\right)>0$ and there exists some $\mu_{0}\in(0,+\infty)$ such that
    (|$\mathbb{P}\left(\absolutevalue{\mathsf{B}^{\star}}\in\left(\mu_{0},+\infty\right)\cup\{0\}\right)=1.$
    Then for any fixed $i$ such that ${\beta}^{\star}_{i}=0$, we have $\lim_{p\rightarrow\infty}\mathbb{P}\left(T_{i}=1\right)=\alpha,$ and the false positive rate satisfies that almost surely $\lim_{p\to\infty}\mathsf{FPR}(p)=\alpha.$ Refer also to Remark B.14 from Appendix for the exact asymptotic limit of TPR.
  • (b
    The false coverage proportion satisfies that almost surely $\lim_{p\to\infty}\mathsf{FCP}(p)=\alpha$.
We demonstrate Corollary 3.9 in Figure 2. We note that the FPR and FCP values obtained from Degrees-of-Freedom Debiasing diverge from the intended $\alpha$ values, showing a clear misalignment with the 45-degree line. In contrast, the Spectrum-Aware Debiasing method aligns rather well with the specified $\alpha$ values, and this occurs without much compromise on the TPR level.
4 PCR-Spectrum-Aware Debiasing
4.1 Outliers and PC alignment
The assumptions made in our previous section exclude important scenarios where the design may contain outlier eigenvalues or the signal may align with an eigenvector of the sample covariance matrix. To address this challenge, we propose an enhanced Spectrum-Aware procedure whic relaxes Assumptions 2 and 3 to Assumption 8 below. To this end, denote $\mathcal{N}:=\left\{i\in[p]:d_{i}^{2}>0\right\},N:=|\mathcal{N}|$. We let $\mathcal{J}$ be a user-chosen, finite index set $\mathcal{J}\subseteq\mathcal{N}$ that contains $\mathcal{J}_{1}\cup\mathcal{J}_{2}$ as subset (See Remark 4.2). We denote its size as $J:=\absolutevalue{\mathcal{J}}$ |.
Assumption 8.
We assume that $\mathcal{J}$ is of finite size [^2] Finite size means that $J_{1}$ does not grow with $n,p$. and for some real-valued vectors $\bm{\upsilon}^{\star}\in\mathbb{R}^{J},\bm{\zeta}^{\star}\in\mathbb{R}^{p}$,
$\bm{\beta}^{\star}=\bm{\beta}^{\star}_{\mathsf{al}}+\bm{\zeta}^{\star},\qquad\bm{\beta}^{\star}_{\mathsf{al}}=\sum_{i=1}^{J}\upsilon^{\star}_{i}\cdot\mathbf{o}_{\mathcal{J}(i)}.$ (30)
where we used $\mathcal{J}(i)$ to denote the $i$-th index in $\mathcal{J}$. Both $\bm{\upsilon}^{\star}$ and $\bm{\zeta}^{\star}$ are unknown, and they can be either deterministic or random independent of $\mathbf{O},\bm{\varepsilon}$. If $\bm{\zeta}^{\star}$ is deterministic, we assume that $\bm{\zeta}^{\star}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{C}^{\star}$ as $n,p\to\infty$, where $\mathsf{C}^{\star}$ is a random variable with finite variance. If $\bm{\zeta}^{\star}$ is random, we assume the same convergence holds almost surely. Furthermore, we assume that Assumption 2 holds except that, instead of (13), we only require $\lim_{p\to\infty}\max_{i\in[p]\setminus\mathcal{J}}d_{i}^{2}<+\infty.$
Under Assumption 8, $\bm{\beta}^{\star}_{\mathsf{al}}$ is the alignment component that aligns to $\mathcal{J}(i)$-th Principal Component (PC) $\mathbf{o}_{\mathcal{J}(i)}$ if the corresponding $\upsilon^{\star}_{i}$ is non-zero, while $\bm{\zeta}^{\star}$ is the complement component that is independent of the design. Note that when $\mathcal{J}=\emptyset$, Assumption 8 reduces to Assumptions 2 and 3 precisely.
Remark 4.1.
Assumption 8 does not impose any constraints on $\bm{\upsilon}^{\star}\in\mathbb{R}^{J}$. For example, it is permitted that $\bm{\upsilon}^{\star}=0$ or that $p^{-1}\norm{\bm{\upsilon}^{\star}}^{2}$ diverges as $p\to\infty$. Note that Assumption 8 also permits $\bm{\zeta}^{\star}=0$ but $p^{-1}\norm{\bm{\zeta}^{\star}}^{2}$ cannot diverge.
Remark 4.2.
$\mathcal{J}$ needs to be a finite index set that contains both $\mathcal{J}_{1}$ and $\mathcal{J}_{2}$ as subsets. The index set $\mathcal{J}_{2}$ can be determined by observing the spectrum of $\mathbf{X}^{\top}\mathbf{X}$. $\mathcal{J}_{1}$ is generally not observed and requires prior information. However, we remark that eigenvectors indexed by $\mathcal{J}_{1}\cap\mathcal{J}_{2}$ tend to distort the debiasing procedure most severely. So often just including $\mathcal{J}_{2}$ (i.e. top few PCs) in $\mathcal{J}$ can significantly improve inference.
We develop a debiasing approach that recovers both components of $\bm{\beta}^{\star}$ from (30). Our approach uses classical PCR to consistently estimate the aligned component $\bm{\beta}^{\star}_{\mathsf{al}}$ and uses Spectrum-Aware Debiasing to produce a debiased estimator of $\bm{\zeta}^{\star}$.
4.2 The PCR algorithms
Figure 3: top-left: Same setting as Figure 1 except for specific changes to the design distribution parameters that lead to more challenging scenarios (see details of $\mathsf{MatrixNormal}$-$\mathsf{B}$,…,$\mathsf{MultiCauchy}$ in Definition E.4 from Appendix). Rows 1–4 correspond to: (i) DF: Degrees-of-Freedom Debiasing as described in [12]; (ii) SA: Spectrum-Aware Debiasing as described in Definition 3.4; (iii): PCRDF: PCR-Degrees-of-Freedom Debiasing, that is, the procedure obtained from Section 4.3.3 on substituting Degrees-of-Freedom Debiasing for Spectrum-Aware Debiasing in the complement PCR step; (iv) PCRSA: PCR-Spectrum-Aware Debiasing as described in Section 4.3.3. The true signals $\bm{\beta}^{\star}$ for these designs are generated such that they align with the 2nd, 4th, and 6th PCs. Concretely, we generate the signal as follows: $\bm{\beta}^{\star}=\bm{\beta}^{\star}_{\mathsf{al}}+\bm{\zeta}^{\star}$ where the components of ${\zeta}^{\star}_{i}$ are i.i.d. draws satisfying ${\zeta}^{\star}_{i}\sim 0.24\cdot N(-20,1)+0.06\cdot N(10,1)+0.7\cdot\delta_{0}$ and $\bm{\beta}^{\star}_{\mathsf{al}}=\sum_{i=1}^{J}\upsilon^{\star}_{i}\cdot\mathbf{o}_{\mathcal{J}(i)}$ with $\upsilon^{\star}_{i}=5\cdot\sqrt{p},i\in\{2,4,6\}$ and 00 otherwise. We set $\mathcal{J}$ to be the top 20 PCs for all designs, except MultiCauchy where we use the top 100 PCs. Penalty $h$ used in complement PCR step (described in Section 4.3.2) is identical to that used in Figure 1. See the corresponding QQ plot in Figure 6 from Appendix. bottom-left: Same setting as top-left except that the designs are taken from real datasets (real data descriptions are in Definition E.5 from Appendix). The dataset sizes are as follows. $\mathsf{Speech}$: $200\times 400$, $\mathsf{DNA}$: $100\times 180$, $\mathsf{SP500}$: $300\times 496$ and $\mathsf{FaceImage}$: $1348\times 2914$ and $\mathsf{Crime}$: $50\times 99$. All designs are centered and standardized (across rows) and then rescaled so that the average of eigenvalues of sample covariance matrix is $1$. The signal is generated in the same way as in the top-left. See the corresponding QQ plot in Figure 7 from Appendix. We set $\mathcal{J}$ to be the top 10 PCs for all designs, except FaceImage where we once again use the top 100 PCs. top-right: Under the setting of top-left, we plot the false coverage proportion (FCP) of the confidence intervals for $({\beta}^{\star}_{i})_{i=1}^{p}$, as we vary the targeted FCP level on the x-axis $\alpha$ from 00 to $1$. The $y$-axis also ranges from 0 to 1. bottom-right: analogous FCP plots under the setting of bottom-left.
4.3 PCR procedures
Given the index set $\mathcal{J}$, we describe PCR procedures that separately estimate the alignment component $\bm{\beta}^{\star}_{\mathsf{al}}$ and the complement component $\bm{\zeta}^{\star}$.
4.3.1 Classical PCR
The alignment component $\bm{\beta}^{\star}_{\mathsf{al}}$ can be readily recovered using the traditional PCR method. The method computes the following:
$\hat{\bm{\theta}}_{\mathsf{pcr}}(\mathcal{J}):=\left(\tilde{\mathbf{X}}_{\mathcal{J}}^{\top}\tilde{\mathbf{X}}_{\mathcal{J}}\right)^{-1}\tilde{\mathbf{X}}_{\mathcal{J}}^{\top}\mathbf{y}\in\mathbb{R}^{J},$ (31)
where $\tilde{\mathbf{X}}_{\mathcal{J}}:=\mathbf{X}\mathbf{O}_{\mathcal{J}}^{\top}\in\mathbb{R}^{n\times J}$ represents the basis-transformed design matrix and $\mathbf{O}_{\mathcal{J}}\in\mathbb{R}^{J\times p}$ comprises rows of $\mathbf{O}$ indexed by $\mathcal{J}$. The alignment PCR estimator is then given by $\hat{\bm{\beta}}_{\mathsf{al}}:=\mathbf{O}_{\mathcal{J}}^{\top}\hat{\bm{\theta}}_{\mathsf{pcr}}(\mathcal{J})\in\mathbb{R}^{p}$. Theorem 4.1 (a) shows that $\hat{\bm{\beta}}_{\mathsf{al}}$ is a consistent estimator of $\bm{\beta}^{\star}_{\mathsf{al}}$. This is the traditional PCR estimator, but it suffers from a shrinkage bias since it only recovers $\bm{\beta}^{\star}_{\mathsf{al}}$. To obtain an asymptotically unbiased estimator for $\bm{\beta}^{\star}$, it is essential to debias $\hat{\bm{\beta}}_{\mathsf{al}}$. We accomplish this in the following section.
4.3.2 Complement PCR
We leverage our Spectrum-Aware Debiasing theory to devise a modified PCR procedure that provides an accurate estimate of the complement component $\bm{\zeta}^{\star}$. We collect the indices in $\mathcal{J}^{c}$ corresponding to positive eigenvalues not used by Alignment PCR
$\bar{\mathcal{J}}:=\{i\in[p]:d_{i}^{2}>0,i\notin\mathcal{J}\}.$
Here, $|\bar{\mathcal{J}}|=N-J$ where $N=\mathrm{rank}(\mathbf{X})$. As a first step, we calculate a PCR estimator using the PCs indexed by $\bar{\mathcal{J}}$. That is, we calculate ${\hat{\bm{\theta}}_{\mathsf{pcr}}}^{\bot}=\hat{\bm{\theta}}_{\mathsf{pcr}}(\bar{\mathcal{J}})$ using the definition in (31). Next, we construct a new dataset as follows
$\mathbf{X}_{\mathsf{new}}:=\quantity(\mathbf{D}_{\bar{\mathcal{J}}}^{\top}\mathbf{D}_{\bar{\mathcal{J}}})^{1/2}\mathbf{O}_{\bar{\mathcal{J}}},\;\;\;\;\mathbf{y}_{\mathsf{new}}:=\quantity(\mathbf{D}_{\bar{\mathcal{J}}}^{\top}\mathbf{D}_{\bar{\mathcal{J}}})^{1/2}\hat{\bm{\theta}}_{\mathsf{pcr}}^{\bot},$ (32)
where $\mathbf{D}_{\bar{\mathcal{J}}}\in\mathbb{R}^{n\times(N-J)}$, $\mathbf{O}_{\bar{\mathcal{J}}}\in\mathbb{R}^{(N-J)\times p}$ respectively consist of the columns of $\mathbf{D}$ and the rows of $\mathbf{O}$ indexed by $\bar{\mathcal{J}}$. We employ Spectrum-Aware Debiasing on this new dataset. The resulting estimator, which we call complement PCR, is $\hat{\bm{\beta}}_{\mathsf{co}}=\hat{\bm{\beta}}^{u}(\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}},h)$, calculated using this new dataset in (3). We establish in Theorem 4.1 (b) that $\hat{\bm{\beta}}_{\mathsf{co}}$ is approximately Gaussian centered at the complement signal component $\bm{\zeta}^{\star}$, with variance $\hat{\tau}_{*}=\hat{\tau}_{*}(\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}},h)$ obtained using (26) on the new dataset.
4.3.3 PCR-Spectrum-Aware Debiasing or Debiased PCR
Combining our estimators from the previous sections, we obtain a debiased estimator for the full signal $\bm{\beta}^{\star}$ given by $\hat{\bm{\beta}}^{u}_{\mathsf{pcr}}:=\hat{\bm{\beta}}_{\mathsf{al}}+\hat{\bm{\beta}}_{\mathsf{co}}$. Since this estimator utilizes ideas from the classical PCR as well as our Spectrum-Aware Debiasing approaches, we name it PCR-Spectrum-Aware Debiasing. If the index set $\mathcal{J}$ includes all outlier PCs and PCs aligned with $\bm{\beta}^{\star}$, the procedure successfully removes shrinkage bias of the classical PCR estimator. It achieves this by “repurposing” discarded PCs to construct the complement component estimator $\hat{\bm{\beta}}_{\mathsf{co}}$. Algorithm 1 from Appendix presents the entire procedure in detail.
4.4 Asymptotic normality
We now state the asymptotic properties of the debiased PCR procedure. The proof of the theorem below is deferred to Appendix C in Appendix.
Theorem 4.1.
Suppose Assumptions 48 hold. Then, almost surely as $p\to\infty$, we have the following: (a) Alignment PCR: $\frac{1}{p}\norm{\hat{\bm{\beta}}_{\mathsf{al}}(\mathcal{J})-\bm{\beta}^{\star}_{\mathsf{al}}}^{2}\to 0$; (b) Complement PCR: $\hat{\tau}_{*}^{-1/2}\quantity(\hat{\bm{\beta}}_{\mathsf{co}}(\bar{\mathcal{J}})-\bm{\zeta}^{\star})\stackrel{{\scriptstyle W_{2}}}{{\to}}N(0,1)$; (c) Debiased PCR: $\hat{\tau}_{*}^{-1/2}\quantity(\hat{\bm{\beta}}^{u}_{\mathsf{pcr}}-\bm{\beta}^{\star})\stackrel{{\scriptstyle W_{2}}}{{\to}}N(0,1).$
Given exchangeability of entries of $\bm{\zeta}^{\star}$, we may obtain results analogous to Corollary 3.8 for finite or single coordinate inference. We defer the results to Section C.3 in Appendix.
We demonstrate Theorem 4.1 using two sets of design matrices. Our first set (top panel of Figure 3) represents more challenging variants of the settings from Figure 1. These designs contain high correlation, heterogeneity, or both. They also contain outlier eigenvalues and the signal $\bm{\beta}^{\star}$ aligns with a few top eigenvectors. Specifically, the top panel presents the following right-rotationally invariant designs: (i) $\mathsf{MatrixNormal}$-$\mathsf{B}$: stronger row- and column-wise correlations than $\mathsf{MatrixNormal}$; (ii) $\mathsf{Spiked}$-$\mathsf{B}$: larger and fewer spikes than $\mathsf{Spiked}$; (iii) $\mathsf{LNN}$-$\mathsf{B}$: matrix product with larger exponents and stronger correlations than $\mathsf{LNN}$; (iv) $\mathsf{VAR}$-$\mathsf{B}$: stronger row dependencies than $\mathsf{VAR}$; (v) $\mathsf{MultiCauchy}$: heavier tails than $\mathsf{Multi}$-$t$. Detailed description of these design distributions are given in Definition E.4 from Appendix. Our second set of experiments (bottom panel of Figure 3) uses real data designs from five domains: speech audio [42] , DNA [70], stock returns (S&P 500) [63], face images [48], and crime metrics [75]. FUrther details about these design matrices are included in Definition E.5 from Appendix.
4.5 Inference
Theorem 4.1 motivates an inference procedure similar to Section 4.5. For a specified level $\alpha\in[0,1]$, the confidence intervals
$\mathsf{CI}_{i}\quantity(\hat{\beta}^{u}_{\mathsf{pcr},i},\hat{\tau}_{*})=\left(\hat{\beta}^{u}_{\mathsf{pcr},i}+a\sqrt{\hat{\tau}_{*}},\hat{\beta}^{u}_{\mathsf{pcr},i}+b\sqrt{\hat{\tau}_{*}}\right)$ (33)
admit the false coverage proportion guarantee $\mathsf{FCP}(p):=\frac{1}{p}\sum_{i=1}^{p}\mathbb{I}\left({\beta}^{\star}_{i}\notin\mathsf{CI}_{i}\right)\rightarrow\alpha$, when $a,b$ satisfy $\Phi(b)-\Phi(a)=1-\alpha$. The right column of Figure 3 displays the FCP of these confidence intervals in the settings discussed following Theorem 4.1. PCR-Spectrum-Aware Debiasing achieves an FCP that aligns exceptionally well with the intended $\alpha$ values across these challenging settings, outperforming other methods.
4.6 Alignment testing
A fundamental challenge that modern data analysis presents relates to alignment of a part of the signal with eigenvectors of the sample covariance matrix. Such alignment distorts the performance of inference procedures unless they explicitly account for it. As a by-product, our Spectrum-Aware Debiasing theory provides a formal test for alignment, in other words, for testing $H_{i,0}^{\bm{\upsilon}^{\star}}:\upsilon^{\star}_{i}=0$ vs $H_{i,1}^{\bm{\upsilon}^{\star}}:\upsilon^{\star}_{i}\neq 0$, where $\upsilon^{\star}_{i}$ is given by (30). Below, $\bm{\upsilon}^{\star}$ refers to the vector with $i$-th entry $\upsilon^{\star}_{i}$. Corollary 4.3 below is proved in Section C.4 from Appendix.
Corollary 4.3.
Suppose that the assumptions in Theorem 4.1 hold. Then as $p\to\infty$,
$\hat{\bm{\Gamma}}^{-1/2}\quantity(\hat{\bm{\theta}}_{\mathsf{pcr}}-\bm{\upsilon}^{\star})\Rightarrow\mathcal{N}(\bm{0},\mathbf{I}_{J}),$
where $\hat{\bm{\theta}}_{\mathsf{pcr}}$ is given by (31), $\hat{\bm{\Gamma}}=\hat{\sigma}^{2}\cdot\quantity(\mathbf{D}^{\top}_{\mathcal{J}}\mathbf{D}_{\mathcal{J}})^{-1}+\hat{\omega}\cdot\mathbf{I}_{J}$ with $\mathbf{D}_{\mathcal{J}}\in\mathbb{R}^{n\times J}$ representing columns of $\mathbf{D}$ indexed by $\mathcal{J}$, $\hat{\sigma}^{2}$ a consistent estimator for the noise variance $\sigma^{2}$ given in (38) and $\hat{\omega}:=p^{-1}{\norm{\hat{\bm{\beta}}_{\mathsf{co}}}^{2}}-\hat{\tau}_{*}$.
Corollary 4.3 motivates the p-values $P_{i}:=2-2\cdot\Phi\quantity(\absolutevalue{\hat{\theta}_{\mathsf{pcr},i}/s_{i}}),s_{i}:=\sqrt{\hat{\sigma}^{2}\cdot d^{-2}_{\mathcal{J}(i)}+\hat{\omega}}$. Since the quantities $\quantity(P_{i})_{i=1}^{J}$ are asymptotically independent, the Benjamini-Hochberg procedure [13] can be used to control the False Discovery Rate (FDR), which is the expected ratio of PCs falsely identified as aligned with $\bm{\beta}^{\star}$ out of all PCs identified as aligned with $\bm{\beta}^{\star}$.
We demonstrate the efficacy of our alignment test in Figure 4. The top-left panel displays Benjamini-Hochberg adjusted p-values for testing alignment in the setting of the real-data designs considered in Figure 3, bottom-row. The bottom-left panel shows the true alignment angles between the underlying signal and the top six PCs. Our test accurately identifies alignment where present. In this setting, alignment detection is relatively easy, as the true alignment angles between the top PCs and signal, where present, are all significantly smaller than $90^{\circ}$ ($\angle(\bm{\beta}^{\star},\mathbf{o}_{i})\lesssim 75^{\circ}$). We illustrate in Figure 4, right-column how our alignment test performs when alignments become less pronounced and therefore harder to detect. Our method remains effective in rejecting all strong alignments present ($\angle(\bm{\beta}^{\star},\mathbf{o}_{i})\lesssim 75^{\circ}$). While it is less decisive in rejecting weak alignments ($75^{\circ}\lesssim\angle(\bm{\beta}^{\star},\mathbf{o}_{i})\lesssim 85^{\circ}$), the overall detected alignment pattern, as reflected in the small p-values shown in the top-right panel, closely matches the true alignment pattern displayed in the bottom-right panel [^3] Note that in the setting of Figure 4, we artificially aligned the signal with the 2nd, 4th and 6th PCs. However, the bottom row of Figure 4 suggests that the signal $\bm{\beta}^{\star}$ also aligns with the 1st PC for the SP500, FaceImage and Crime designs. This additional alignment was not introduced deliberately; however, it exists due to the following reason. Recall the signal decomposition from Equation 30 given by $\bm{\beta}^{\star}=\bm{\beta}^{\star}_{\mathsf{al}}+\bm{\zeta}^{\star}$. In our setting here, $\bm{\beta}^{\star}_{\mathsf{al}}$ is a linear combination of the 2nd, 4th, 6th PCs, while we generated $\bm{\zeta}^{\star}$ such that its entries have non-zero mean ($\bm{1}_{p}^{\top}\bm{\zeta}^{\star}\neq 0$). Coincidentally, the top PC of the SP500, FaceImage and Crime designs aligns with $\bm{1}_{p}$ due to correlation among the covariates. Thus, the intended alignment-complement decomposition of the signal is mis-specified, and the model is able to correctly identify the alignment of $\bm{\zeta}^{\star}$ with $\bm{1}_{p}$. We conducted similar experiments for the simulated designs in the top row of Figure 3 in Section E.4 from Appendix.
Figure 4: left-column: Same setting as Figure 3, bottom-left experiment. right-column: Same setting as Figure 3, bottom-left experiment, except that we increase difficulty of alignment detection by considering $v_{i}^{*}=\sqrt{p},i\in\{2,4,6\}$ as opposed to $v_{i}^{*}=5\cdot\sqrt{p},i\in\{2,4,6\}$. top-row: Benjamini-Hochberg adjusted p-values for $H_{i,0}^{\bm{\upsilon}^{\star}},i=1,...,6$ based on Corollary 4.3. ** indicates rejection under FDR level 0.05 and * rejection under FDR level 0.1. bottom-row: True alignment angles between the signal $\bm{\beta}^{\star}$ and $i$-th PC $\mathbf{o}_{i}$ calculated using $\mathrm{argcos}\{\mathbf{o}_{i}^{\top}\bm{\beta}^{\star}/(\norm{\mathbf{o}_{i}}_{2}.\norm{\bm{\beta}^{\star}}_{2})\}$.
5 Proof outline and novelties
Our main result relies on three main steps: (i) a characterization of the empirical distribution of a population version of $\hat{\bm{\beta}}$, (ii) connecting this population version with our data-driven Spectrum-Aware estimator, (iii) developing a consistent estimator of the asymptotic variance. We next describe our main technical novelties for step (i) in Section 5.1, and that for steps (ii) and (iii) in Section 5.2.
5.1 Result A: Distributional characterizations
Theorem 3.1 relies on the characterization of certain properties of $\hat{\bm{\beta}}$ and the following two quantities:
$\mathbf{r}_{*}:=\hat{\bm{\beta}}+\frac{1}{\gamma_{*}}\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}}),\quad\mathbf{r}_{**}:=\hat{\bm{\beta}}+\frac{1}{\eta_{*}-\gamma_{*}}\mathbf{X}^{\top}(\mathbf{X}\hat{\bm{\beta}}-\mathbf{y}).$ (34)
Here, $\mathbf{r}_{*}$ can be interpreted as the population version of the debiased estimator $\hat{\bm{\beta}}^{u}$ and $\mathbf{r}_{**}$ as an auxiliary quantity that arises in the intermediate steps in our proof. The following theorem characterizes the empirical distribution of the entries of $\hat{\bm{\beta}}$ and $\mathbf{r}_{*}$. We prove it in Section B.1 from Appendix.
Theorem 5.1 (Distributional characterizations).
Under Assumptions 26, almost surely as $n,p\to\infty$,
$\left(\hat{\bm{\beta}},\mathbf{r}_{*},\bm{\beta}^{\star}\right)\stackrel{{\scriptstyle W_{2}}}{{\rightarrow}}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right),\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star},\mathsf{B}^{\star}\right),$ (35)
where $\mathsf{Z}\sim N(0,1)$ is independent of $\mathsf{B}^{\star}$. Furthermore, almost surely as $p\to\infty$
$\frac{1}{p}\left\|\mathbf{X}\mathbf{r}_{**}-\mathbf{y}\right\|^{2}\to\tau_{**}\cdot\mathbb{E}\mathsf{D}^{2}+\sigma^{2}\cdot\delta,$ (36)
$\frac{1}{p}\norm{\mathbf{y}-\mathbf{X}\hat{\bm{\beta}}}^{2}\to\tau_{**}\cdot\mathbb{E}\frac{\mathsf{D}^{2}(\eta_{*}-\gamma_{*})^{2}}{(\mathsf{D}^{2}+\eta_{*}-\gamma_{*})^{2}}+\sigma^{2}\cdot\quantity(\frac{n-p}{p}+\mathbb{E}\quantity(\frac{\eta_{*}-\gamma_{*}}{\mathsf{D}^{2}+\eta_{*}-\gamma_{*}})^{2})$
5.2 Result B: Consistent estimation of fixed points
Note that the population debiased estimator $\mathbf{r}_{*}$ cannot be used to conduct inference since $\gamma_{*}$ is unknown. Furthermore, the previous theorem says roughly that $\mathbf{r}_{*}-\bm{\beta}^{\star}$ behaves as a standard Gaussian with variance $\tau_{*}$, without providing any estimator for $\tau_{*}$. We address these two points here. In particular, we will see that addressing these points ties us to establishing consistent estimators for the solution to the fixed points defined in (20). The theorem below shows that $(\widehat{\mathsf{adj}},\hat{\eta}_{*},\hat{\tau}_{*},\hat{\tau}_{**})$ from (26) serve as consistent estimators of the fixed points $(\gamma_{*},\eta_{*},\tau_{*},\tau_{**})$, and $\hat{\bm{\beta}}^{u},\hat{\mathbf{{r}}}_{**}$ as consistent estimators of $\mathbf{r}_{*}$ and $\mathbf{r}_{**}$, where $\hat{\mathbf{{r}}}_{**}$ is defined as in (37) below. For the purpose of the discussion below, we note that $\hat{\tau}_{**}$ from (26) can be written as follows.
$\hat{\tau}_{**}(p):=\frac{\frac{1}{p}\left\|\mathbf{X}\hat{\mathbf{{r}}}_{**}-\mathbf{y}\right\|^{2}-\frac{n}{p}\cdot\sigma^{2}}{\frac{1}{p}\sum_{i=1}^{p}d_{i}^{2}};\quad\hat{\mathbf{{r}}}_{**}:=\hat{\bm{\beta}}+\frac{1}{\hat{\eta}_{*}-\widehat{\mathsf{adj}}}\mathbf{X}^{\top}(\mathbf{X}\hat{\bm{\beta}}-\mathbf{y}).$ (37)
Furthermore, recall that when the noise level $\sigma^{2}$ is unknown, one requires an estimator for $\sigma^{2}$ to calculate $\hat{\tau}_{*},\hat{\tau}_{**}$ in (26). We define such an estimator below and show that that it estimates $\sigma^{2}$ consistently.
$\hat{\sigma}^{2}(\mathbf{X},\mathbf{y},h)\leftarrow\frac{\|\mathbf{y}-\mathbf{X}\widehat{\bm{\beta}}\|^{2}-\frac{\left\|\left(\mathbf{I}_{n}+\frac{1}{\hat{\eta}_{*}-\mathrm{adj}}\mathbf{XX}^{\top}\right)(\mathbf{y}-\mathbf{X}\widehat{\bm{\beta}})\right\|^{2}}{\sum_{i=1}^{p}d_{i}^{2}}\sum_{i=1}^{p}\frac{\left(\hat{\eta}_{*}-\widehat{\mathsf{adj}}\right)^{2}d_{i}^{2}}{\left(d_{i}^{2}+\hat{\eta}_{*}-\widehat{\mathsf{adj}}\right)^{2}}}{\sum_{i=1}^{p}\frac{\left(\hat{\eta}_{*}-\widehat{\mathsf{adj}}\right)^{2}\cdot\left(\sum_{j=1}^{p}d_{j}^{2}-nd_{i}^{2}\right)}{\left(d_{i}^{2}+\hat{\eta}_{*}-\widehat{\mathsf{adj}}\right)^{2}\cdot\left(\sum_{j=1}^{p}d_{j}^{2}\right)}+n-p}.$ (38)
Note this is well-defined when
$\frac{n}{p}\cdot\frac{\frac{1}{p}\sum_{i=1}^{p}d_{i}^{2}\cdot\quantity(1-\quantity(\frac{\hat{\eta}_{*}-\widehat{\mathsf{adj}}}{d_{i}^{2}+\hat{\eta}_{*}-\widehat{\mathsf{adj}}})^{2})}{\frac{1}{p}\sum_{i=1}^{p}d_{i}^{2}\cdot\frac{1}{p}\sum_{i=1}^{p}\quantity(1-\quantity(\frac{\hat{\eta}_{*}-\widehat{\mathsf{adj}}}{d_{i}^{2}+\hat{\eta}_{*}-\widehat{\mathsf{adj}}})^{2})}\neq 1.$ (39)
In particular, the LHS of (39) consistently estimates the LHS of (21) in Assumption 7.
Theorem 5.2 (Consistent estimation of fixed points).
Suppose that Assumption 27 hold. Then, the estimators in (26) and (37) are well-defined for any $p$ and we have that almost surely as $p\to\infty$,
$\widehat{\mathsf{adj}}\left(p\right)\rightarrow\gamma_{*},\quad\hat{\eta}_{*}\left(p\right)\rightarrow\eta_{*},\quad\hat{\tau}_{*}\left(p\right)\rightarrow\tau_{*},\quad\hat{\tau}_{**}\left(p\right)\rightarrow\tau_{**},\quad\hat{\sigma}^{2}\left(p\right)\rightarrow\sigma^{2}$
$\frac{1}{p}\norm{\hat{\bm{\beta}}^{u}(p)-\mathbf{r}_{*}}^{2}\rightarrow 0,\quad\frac{1}{p}\norm{\hat{\mathbf{{r}}}_{**}(p)-\mathbf{r}_{**}}^{2}\rightarrow 0.$
We prove this theorem in Section B.2 from Appendix. It is not hard to see that Theorem 5.1 combined with Theorem 5.2 proves our main result Theorem 3.1.
5.3 Proof of result A
In this section, we discuss our proof novelties for Theorem 5.1. Section B.1 from Appendix contains this proof.
We base our proof on the approximate message passing (AMP) machinery (cf. [24, 104, 81, 36, 68] for a non-exhaustive list of references). In this approach, one constructs an AMP algorithm in terms of fixed points ($\eta_{*},\gamma_{*},\tau_{*},\tau_{**}$ in our case) and shows that its iterates $\hat{\mathbf{v}}^{t}$ converge to our objects of interest $\hat{\mathbf{v}}$ ($\hat{\mathbf{v}}$ can be $\hat{\bm{\beta}}$ or $\mathbf{r}_{*}$ in our case) in the following sense: almost surely
$\lim_{t\rightarrow\infty}\lim_{p\to\infty}\frac{\left\|\hat{\mathbf{v}}^{t}-\hat{\mathbf{v}}\right\|^{2}}{p}=0.$ (40)
AMP theory provides a precise characterization of the following limit involving the algorithmic iterates for any fixed $t$: $\lim_{p\rightarrow\infty}\|\hat{\mathbf{v}}^{t}-\mathbf{v}_{0}\|^{2}/p,$ where $\mathbf{v}_{0}$ is usually a suitable function of $\bm{\beta}_{\star}$ around which one expects $\hat{\mathbf{v}}$ should be centered. Thus plugging this in (40) yields properties of the object of interest $\hat{\mathbf{v}}$. Within this theory, the framework that characterizes $\lim_{p\rightarrow\infty}\|\hat{\mathbf{v}}^{t}-\mathbf{v}_{0}\|^{2}/p$ is known as state evolution [7, 52]. Despite the existence of this solid machinery, (40) requires a case-by-case proof, and for many settings, this presents deep challenges.
We use the above algorithmic proof strategy, but in case of our right-rotationally invariant designs that, the original AMP algorithms fail to apply. To alleviate this, [74] proposed vector approximate message passing algorithms. We use these algorithms to create our $\hat{\mathbf{v}}^{t}$’s. Subsequently, proving (40) presents the main challenge. To this end, one is required to show the following Cauchy convergence property of the VAMP iterates: almost surely, $\lim_{(s,t)\rightarrow\infty}\left(\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\hat{\mathbf{v}}^{t}-\hat{\mathbf{v}}^{s}\right\|^{2}\right)=0.$ We prove this using a Banach contraction argument (cf. (81) from Appendix). Such an argument saw prior usage in the context of Bayes optimal learning in [59]. However, they studied a “matched” problem where the signal prior (analogous to $\mathsf{B}^{\star}$ in our setting) is known to the statistician and she uses this exact prior during the estimation process. Arguments under such matched Bayes optimal problems do not translate to our case, and proving (81) presents novel difficulties in our setting. To mitigate this, we leverage a fundamental property of the R-transform, specifically that $-zR^{\prime}(z)/R(z)<1$ for all $z$, and discover and utilize a crucial interplay of this property with the non-expansiveness of the proximal map (see Proposition A.7 (b) from Appendix).
Remark 5.1 (Comparison with [40, 41]).
In their seminal works, [40, 41] initiated the first study of the risk of regularized estimators under right-rotationally invariant designs. They stated a version of Theorem 5.1 with a partially non-rigorous argument. In their approach, an auxiliary $\ell_{2}$ penalty of sufficient magnitude is introduced to ensure contraction of AMP iterates. Later, they remove this penalty through an analytical continuation argument. However, this proof suffers two limitations. The first one relates to the non-rigorous applications of the AMP state evolution results. For instance, [41, Lemma 3] shows that for each fixed value of $p$, $\lim_{t\rightarrow\infty}\frac{\left\|\hat{x}^{t}-\hat{x}\right\|^{2}}{p}=0.$ However, in [41, Proof of Lemma 4], the authors claim that this would imply (40) upon exchanging limits with respect to $t$ and $p$. Such an exchange of limits is non-rigorous since the correctness of AMP state evolution is established for a finite number of iterations ($t<T,T$ fixed) as $p\to\infty$. The limit in $T$ is taken after $p$. The other limitation lies in the analytic continuation approach that requires multiple exchanges of limit operations [41, Appendix H] that seem difficult to justify and incur intractable assumptions [41, Assumption 1 (c), (e)] (in particular, it is unclear how to verify the existence claim in Assumption 1 (c) beyond Gaussian designs). Our alternative approach establishes contraction without the need for a sufficiently large $\ell_{2}$-regularization component, as in [40, 41], and thereby avoids the challenges associated with the analytic continuation argument.
5.4 Proof of result B
In this section, we discuss the proof of Theorem 5.2. See Section B.2 from Appendix for the proof details.
First, let us present some heuristics for how one might derive the consistent estimators $\quantity(\widehat{\mathsf{adj}},\hat{\eta}_{*},\hat{\tau}_{*},\hat{\tau}_{**})$. We start from (20a). Using Lemma 2.7, it can be written as
$\frac{\gamma_{*}}{\eta_{*}}=\mathbb{E}\frac{1}{1+\gamma_{*}^{-1}h^{\prime\prime}\quantity(\operatorname{Prox}_{\gamma_{*}^{-1}h}(\mathsf{B}^{\star}+\sqrt{\tau_{*}}\mathsf{Z}))}.$ (41)
Recall that we have established Theorem 5.1 that shows $p\to\infty$, almost surely, $\hat{\bm{\beta}}\stackrel{{\scriptstyle W_{2}}}{{\to}}\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right).$ Combining this and (41), we expect that
$\frac{1}{\eta_{*}}\approx\frac{1}{p}\sum_{i=1}^{p}\frac{\gamma_{*}^{-1}}{1+\gamma_{*}^{-1}h^{\prime\prime}(\hat{{\beta}}_{i})}.$ (42)
Using the definition of R-transform, we can rewrite (20c) as $\eta_{*}^{-1}=\mathbb{E}\frac{1}{\mathsf{D}^{2}+\eta_{*}-\gamma_{*}}$ which, along with (12), implies that $\frac{1}{\eta_{*}}\approx\frac{1}{p}\sum_{i=1}^{p}\frac{1}{d_{i}^{2}+\eta_{*}-\gamma_{*}}.$ Combining this and (42) to eliminate $\eta_{*}$, we obtain that
$\frac{1}{p}\sum_{i=1}^{p}\frac{1}{\left(d_{i}^{2}-\gamma_{*}\right)\left(\frac{1}{p}\sum_{j=1}^{p}\quantity(\gamma_{*}+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right))^{-1}\right)+1}\approx 1.$ (43)
Setting $\approx$ above to equality, we obtain our exact equation for the Spectrum-Aware adjustment factor, i.e.  (23). One thus expects intuitively that $\widehat{\mathsf{adj}}$ consistently estimates $\gamma_{*}$. To establish the consistency rigorously, we recognize and establish the monotonicity of the LHS of (43) as a function of $\gamma_{*}$, and study its point-wise limit. We direct the reader to Lemma B.11 and Proposition B.13 from Appendix for more details.
Once we have established the consistency of $\widehat{\mathsf{adj}}$ as an estimator for $\gamma_{*}$, we substitute $\widehat{\mathsf{adj}}$ back into (42) to obtain a consistent estimator $\hat{\eta}_{*}$ for $\eta_{*}$. It is important to note that the definition of $\mathbf{r}_{**}$, as given in (34), only involves the fixed points $\eta_{*}$ and $\gamma_{*}$. As a result, we can utilize $\widehat{\mathsf{adj}}$ and $\hat{\eta}_{*}$ to produce a consistent estimator $\hat{\mathbf{{r}}}_{**}$ for $\mathbf{r}_{**}$. Now note that (36) would give us a system of linear equation
$\matrixquantity(\frac{1}{p}\left\|\mathbf{X}\mathbf{r}_{**}-\mathbf{y}\right\|^{2}\\ \frac{1}{p}\norm{\mathbf{y}-\mathbf{X}\hat{\mathbf{\beta}}}^{2})\approx\matrixquantity(\mathbb{E}\mathsf{D}^{2}&\delta\\ \mathbb{E}\frac{\mathsf{D}^{2}(\eta_{*}-\gamma_{*})^{2}}{(\mathsf{D}^{2}+\eta_{*}-\gamma_{*})^{2}}&\quad\frac{n-p}{p}+\mathbb{E}\quantity(\frac{\eta_{*}-\gamma_{*}}{\mathsf{D}^{2}+\eta_{*}-\gamma_{*}})^{2})\matrixquantity(\tau_{**}\\ \sigma^{2})$ (44)
The estimators $(\hat{\tau}_{**},\hat{\sigma}^{2})$ in (26) for $(\tau_{**},\sigma^{2})$ are solved from the two linear equations above with the 2-by-2 matrix on RHS replaced by its sample version. Note that (21) is required to ensure the 2-by-2 matrix is non-singular. Now with estimators for $\gamma_{*},\eta_{*},\sigma^{2}$ and $\tau_{**}$, we can construct the estimator $\hat{\tau}_{*}$ for $\tau_{*}$ using (20d) and (12).
6 Discussion
We conclude our paper with a discussion of two main points. First, we clarify that although our method can handle various dependencies in the design through the right-rotational invariance assumption, it does not extend to anisotropic Gaussian designs where the rows of $\mathbf{X}$ are sampled from $\mathcal{N}(\mathbf{0},\mathbf{\Sigma})$ with an arbitrary covariance matrix $\mathbf{\Sigma}$ (unless $\mathbf{\Sigma}$ is right-rotationally invariant). Moreover, contrasting with [12], our Spectrum-Aware adjustment (4) does not apply directly to non-separable penalties, e.g. SLOPE, group Lasso, etc. We note that the current framework can be expanded to address both these issues. In Appendix D from Appendix, we suggest a debiased estimator for “ellipsoidal designs” $\mathbf{X}=\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}\mathbf{\Sigma}^{1/2}$ and non-separable convex penalties. We also conjecture its asymptotic normality using the non-separable VAMP formalism [37]. We leave a detailed study of this extensive class of estimators to future works.
We discuss another potential direction of extension, that of relaxing the exchangeability assumption in Corollary 3.8 and Corollary C.1 from Appendix that establish inference guarantees on finite-dimensional marginals. One may raise a related question, that of constructing confidence intervals for $\mathbf{a}^{\top}\bm{\beta}^{\star}$ for a given choice of $\mathbf{a}$. Under Gaussian design assumptions, such guarantees were obtained using the leave-one-out method as in [18, Section 4.6] or Stein’s method as in [12] without requiring the exchangeability assumption (at the cost of other assumptions on $\bm{\beta}^{\star}$ and/or $\bm{\Sigma}$). Unfortunately, these arguments no longer apply under right-rotational invariant designs owing to the presence of a global dependence structure. Thus, establishing such guarantees without exchangeability can serve as an exciting direction for future research.
{acks}
[Acknowledgments] P.S. was funded partially by NSF DMS-2113426. The authors would like to thank Florent Krzakala and Cedric Gerbelot for clarification on the contributions in [40, 41], and Boris Hanin for references on linear neural networks.
Appendix A Preliminary
A.1 Empirical Wasserstein-2 convergence
We will use below the following fact. See [32, Appendix E] and references within for its justification.
Proposition A.1.
To verify $(\mathbf{v}_{1},\ldots,\mathbf{v}_{k})\stackrel{{\scriptstyle W_{2}}}{{\rightarrow}}(\mathsf{V}_{1},\ldots,\mathsf{V}_{k})$, it suffices to check that
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}f\left(v_{i,1},\ldots,v_{i,k}\right)=\mathbb{E}\left[f\left(\mathsf{V}_{1},\ldots,\mathsf{V}_{k}\right)\right]$
holds for every function $f:\mathbb{R}^{k}\rightarrow\mathbb{R}$ satisfying, for some constant $C>0$, the pseudo-Lipschitz condition $\left|f(\mathbf{v})-f\left(\mathbf{v}^{\prime}\right)\right|\leq C\left(1+\|\mathbf{v}\|_{2}+\left\|\mathbf{v}^{\prime}\right\|_{2}\right)\left\|\mathbf{v}-\mathbf{v}^{\prime}\right\|_{2}.$ Meanwhile, this condition implies (11).
The following results are from [32, Appendix E].
Proposition A.2.
Suppose $\mathbf{V}\in\mathbb{R}^{n\times t}$ has i.i.d. rows equal in law to $\mathsf{V}\in\mathbb{R}^{t}$, which has finite mixed moments of all orders. Then $\mathbf{V}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{V}$ almost surely as $n\to\infty$. Furthermore, if $\mathbf{E}\in\mathbb{R}^{n\times k}$ is deterministic with $\mathbf{E}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{E}$, then $(\mathbf{V},\mathbf{E})\stackrel{{\scriptstyle W_{2}}}{{\to}}(\mathsf{V},\mathsf{E})$ almost surely where $\mathsf{V}$ is independent of $\mathsf{E}$.
Proposition A.3.
Suppose $\mathbf{V}\in\mathbb{R}^{n\times k}$ satisfies $\mathbf{V}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{V}$ as $n\to\infty$, and $g:\mathbb{R}^{k}\to\mathbb{R}^{l}$ is continuous with $\|g(\mathbf{v})\|\leq C(1+\|\mathbf{v}\|)^{\mathfrak{p}}$ for some $C>0$ and $\mathfrak{p}\geq 1$. Then $g(\mathbf{V})\stackrel{{\scriptstyle W_{2}}}{{\to}}g(\mathsf{V})$ where $g(\cdot)$ is applied row-wise to $\mathbf{V}$.
Proposition A.4.
Suppose $\mathbf{V}\in\mathbb{R}^{n\times k}$, $\mathbf{W}\in\mathbb{R}^{n\times l}$, and $\mathbf{M}_{n},\mathbf{M}\in\mathbb{R}^{k\times l}$ satisfy $\mathbf{V}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{V}$, $\mathbf{W}\stackrel{{\scriptstyle W_{2}}}{{\to}}0$, and $\mathbf{M}_{n}\to\mathbf{M}$ entrywise as $n\to\infty$. Then $\mathbf{V}\mathbf{M}_{n}+\mathbf{W}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{V}^{\top}\cdot\mathbf{M}$.
Proposition A.5.
Fix $\mathfrak{p}\geq 1$ and $k\geq 0$. Suppose $\mathbf{V}\in\mathbb{R}^{n\times k}$ satisfies ${\mathbf{V}}\stackrel{{\scriptstyle W_{2}}}{{\rightarrow}}\mathsf{V}$, and $f:\mathbb{R}^{k}\rightarrow\mathbb{R}$ is a function satisfying (11) that is continuous everywhere except on a set having probability 0 under the law of $\mathsf{V}$. Then $\frac{1}{n}\sum_{i=1}^{n}f({\mathbf{V}})_{i}\rightarrow\mathbb{E}[f(\mathsf{V})].$
Proposition A.6.
Fix $l\geq 0$, let $\mathbf{O}\sim\operatorname{Haar}(\mathbb{O}(n-l))$, and let $\mathbf{v}\in\mathbb{R}^{n-l}$ and $\bm{\Pi}\in\mathbb{R}^{n\times(n-l)}$ be deterministic, where $\bm{\Pi}$ has orthonormal columns and $n^{-1}\|\mathbf{v}\|^{2}\to\sigma^{2}$ as $n\to\infty$. Then $\bm{\Pi}\mathbf{O}\mathbf{v}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{Z}\sim N(0,\sigma^{2})$ almost surely. Furthermore, if $\mathbf{E}\in\mathbb{R}^{n\times k}$ is deterministic with $\mathbf{E}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{E}$, then $(\bm{\Pi}\mathbf{O}\mathbf{v},\mathbf{E})\stackrel{{\scriptstyle W_{2}}}{{\to}}(\mathsf{Z},\mathsf{E})$ almost surely where $\mathsf{Z}$ is independent of $\mathsf{E}$.
A.2 Proximal map
We collect a few useful properties of proximal map.
Proposition A.7.
Under Assumption 4, we have that for any $v>0$,
  • (a)
    for any $x,y\in\mathbb{R}$, $y=\operatorname{Prox}_{vh}(x)\iff x-y\in v\partial h(y)$ where $\partial h$ is the subdifferential of $h$;
  • (b)
    Proximal map is firmly non-expansive: for any $x,y\in\mathbb{R}$, $\quantity|\operatorname{Prox}_{vh}(x)-\operatorname{Prox}_{vh}(y)|^{2}\leq(x-y)(\operatorname{Prox}_{vh}(x)-\operatorname{Prox}_{vh}(y))$. This implies that $x\mapsto\operatorname{Prox}_{vh}(x)$ is 1-Lipschitz continuous.
Proof of Lemma 2.7.
Under Assumption 4, for any $v>0$, $x\mapsto\operatorname{Prox}_{vh}(x)$ is continuous, monotone increasing in $x$, and continuously differentiable at any $x$ such that $\operatorname{Prox}_{vh}(x)\notin\mathfrak{D}$ and
$\operatorname{Prox}_{vh}^{\prime}(x)=\frac{1}{1+vh^{\prime\prime}\left(\operatorname{Prox}_{vh}(x)\right)}.$ (45)
This follows from the assumption that $h(x)$ is twice continuously differentiable on $\mathfrak{D}^{c}$ and the implicit differentiation calculation shown in [40, Appendix B1]. For $x\in\{x:\operatorname{Prox}_{vh}(x)\in\mathfrak{D}\}$, $\operatorname{Prox}_{vh}(x)$ is differentiable and has derivative equal to 0 except for a finite set of points. To see this, note that preimage $\operatorname{Prox}^{-1}_{vh}(\mathbf{y})$ for $y\in\mathfrak{D}$ is either a singleton set or a closed interval of the form $[x_{1},x_{2}]$ for $x_{1}\in\mathbb{R}\cup\{-\infty\},x_{2}\in\mathbb{R}\cup\{+\infty\}$ and $x_{1}<x_{2}$, using continuity and monotonicity of $x\mapsto\operatorname{Prox}_{vh}(x)$. This implies that $\{x:\operatorname{Prox}_{vh}(x)\in\mathfrak{D}\}$ is a union of finite number of singleton sets and a finite number of closed intervals. Furthermore, $\operatorname{Prox}_{vh}(x)$ is constant on each of the closed intervals. It follows that $\operatorname{Prox}_{vh}(x)$ is differentiable and has derivative equal to 00 on the interiors of the closed intervals, and that $\mathcal{C}$ is union of some of the singleton sets and all of the finite-valued endpoints of the closed intervals.
We extend functions $h^{\prime\prime}(x)$ and $\operatorname{Prox}_{vh}^{\prime}(x)$ on $\mathfrak{D}$ and $\mathcal{C}$ respectively in the following way: (i) For $y_{0}\in\mathfrak{D}$ such that $\operatorname{Prox}^{-1}_{vh}(y_{0})$ is a closed interval with endpoints $x_{1}\in\mathbb{R}\cup\{-\infty\},x_{2}\in\mathbb{R}\cup\{+\infty\}$ and $x_{1}<x_{2}$, we set $h^{\prime\prime}(y_{0})\leftarrow+\infty$ and $\operatorname{Prox}^{\prime}_{vh}(x)\leftarrow 0$ for all $x\in[x_{1},x_{2}]$ (ii) For $y_{0}\in\mathfrak{D}$ such that $\operatorname{Prox}^{-1}_{vh}(y_{0})$ is a singleton set and its sole element $x_{0}$ is contained in $\mathcal{C}$, we set $h^{\prime\prime}(y_{0})\leftarrow+\infty,\operatorname{Prox}_{vh}^{\prime}(x_{0})\leftarrow 0$; (iii) For $y_{0}\in\mathfrak{D}$ such that $\operatorname{Prox}^{-1}_{vh}(y_{0})$ is a singleton set $\{x_{0}\}$ and that $x\mapsto\operatorname{Prox}_{vh}(x)$ is differentiable at $x_{0}$ with 0 derivative, we set $h(y_{0})\leftarrow+\infty$.
We show that it is impossible to have some $y_{0}\in\mathfrak{D}$ such that $\operatorname{Prox}_{vh}^{-1}\left(y_{0}\right)$ is a singleton set $\left\{x_{0}\right\}$ and that $x\mapsto\operatorname{Prox}_{vh}(x)$ is differentiable at $x_{0}$ with non-zero derivative. This means that all $\mathbf{y}\in\mathfrak{D}$ belongs to cases (i), (ii) and (iii) above. Suppose to the contrary. We know from the above discussion that there exists some $\mathfrak{e}>0$ such that $\operatorname{Prox}_{vh}^{\prime}(x)$ is continuous on $\left(x_{0},x_{0}+\mathfrak{e}\right)$ and $\left(x_{0}-\mathfrak{e},x_{0}\right)$. We claim that $x\mapsto\operatorname{Prox}_{vh}^{\prime}(x)$ is continuous at $x_{0}$. To see this, note that for any $\Delta>0$, we can find $\bm{\varepsilon}\in(0,\mathfrak{e})$ such that
  • there exists some $x_{+}\in\left(x_{0},x_{0}+\epsilon\right)$ such that for any $x\in\left(x_{0},x_{0}+\epsilon\right)$,
    $\left|\operatorname{Prox}_{vh}^{\prime}(x)-\operatorname{Prox}_{vh}^{\prime}\left(x_{+}\right)\right|<\frac{\Delta}{5},\quad\left|\frac{\operatorname{Prox}_{vh}\left(x_{0}\right)-\operatorname{Prox}_{vh}\left(x_{+}\right)}{x_{0}-x_{+}}-\operatorname{Prox}_{vh}^{\prime}\left(x_{+}\right)\right|<\frac{\Delta}{5}$
  • there exists some $x_{-}\in\left(x_{0}-\epsilon,x_{0}\right)$ such that for any $x\in\left(x_{0}-\epsilon,x_{0}\right)$,
    $\left|\operatorname{Prox}_{vh}^{\prime}(x)-\operatorname{Prox}_{vh}^{\prime}\left(x_{-}\right)\right|<\frac{\Delta}{5},\quad\left|\frac{\operatorname{Prox}_{vh}\left(x_{0}\right)-\operatorname{Prox}_{vh}\left(x_{-}\right)}{x_{0}-x_{-}}-\operatorname{Prox}_{vh}^{\prime}\left(x_{-}\right)\right|<\frac{\Delta}{5}$
  • for any $x\in\left(x_{0}-\epsilon,x_{0}\right)\cup\left(x_{0},x_{0}+\epsilon\right)$,
    $\left|\operatorname{Prox}_{vh}^{\prime}\left(x_{0}\right)-\frac{\operatorname{Prox}_{vh}\left(x_{0}\right)-\operatorname{Prox}_{vh}(x)}{x_{0}-x}\right|<\frac{\Delta}{5}.$
Then for any $x\in\left(x_{0}-\epsilon,x_{0}+\epsilon\right)$, we have $\left|\operatorname{Prox}_{vh}^{\prime}\left(x_{0}\right)-\operatorname{Prox}_{vh}^{\prime}(x)\right|<\Delta$ by triangle inequality. This proves the claim. Now, since $x\mapsto\operatorname{Prox}_{vh}(x)$ is continuously differentiable on $\left(x_{0}-\mathfrak{e},x_{0}+\mathfrak{e}\right)$ and $\operatorname{Prox}_{vh}^{\prime}\left(x_{0}\right)\neq 0$, inverse function theorem implies that $y\mapsto\operatorname{Prox}_{vh}^{-1}(y)$ is a well defined, real-valued function and it is continuous differentiable on some open interval $U$ containing $y_{0}$. This implies that $h$ is differentiable at any $y\in U$ and that $y\mapsto\operatorname{Prox}_{vh}^{-1}(y)=y+vh^{\prime}(y)$ is continuously differentiable. But this would imply that $h$ is twice continuously differentiable on $U$ which contradicts the assumption that $y_{0}\in\mathfrak{D}$.
Note that we have assigned +$\infty$+ ∞ to $h^{\prime\prime}$ on $\mathfrak{D}$ and 00 to $\operatorname{Prox}^{\prime}_{vh}$ on $\mathcal{C}$. Piecewise continuity of $x\mapsto\frac{1}{w+h^{\prime\prime}\left(\operatorname{Prox}_{vh}(x)\right)}$ for any $w>0$ follows from the discussion above. ∎
A.3 Properties of R- and Cauchy transform
The following shows that the Cauchy- and R-transforms of $-\mathsf{D}^{2}$ are well-defined by (14), and reviews their properties.
Lemma A.8.
Let $G(\cdot)$ and $R(\cdot)$ be the Cauchy- and R-transforms of $-\mathsf{D}^{2}$ under Assumption 2.
  • (a)
    The function $G:(-d_{-},\infty)\to\mathbb{R}$ is positive and strictly decreasing. Setting $G\left(-d_{-}\right):=\lim_{z\to-d_{-}}G(z)\in(0,\infty]$, $G$ admits a functional inverse $G^{-1}:(0,G(-d_{-}))\to(-d_{-},\infty)$.
  • (b)
    The function $R:\left(0,G\left(-d_{-}\right)\right)\to\mathbb{R}$ is negative and strictly increasing.
  • (c)
    For any $z\in\left(0,G\left(-d_{-}\right)\right),R^{\prime}(z)=-\left(\mathbb{E}\frac{1}{\left(\mathsf{D}^{2}+R(z)+\frac{1}{z}\right)^{2}}\right)^{-1}+\frac{1}{z^{2}}$.
  • (d)
    For any $z\in\left(0,G\left(-d_{-}\right)\right),-\frac{zR^{\prime}(z)}{R(z)}\in(0,1)$.
  • (e)
    For any $z\in\left(0,G\left(-d_{-}\right)\right),z^{2}R^{\prime}(z)\in(0,1)$.
  • (f)
    For all sufficiently small $z\in(0,G(-d_{-}))$, R-transform admits convergent series expansion given by
    $R(z)=\sum_{k\geq 1}\kappa_{k}z^{k-1}$ (46)
    where $\left\{\kappa_{k}\right\}_{k\geq 1}$ are the free cumulants of the law of $-\mathsf{D}^{2}$ and $\kappa_{1}=-\mathbb{E}\mathsf{D}^{2}$ and $\kappa_{2}=\mathbb{V}(\mathsf{D}^{2})$.
Proof.
See [59, Lemma G.6] for (a) and (b), To see (c), for any $z\in\left(0,G\left(-d_{-}\right)\right)$, differentiating $R(z)=G^{-1}(z)-z^{-1}$ yields
$-zR^{\prime}(z)=z\left(\mathbb{E}\frac{1}{\left(\mathsf{D}^{2}+G^{-1}(z)\right)^{2}}\right)^{-1}-\frac{1}{z}$
To see (d),
$-\frac{zR^{\prime}(z)}{R(z)}=\frac{z\left(\mathbb{E}\frac{1}{\left(\mathsf{D}^{2}+G^{-1}(z)\right)^{2}}\right)^{-1}-\frac{1}{z}}{G^{-1}(z)-\frac{1}{z}}<1$
$\Leftrightarrow z\left(\mathbb{E}\frac{1}{\left(\mathsf{D}^{2}+G^{-1}(z)\right)^{2}}\right)>G^{-1}(z)$
$\Leftrightarrow\mathbb{E}\frac{G^{-1}(z)}{\left(\mathsf{D}^{2}+G^{-1}(z)\right)^{2}}<z=\mathbb{E}\frac{1}{\mathsf{D}^{2}+G^{-1}(z)}$
$\Leftrightarrow\mathbb{E}\frac{-\mathsf{D}^{2}}{\left(\mathsf{D}^{2}+G^{-1}(z)\right)^{2}}<0$
where we used in the second line that $R(z)=G^{-1}(z)-1/z<0$ from (b). Note that the last line is true since $\mathsf{D}^{2}\neq 0$ with positive probability. (e) trivially follows from (c). (f) follows from [69, Notation 12.6, Proposition 13.15]. ∎
A.4 DF adjustment coincide with Spectrum-Aware adjustment under Marchenko-Pastur law
Lemma A.9.
If the empirical distribution of the eigenvalues of $\mathbf{X}^{\top}\mathbf{X}$ weakly converges Marchenko-Pastur law, then |$\absolutevalue{\widehat{\mathsf{adj}}-\breve{\mathsf{adj}}}\to 0.$
Proof of Lemma A.9.
By weak convergence,
$\frac{1}{p}\sum_{i=1}^{p}\frac{-1}{d_{i}^{2}+\lambda_{2}}\rightarrow G\left(-\lambda_{2}\right)$
where $z\mapsto G(z)$ is the Cauchy transform of Marchenko-Pastur law [^4] Here, $G(z):=$ $\int\frac{1}{z-x}\mu(dx)$ where $\mu(\cdot)$ is measure associated to Marchenko-Pasteur law. . Then we have that
$\widehat{\mathsf{adj}}\rightarrow\lambda_{2}\left(\frac{1}{1+\lambda_{2}G\left(-\lambda_{2}\right)}-1\right)^{-1},\quad\breve{\mathsf{adj}}\rightarrow 1-\delta^{-1}\left(1+\lambda_{2}G\left(-\lambda_{2}\right)\right)$ (47)
Observe that the limiting values of $\widehat{\mathsf{adj}}$ and $\breve{\mathsf{adj}}$ above are equal if and only if the following holds
$01+\left(\lambda_{2}+1-\delta^{-1}\right)G\left(-\lambda_{2}\right)-\delta^{-1}\lambda_{2}\quantity(G\left(-\lambda_{2}\right))^{2}=$ 0 . (48)
Here, (48) indeed holds true since $G\left(-\lambda_{2}\right)$ is one of the root of the quadratic equation (48). This is by referencing the explicit expression of the Cauchy transform of the Marchenko-Pastur law (cf. [3, Lemma 3.11]). ∎
A.5 Fixed point equation
A.5.1 An auxiliary lemma
Lemma A.10.
Under Assumption 4 and 6,
$\mathbb{P}(\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)\neq 0)>0,\quad\mathbb{P}(\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)\neq 1)>0$ (49)
$\mathbb{P}(h^{\prime\prime}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)\right)\neq+\infty)>0,\quad\mathbb{P}(h^{\prime\prime}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)\right)\neq 0)>0$
where $\mathsf{Z}\sim N(0,1)$ is independent of $\mathsf{B}^{\star}$.
Proof of Lemma A.10.
Note that $\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)\neq 0$ with positive probability or else $\frac{\gamma_{*}}{\eta_{*}}=\mathbb{E}\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)=0$ which violates Assumption 6. Meanwhile, $\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)\neq 1$ with positive probability or else $\frac{\gamma_{*}}{\eta_{*}}=\mathbb{E}\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)=1$ which also violates Assumption 6. The inequalities in the second line follows immediately from (16) and that $\gamma_{*}>0$. ∎
A.5.2 Uniqueness of fixed points given existence
Suppose that Assumption 26 hold. Our proof of Theorem 5.1 and Theorem 5.2 does not require $(\gamma_{*},\eta_{*},\tau_{*},\tau_{**})$ to be a unique solution of (20), only that it is one of the solutions. However, if there are two different solutions of (20), it would lead to a contradiction in Theorem 5.2. More concretely, suppose that there exists two different solutions of (20): $x^{(1)}:=\quantity(\gamma_{*}^{(1)},\eta_{*}^{(1)},\tau_{*}^{(1)},\tau_{**}^{(1)})$ and $x^{(2)}:=\quantity(\gamma_{*}^{(2)},\eta_{*}^{(2)},\tau_{*}^{(2)},\tau_{**}^{(2)})$. By Theorem 5.2, we would have $\quantity(\widehat{\mathsf{adj}},\hat{\eta}_{*},\hat{\tau}_{*},\hat{\tau}_{**})$ converges almost surely to both $x^{(1)}$ and $x^{(2)}$, hence the contradiction.
A.5.3 Existence of fixed points for Elastic Net
Given that $h$ is the Elastic Net penalty as in (17) with $\lambda_{2}>0$, we now prove show that Assumption 6 holds under Assumption 2 and 3. First, eliminate the variable $\tau_{**}$ from (20) via (20b) and introduce change of variable $\tau_{*}=\gamma_{*}^{-2}\alpha_{*}^{-2}$ for some new variable $\alpha_{*}>0$. We then obtain a new system of fixed equation
$\gamma_{*}^{-1}=\frac{1}{-R\left(\eta_{*}^{-1}\right)}$ (50a)
$\eta_{*}^{-1}=\gamma_{*}^{-1}\mathbb{E}\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}\left(\mathsf{B}^{\star}+\frac{\gamma_{*}^{-1}}{\alpha_{*}}\mathsf{Z}\right)$ (50b)
$1=\alpha_{*}^{2}R^{\prime}\left(\eta_{*}^{-1}\right)\mathbb{E}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\mathsf{B}^{\star}+\frac{\gamma_{*}^{-1}}{\alpha_{*}}\mathsf{Z}\right)-\mathsf{B}^{\star}\right)^{2}+\sigma^{2}\frac{\alpha_{*}^{2}}{\gamma_{*}^{-1}}\left[1+\frac{\eta_{*}^{-1}R^{\prime}\left(\eta_{*}^{-1}\right)}{R\left(\eta_{*}^{-1}\right)}\right]$ (50c)
Note that Assumption 6 holds if and only if we can find a solution $\gamma_{*}^{-1},\eta_{*}^{-1},\alpha_{*}>0$ for the above.
Denote $\gamma_{+}^{-1}=\lim_{z\rightarrow G\left(-d_{-}\right)}\frac{1}{-R(z)}=\frac{1}{\frac{1}{G\left(-d_{-}\right)}+d_{-}}$. Note that $\gamma_{+}^{-1}\in\left(\frac{1}{\mathbb{E}\mathsf{D}^{2}},G\left(-d_{-}\right)\right]$since $x\mapsto\frac{1}{-R(z)}$ is strictly increasing on its domain $\left(0,G\left(-d_{-}\right)\right)$ using Lemma A.8, (b) and $\lim_{z\rightarrow 0}\frac{1}{-R(z)}=\frac{1}{\mathbb{E}\mathsf{D}^{2}}$ using Lemma A.8, (f). Let us define the functions $f_{1}:\left[\frac{1}{\mathbb{E}\mathsf{D}^{2}},\gamma_{+}^{-1}\right)\mapsto\left[0,G\left(-d_{-}\right)\right)$ as the inverse function of $x\mapsto\frac{1}{-R(z)}$. Note that $f_{1}$ is well-defined and strictly increasing on its domain. It also satisfies
$f_{1}\left(\frac{1}{\mathbb{E}\mathsf{D}^{2}}\right)=0,\qquad\lim_{\gamma^{-1}\rightarrow\gamma_{+}^{-1}}f_{1}\left(\gamma^{-1}\right)=G\left(-d_{-}\right).$
Let us define function $f_{2}:(0,+\infty)\times(0,+\infty)\mapsto(0,+\infty)$ such that
$f_{2}\left(\gamma^{-1},\alpha\right)=\gamma^{-1}\mathbb{E}\operatorname{Prox}_{\gamma^{-1}h}^{\prime}\left(\mathsf{B}^{\star}+\frac{\gamma^{-1}}{\alpha}\mathsf{Z}\right)=\frac{\gamma^{-1}}{1+\lambda_{2}\gamma^{-1}}\mathbb{P}\left(\left|\frac{1}{\gamma^{-1}}\mathsf{B}^{\star}+\frac{1}{\alpha}\mathsf{Z}\right|>\lambda_{1}\right).$
Now we study the equation (in terms of $\gamma^{-1}$ )
$f_{1}\left(\gamma^{-1}\right)=f_{2}\left(\gamma^{-1},\alpha\right).$
Observe that this equation amounts to eliminating $\eta_{*}^{-1}$ and solving for $\gamma_{*}^{-1}$ in terms $\alpha_{*}$ from (50a) and (50b). We claim that for any fixed $\alpha>0$, there is at least one solution $\gamma^{-1}(\alpha)\in\left[\frac{1}{\mathbb{E}\mathsf{D}^{2}},\gamma_{+}^{-1}\right)$. To see the claim, note that since
$f_{2}\left(\frac{1}{\mathbb{E}\mathsf{D}^{2}},\alpha\right)=\frac{\frac{1}{\mathbb{E}\mathsf{D}^{2}}}{1+\lambda_{2}\frac{1}{\mathbb{E}\mathsf{D}^{2}}}\mathbb{P}\left(\left|\mathsf{B}^{\star}\mathbb{E}\mathsf{D}^{2}+\frac{1}{\alpha}\mathsf{Z}\right|>\lambda_{1}\right)>0=f_{1}\left(\frac{1}{\mathbb{E}\mathsf{D}^{2}}\right),$
a sufficient condition for $f_{1}\left(\gamma^{-1}\right)=f_{2}\left(\gamma^{-1},\alpha\right)$ to have a solution on $\left[\frac{1}{\mathbb{E}\mathsf{D}^{2}},\gamma_{+}^{-1}\right)$ is $f_{2}\left(\gamma_{+}^{-1},\alpha\right)<$ $f_{1}\left(\gamma_{+}^{-1}\right)$. When $\gamma_{+}^{-1}<+\infty$, it suffices that
$\frac{\gamma_{+}^{-1}}{1+\lambda_{2}\gamma_{+}^{-1}}\mathbb{P}\left(\left|\frac{1}{\gamma_{+}^{-1}}\mathsf{B}^{\star}+\frac{1}{\alpha}\mathsf{Z}\right|>\lambda_{1}\right)<G\left(-d_{-}\right)$
which holds since
$\frac{\gamma_{+}^{-1}}{1+\lambda_{2}\gamma_{+}^{-1}}\mathbb{P}\left(\left|\frac{1}{\gamma_{+}^{-1}}\mathsf{B}^{\star}+\frac{1}{\alpha}\mathsf{Z}\right|>\lambda_{1}\right)<\gamma_{+}^{-1}=\frac{1}{\frac{1}{G\left(-d_{-}\right)}+d_{-}}\leq G\left(-d_{-}\right)$
and when $\gamma_{+}^{-1}=+\infty$, it suffices that
$\frac{1}{\lambda_{2}}\mathbb{P}\left(\left|\frac{1}{\alpha}\mathsf{Z}\right|>\lambda_{1}\right)<G\left(-d_{-}\right)$
which holds because $\gamma_{+}^{-1}=+\infty$ if and only if $d_{-}=0$ and $G\left(-d_{-}\right)=+\infty$. This means that for any $\alpha>0$, we can find $\gamma^{-1}(\alpha)$ and $\eta^{-1}(\alpha)=f_{1}(\gamma^{-1}(\alpha))=f_{2}(\gamma^{-1}(\alpha),\alpha)$ that solves (50a) and (50b). The plan is to plug $\gamma^{-1}(\alpha)$ and $\eta^{-1}(\alpha)$ into the RHS of (50c) to obtain the function $v:(0,+\infty)\mapsto(0,+\infty)$
$\begin{gathered}v(\alpha)=\alpha^{2}R^{\prime}\left(\eta^{-1}(\alpha)\right)\left[\mathbb{E}\left(\operatorname{Prox}_{\gamma^{-1}(\alpha)h}\left(\mathsf{B}^{\star}+\frac{\gamma^{-1}(\alpha)}{\alpha}\mathsf{Z}\right)-\mathsf{B}^{\star}\right)^{2}\right]\\ +\sigma^{2}\alpha^{2}\frac{1}{\gamma^{-1}(\alpha)}\left[1+\frac{\eta^{-1}(\alpha)R^{\prime}\left(\eta^{-1}(\alpha)\right)}{R\left(\eta^{-1}(\alpha)\right)}\right]\end{gathered}$
and show that the RHS of (50c), i.e. $v(\alpha)$, diverges to +$\infty$+ ∞ as $\alpha\to+\infty$ and goes to some value less than $1$ as $\alpha\to 0$.
First consider any positive increasing sequence $\left(\alpha_{m}\right)_{m=1}^{+\infty}$ such that $\alpha_{m}\rightarrow+\infty$ as $m\rightarrow\infty$. We claim that one can pick solution $\gamma^{-1}\left(\alpha_{m}\right)\in\left[\frac{1}{\mathbb{E}\mathsf{D}^{2}},\gamma_{+}^{-1}\right)$ for each $m$ such that $C_{1}:=$ $\limsup_{m\rightarrow\infty}\gamma^{-1}\left(\alpha_{m}\right)<\gamma_{+}^{-1}$. For the case where $\gamma_{+}^{-1}<+\infty$, suppose the claim is false by which we would have a subsequence $\left(\alpha_{m_{i}}\right)_{i=1}^{+\infty}$ of $\left(\alpha_{m}\right)_{m=1}^{+\infty}$ such that as $i\rightarrow\infty$,
$f_{2}\left(\gamma_{m_{i}}^{-1},\alpha_{m_{i}}\right)=$ $\frac{\gamma^{-1}\left(\alpha_{m_{i}}\right)}{1+\lambda_{2}\gamma^{-1}\left(\alpha_{m_{i}}\right)}\mathbb{P}\left(\left|\frac{1}{\gamma^{-1}\left(\alpha_{m_{i}}\right)}\mathsf{B}^{\star}+\frac{1}{\alpha_{m_{i}}}\mathsf{Z}\right|>\lambda_{1}\right)=f_{1}\left(\gamma^{-1}\left(\alpha_{m_{i}}\right)\right)$
$\rightarrow G\left(-d_{-}\right)$
which is impossible since we also have that as $i\rightarrow\infty$,
$\frac{\gamma^{-1}\left(\alpha_{m_{i}}\right)}{1+\lambda_{2}\gamma^{-1}\left(\alpha_{m_{i}}\right)}\mathbb{P}\left(\left|\frac{1}{\gamma^{-1}\left(\alpha_{m_{i}}\right)}\mathsf{B}^{\star}+\frac{1}{\alpha_{m_{i}}}\mathsf{Z}\right|>\lambda_{1}\right)\rightarrow\frac{\gamma_{+}^{-1}}{1+\lambda_{2}\gamma_{+}^{-1}}\mathbb{P}\left(\left|\frac{1}{\gamma_{+}^{-1}}\mathsf{B}^{\star}\right|>\lambda_{1}\right)$
$<\gamma_{+}^{-1}\leq G\left(-d_{-}\right)$
For the case $\gamma_{+}^{-1}=+\infty$, suppose the claim is false by which we would have a subsequence $\left(\alpha_{m_{i}}\right)_{i=1}^{+\infty}$ of $\left(\alpha_{m}\right)_{m=1}^{+\infty}$ such that as $i\rightarrow\infty$,
$f_{2}\left(\gamma_{m_{i}}^{-1},\alpha_{m_{i}}\right)$ $=\frac{\gamma^{-1}\left(\alpha_{m_{i}}\right)}{1+\lambda_{2}\gamma^{-1}\left(\alpha_{m_{i}}\right)}\mathbb{P}\left(\left|\frac{1}{\gamma^{-1}\left(\alpha_{m_{i}}\right)}\mathsf{B}^{\star}+\frac{1}{\alpha_{m_{i}}}\mathsf{Z}\right|>\lambda_{1}\right)$
$=f_{1}\left(\gamma^{-1}\left(\alpha_{m_{i}}\right)\right)\rightarrow+\infty$
because $\lim_{\gamma^{-1}\rightarrow\gamma_{+}^{-1}}f_{1}\left(\gamma^{-1}\right)=G\left(-d_{-}\right)=+\infty$ when $\gamma_{+}^{-1}=+\infty$. But this is impossible since the LHS is strictly less than $\lambda_{2}^{-1}$.
Recall that $\eta^{-1}\left(\alpha_{m}\right)=f_{1}\left(\gamma^{-1}\left(\alpha_{m}\right)\right)$. It follows from the above discussion that
$\limsup_{m\rightarrow\infty}\eta^{-1}\left(\alpha_{m}\right)<G\left(-d_{-}\right)$
from which we conclude that
$C_{2}:=\liminf_{m\rightarrow\infty}1+\frac{\eta^{-1}\left(\alpha_{m}\right)R^{\prime}\left(\eta^{-1}\left(\alpha_{m}\right)\right)}{R\left(\eta^{-1}\left(\alpha_{m}\right)\right)}>0$
This follows from the fact that $\lim_{x\rightarrow 0}1+\frac{xR^{\prime}(x)}{R(x)}=1$ using Lemma A.8, (f) and continuity of the function $x\mapsto$ $1+\frac{xR^{\prime}(x)}{R(x)}$ on $\left(0,G\left(-d_{-}\right)\right)$. Note that by the above discussion, we have $\liminf_{\alpha\rightarrow+\infty}\frac{v(\alpha)}{\alpha^{2}}\geq\frac{c_{2}}{c_{1}}$ which then implies that
$\liminf_{\alpha\rightarrow+\infty}v(\alpha)\rightarrow+\infty.$ (51)
Now consider any positive decreasing sequence $\left(\alpha_{m}\right)_{m=1}^{+\infty}$ such that $\alpha_{m}\rightarrow 0$ as $m\rightarrow\infty$. We again claim that one can pick the solution $\gamma^{-1}\left(\alpha_{m}\right)\in\left[\frac{1}{\mathbb{E}\mathsf{D}^{2}},\gamma_{+}^{-1}\right)$ for each $m$ such that $C_{3}:=$ $\limsup_{m\rightarrow\infty}\gamma^{-1}\left(\alpha_{m}\right)\in\left[\frac{1}{\mathbb{E}\mathsf{D}^{2}},+\infty\right)$ and $\limsup_{m\rightarrow\infty}\eta^{-1}\left(\alpha_{m}\right)<G\left(-d_{-}\right)$. The proofs are analogous to the previous case. We then have that
$C_{4}:=\limsup_{m\rightarrow\infty}1+\frac{\eta^{-1}\left(\alpha_{m}\right)R^{\prime}\left(\eta^{-1}\left(\alpha_{m}\right)\right)}{R\left(\eta^{-1}\left(\alpha_{m}\right)\right)}<+\infty$
This follows from the fact that $\lim_{x\rightarrow 0}1+\frac{xR^{\prime}(x)}{R(x)}=1$ by Lemma A.8, (f), $\limsup_{m\rightarrow\infty}\eta^{-1}\left(\alpha_{m}\right)<G\left(-d_{-}\right)$ as just shown, and continuity of the function $x\mapsto 1+\frac{xR^{\prime}(x)}{R(x)}$ on $\left(0,G\left(-d_{-}\right)\right)$. We can then conclude that
$\limsup_{m\rightarrow+\infty}\frac{1}{\gamma^{-1}\left(\alpha_{m}\right)}\left[1+\frac{\eta^{-1}\left(\alpha_{m}\right)R^{\prime}\left(\eta^{-1}\left(\alpha_{m}\right)\right)}{R\left(\eta^{-1}\left(\alpha_{m}\right)\right)}\right]\leq C_{4}\mathbb{E}\mathsf{D}^{2}$
which then implies that
$\limsup_{m\rightarrow+\infty}\frac{\sigma^{2}\alpha_{m}^{2}}{\gamma^{-1}\left(\alpha_{m}\right)}\left[1+\frac{\eta^{-1}\left(\alpha_{m}\right)R^{\prime}\left(\eta^{-1}\left(\alpha_{m}\right)\right)}{R\left(\eta^{-1}\left(\alpha_{m}\right)\right)}\right]=0.$ (52)
Note also that
$\limsup_{m\rightarrow+\infty}\left|R^{\prime}\left(\eta^{-1}\left(\alpha_{m}\right)\right)\right|<+\infty$ (53)
since $\underset{m\rightarrow\infty}{\limsup}\eta^{-1}\left(\alpha_{m}\right)<G\left(-d_{-}\right)$. For each $m$, we have
$\eta^{-1}\left(\alpha_{m}\right)=$ $f_{2}\left(\gamma^{-1}\left(\alpha_{m}\right),\alpha_{m}\right)$
$=\frac{\gamma^{-1}\left(\alpha_{m}\right)}{1+\lambda_{2}\gamma^{-1}\left(\alpha_{m}\right)}\mathbb{P}\left(\left|\alpha_{m}\mathsf{B}^{\star}+\gamma^{-1}\left(\alpha_{m}\right)\mathsf{Z}\right|>\alpha_{m}\gamma^{-1}\left(\alpha_{m}\right)\lambda_{1}\right)$
It follows that
$\limsup_{m\rightarrow+\infty}\left|\left(\frac{\gamma^{-1}\left(\alpha_{m}\right)}{1+\lambda_{2}\gamma^{-1}\left(\alpha_{m}\right)}\right)^{2}-\eta^{-2}\left(\alpha_{m}\right)\right|=0$ (54)
Also note that
$\begin{gathered}\alpha_{m}^{2}\mathbb{E}\left(\operatorname{Prox}_{\gamma^{-1}\left(\alpha_{m}\right)h}\left(\mathsf{B}^{\star}+\frac{\gamma^{-1}\left(\alpha_{m}\right)}{\alpha_{m}}\mathsf{Z}\right)-\mathsf{B}^{\star}\right)^{2}\\ =\mathbb{E}\left(\frac{\operatorname{sgn}(\alpha_{m}\mathsf{B}^{\star}+\gamma_{*}^{-1}\mathsf{Z})}{1+\lambda_{2}\gamma^{-1}(\alpha_{m})}\left(\left|\alpha_{m}\mathsf{B}^{\star}+\gamma^{-1}\left(\alpha_{m}\right)\mathsf{Z}\right|\right.\right.\left.\left.-\alpha_{m}\gamma^{-1}\left(\alpha_{m}\right)\lambda_{1}\right)_{+}-\alpha_{m}\mathsf{B}^{\star}\right)^{2}\end{gathered}$
Since $\gamma^{-1}\left(\alpha_{m}\right)$ is bounded between $\frac{1}{\mathbb{E}\mathsf{D}^{2}}$ and $C_{3}$ for all sufficiently large $m$, it follows from the dominated convergence theorem that
$\limsup_{m\rightarrow+\infty}\left|\left(\frac{\gamma^{-1}\left(\alpha_{m}\right)}{1+\lambda_{2}\gamma^{-1}\left(\alpha_{m}\right)}\right)^{2}-\alpha_{m}^{2}\mathbb{E}\left(\operatorname{Prox}_{\gamma^{-1}\left(\alpha_{m}\right)h}\left(\mathsf{B}^{\star}+\frac{\gamma^{-1}\left(\alpha_{m}\right)}{\alpha_{m}}\mathsf{Z}\right)-\mathsf{B}^{\star}\right)^{2}\right|=0$ (55)
Combining (53), (54) and (55), we obtain that
$\limsup_{m\rightarrow+\infty}\bigg{|}\eta^{-2}$ $\left(\alpha_{m}\right)R^{\prime}\left(\eta^{-1}\left(\alpha_{m}\right)\right)$ (56)
$-\alpha_{m}^{2}R^{\prime}\left(\eta^{-1}\left(\alpha_{m}\right)\right)\mathbb{E}\left(\operatorname{Prox}_{\gamma^{-1}\left(\alpha_{m}\right)h}\left(\mathsf{B}^{\star}+\frac{\gamma^{-1}\left(\alpha_{m}\right)}{\alpha_{m}}\mathsf{Z}\right)-\mathsf{B}^{\star}\right)^{2}\bigg{|}=0.$
Now note that
$\limsup_{m\rightarrow+\infty}\eta^{-2}\left(\alpha_{m}\right)R^{\prime}\left(\eta^{-1}\left(\alpha_{m}\right)\right)<1$
using Lemma A.8, (e), $\limsup_{m\rightarrow\infty}\eta^{-1}\left(\alpha_{m}\right)^{m\rightarrow+\infty}<G\left(-d_{-}\right)$ from the above and the continuity of $x\mapsto x^{2}R^{\prime}(x)$ on $\left(0,G\left(-d_{-}\right)\right)$. Using this and (56), we may then conclude that
$\limsup_{m\rightarrow+\infty}\alpha_{m}^{2}R^{\prime}\left(\eta^{-1}\left(\alpha_{m}\right)\right)\mathbb{E}\left(\operatorname{Prox}_{\gamma^{-1}\left(\alpha_{m}\right)h}\left(\mathsf{B}^{\star}+\frac{\gamma^{-1}\left(\alpha_{m}\right)}{\alpha_{m}}\mathsf{Z}\right)-\mathsf{B}^{\star}\right)^{2}<1$
which along with (52) implies that
$\limsup_{\alpha\rightarrow 0}v(\alpha)<1.$ (57)
Combine (51) and (57). By continuity of $\alpha\mapsto v(\alpha)$ on $(0,+\infty)$, we know that there exists a solution $\alpha_{*}\in(0,+\infty)$ to the equation $v\left(\alpha_{*}\right)=1$. Therefore, a solution of (50) is $(\gamma,\eta,\alpha)=\left(\gamma^{-1}\left(\alpha_{*}\right),\eta^{-1}\left(\alpha_{*}\right),\alpha_{*}\right)$ by construction. This concludes the proof.
A.6 VAMP algorithm
For $\sigma^{2}=1$, the VAMP algorithm consists of iteration as follows: for $t\geq 1$,
$\hat{\mathbf{x}}_{1t}=\operatorname{Prox}_{\gamma_{1,t-1}^{-1}}\left(\mathbf{r}_{1,t-1}\right),\quad\eta_{1t}^{-1}=\gamma_{1,t-1}^{-1}\nabla\cdot\operatorname{Prox}_{\gamma_{1,t-1}^{-1}h}\left(\mathbf{r}_{1,t-1}\right)$
$\gamma_{2t}=\eta_{1t}-\gamma_{1,t-1},\quad\mathbf{r}_{2t}=\left(\eta_{1t}\hat{\mathbf{x}}_{1t}-\gamma_{1,t-1}\mathbf{r}_{1,t-1}\right)/\gamma_{2t}$
$\hat{\mathbf{x}}_{2t}=\left(\mathbf{X}^{\top}\mathbf{X}+\gamma_{2t}\mathbf{I}_{p}\right)^{-1}\left(\mathbf{X}^{\top}\mathbf{y}+\gamma_{2t}\mathbf{r}_{2t}\right),\quad\eta_{2t}^{-1}=\frac{1}{p}\operatorname{Tr}\left[\left(\mathbf{X}^{\top}\mathbf{X}+\gamma_{2t}\mathbf{I}_{p}\right)^{-1}\right]$
$\gamma_{1t}=\eta_{2t}-\gamma_{2t},\quad\mathbf{r}_{1t}=\left(\eta_{2t}\hat{\mathbf{x}}_{2t}-\gamma_{2t}\mathbf{r}_{2t}\right)/\gamma_{1t}$
The algorithm can be initialized at $r_{10}\in\mathbb{R}^{p},\gamma_{10},\tau_{10}>0$ such that $\left(r_{10},\bm{\beta}^{\star}\right)\stackrel{{\scriptstyle W_{2}}}{{\to}}\left(\mathsf{R}_{10},\mathsf{B}^{\star}\right)$ and $\mathsf{R}_{10}-\mathsf{B}^{\star}\sim N\left(0,\tau_{10}\right)$. This algorithm is first introduced in [74] and the iterates $\hat{\mathbf{x}}_{1t},\hat{\mathbf{x}}_{2t}$ are supposed to track $\hat{\bm{\beta}}$. The performance of this algorithm is characterized by state evolution iterations: for $t\geq 1$,
$\bar{\alpha}_{1t}=\mathbb{E}\operatorname{Prox}_{\gamma_{1,1-1}^{\prime}}^{\prime}\left(\mathsf{B}^{\star}+N\left(0,\tau_{1,t-1}\right)\right),\quad\bar{\eta}_{1t}^{-1}=\bar{\gamma}_{1,t-1}^{-1}\bar{\alpha}_{1t}$ (58)
$\bar{\gamma}_{2t}=\bar{\eta}_{1t}-\bar{\gamma}_{1,t-1},\quad\tau_{2t}=\frac{1}{\left(1-\bar{\alpha}_{1t}\right)^{2}}\left[\mathcal{E}_{1}\left(\bar{\gamma}_{1,t-1},\tau_{1,t-1}\right)-\bar{\alpha}_{1t}^{2}\tau_{1,t-1}\right]$
$\bar{\alpha}_{2t}=\bar{\gamma}_{2t}\mathbb{E}\frac{1}{\mathsf{D}^{2}+\bar{\gamma}_{2t}},\quad\bar{\eta}_{2t}^{-1}=\bar{\gamma}_{2t}^{-1}\bar{\alpha}_{2t}$
$\bar{\gamma}_{1,t}=\bar{\eta}_{2t}-\bar{\gamma}_{2t},\quad\tau_{t}=\frac{1}{\left(1-\bar{\alpha}_{2t}\right)^{2}}\left[\mathcal{E}_{2}\left(\bar{\gamma}_{2t},\tau_{2t}\right)-\bar{\alpha}_{2t}^{2}\tau_{2t}\right]$
where
$\mathcal{E}_{1}\left(\gamma_{1},\tau\right):=\mathbb{E}\left(\operatorname{Prox}_{\gamma_{1}^{-1}h}\left(\mathsf{B}^{\star}+N\left(0,\tau\right)\right)-\mathsf{B}^{\star}\right)^{2},\quad\mathcal{E}_{2}\left(\gamma_{2},\tau_{2}\right):=\mathbb{E}\left[\frac{\mathsf{D}^{2}+\tau_{2}\gamma_{2}^{2}}{\left(\mathsf{D}^{2}+\gamma_{2}\right)^{2}}\right].$
Appendix B Spectrum-Aware Debiasing
B.1 Proof result A: Distribution characterization
In this section, we prove Theorem 5.1 using VAMP algorithm as proof device. We define the version of VAMP algorithm we will use in Section B.1.1, prove Cauchy convergence of its iterates in Section B.1.2, and prove Theorem 5.1 in Section B.1.3. To streamline the presentation, proofs of intermediate claims are collected in Section B.3. We also assume without loss of generality that
$\sigma^{2}=1$
for the remainder of this section. The general case for arbitrary $\sigma^{2}>0$ follows from a simple rescaling argument.
B.1.1 The oracle VAMP algorithm
We review the oracle VAMP algorithm defined in [40] and present an extended state evolution result for the algorithm. This algorithm is obtained by initializing the VAMP algorithm introduced in [74] at stationarity $\mathbf{r}_{10}=\bm{\beta}^{\star}+N(\rm{0},\tau_{*}\mathbf{I}_{p}),\gamma_{10}^{-1}=\gamma_{*}^{-1}$. See Section A.6 for a review. Then for $t\geq 1$, we have iterates
$\hat{\mathbf{x}}_{1t}=\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\mathbf{r}_{1,t-1}\right)$ (59a)
$\mathbf{r}_{2t}=\frac{1}{\eta_{*}-\gamma_{*}}\left(\eta_{*}\hat{\mathbf{x}}_{1t}-\gamma_{*}\mathbf{r}_{1,t-1}\right)$ (59b)
$\hat{\mathbf{x}}_{2t}=\left(\mathbf{X}^{\top}\mathbf{X}+\left(\eta_{*}-\gamma_{*}\right)\mathbf{I}_{p}\right)^{-1}\left(\mathbf{X}^{\top}\mathbf{y}+\left(\eta_{*}-\gamma_{*}\right)\mathbf{r}_{2t}\right)$ (59c)
$\mathbf{r}_{1t}=\frac{1}{\gamma_{*}}\left(\eta_{*}\hat{\mathbf{x}}_{2t}-\left(\eta_{*}-\gamma_{*}\right)\mathbf{r}_{2t}\right)$ (59d)
Let us define functions $F:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ and $F^{\prime}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$
$F(q,x):=\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\operatorname{Prox}_{\gamma_{*}^{-1}h}(q+x)-\frac{\gamma_{*}}{\eta_{*}-\gamma_{*}}q-\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}x$ (60)
$F^{\prime}(q,x):=\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\operatorname{Prox}^{\prime}_{\gamma_{*}^{-1}h}(q+x)-\frac{\gamma_{*}}{\eta_{*}-\gamma_{*}}$
Note that for any fixed $x$, $F^{\prime}(q,x)$ equals to the derivative of $q\mapsto F(q,x)$ whenever the derivative exists, and at the finitely many points where $q\mapsto F(q,x)$ is not differentiable $F^{\prime}(q,x)$ equals to 00 (cf. Lemma 2.7). We also define some quantities
$\bm{\Lambda}:=\frac{\eta_{*}\left(\eta_{*}-\gamma_{*}\right)}{\gamma_{*}}\left(\mathbf{D}^{\top}\mathbf{D}+\left(\eta_{*}-\gamma_{*}\right)\mathbf{I}_{p}\right)^{-1}-\left(\frac{\eta_{*}-\gamma_{*}}{\gamma_{*}}\right)\cdot\mathbf{I}_{p}$ (61)
$\bm{\xi}:=\mathbf{Q}\bm{\varepsilon},\quad\mathbf{e}_{b}:=\frac{\eta_{*}}{\gamma_{*}}\left(\mathbf{D}^{\top}\mathbf{D}+\left(\eta_{*}-\gamma_{*}\right)\mathbf{I}_{p}\right)^{-1}\mathbf{D}^{\top}\bm{\xi},\quad\mathbf{e}:=\mathbf{O}^{\top}\mathbf{e}_{b}$
We note some important properties of these quantities, which are essentially consequence of Assumption 2 and (20). We defer the proof of Proposition B.1 to Section B.3.1.
Proposition B.1.
Under Assumption 24, almost surely,
$\lim_{p\rightarrow\infty}\frac{1}{p}\operatorname{Tr}(\bm{\Lambda})=0,\quad\kappa_{*}:=\lim_{p\rightarrow\infty}\frac{1}{p}\operatorname{Tr}\left(\bm{\Lambda}^{2}\right)=\mathbb{E}\left(\frac{\eta_{*}\left(\eta_{*}-\gamma_{*}\right)}{\gamma_{*}\left(\mathsf{D}^{2}+\left(\eta_{*}-\gamma_{*}\right)\right)}-\frac{\eta_{*}-\gamma_{*}}{\gamma_{*}}\right)^{2}$ (62a)
$b_{*}:=\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\mathbf{e}_{b}\right\|^{2}=\frac{1}{\gamma_{*}}-\frac{\kappa_{*}}{\eta_{*}-\gamma_{*}}=\left(\frac{\eta_{*}}{\gamma_{*}}\right)^{2}\mathbb{E}\frac{\mathsf{D}^{2}}{\left(\mathsf{D}^{2}+\eta_{*}-\gamma_{*}\right)^{2}},\quad\tau_{*}=b_{*}+\kappa_{*}\tau_{**}$ (62b)
$\mathbb{E}F^{\prime}\left(\sqrt{\tau_{*}}\mathsf{Z},\mathsf{B}^{\star}\right)=0,\quad\mathbb{E}F\left(\sqrt{\tau_{*}}\mathsf{Z},\mathsf{B}^{\star}\right)^{2}=\tau_{**}$ (62c)
where $\mathsf{Z}\sim N(0,1)$ is independent of $\mathsf{B}^{\star}$. Moreover, the function $(q,x)\mapsto F(q,x)$ is Lipschitz continuous on $\mathbb{R}\times\mathbb{R}$.
Then, one can show that by eliminating $\hat{\mathbf{x}}_{1t},\hat{\mathbf{x}}_{2t}$ and introducing a change of variables
$\mathbf{x}^{t}=\mathbf{r}_{2t}-\bm{\beta}^{\star},\quad\mathbf{y}^{t}=\mathbf{r}_{1t}-\bm{\beta}^{\star}-\mathbf{e},\quad\mathbf{s}^{t}=\mathbf{O}\mathbf{x}^{t}$ (63)
(59) is equivalent to the following iterations: with initialization $\mathbf{q}^{0}\sim N(\rm{0},\tau_{*}\cdot\mathbf{I}_{p}),{\mathbf{x}}^{1}=F(q_{0},\bm{\beta}^{\star})$, for $t=1,2,3,\ldots,$
$\mathbf{s}^{t}=\mathbf{O}\mathbf{x}^{t},\qquad\mathbf{y}^{t}=\mathbf{O}^{\top}\bm{\Lambda}\mathbf{s}^{t},\qquad\mathbf{x}^{t+1}=F(\mathbf{y}^{t}+\mathbf{e},\bm{\beta}^{\star}).$ (64)
The following Proposition will be needed later. Its proof is deferred to Section B.3.1.
Proposition B.2.
Suppose Assumptions 23 hold. Define random variables
$\Xi\sim N(0,1),\qquad\mathsf{P}_{0}\sim N(0,\tau_{*}),\qquad\mathsf{E}\sim N(0,b_{*})$
independent of each other and of $\mathsf{D}$, and set
$\mathsf{L}=\frac{\eta_{*}-\gamma_{*}}{\gamma_{*}}\left(\frac{\eta_{*}}{\mathsf{D}^{2}+\eta_{*}-\gamma_{*}}-1\right),\;\mathsf{E}_{b}=\frac{\eta_{*}}{\gamma_{*}}\frac{\mathsf{D}\Xi}{\mathsf{D}^{2}+\eta_{*}-\gamma_{*}},\;\mathsf{H}=(\mathsf{B}^{\star},\mathsf{D},\mathsf{D}\Xi,\mathsf{L},\mathsf{E}_{b},\mathsf{E},\mathsf{P}_{0}).$
Then $\kappa_{*}=\mathbb{E}\mathsf{L}^{2}$ and $b_{*}=\mathbb{E}\mathsf{E}_{b}^{2}$. Furthermore, almost surely as $n,p\to\infty$,
$\mathbf{H}:=\left(\bm{\beta}^{\star},\mathbf{D}^{\top}\bm{1}_{n\times 1},\mathbf{D}^{\top}\bm{\xi},\operatorname{diag}(\bm{\Lambda}),\mathbf{e}_{b},\mathbf{e},\mathbf{q}^{0}\right)\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{H}.$
Now we state the state evolution for the VAMP algorithm. Its proof is deferred to Section B.3.1.
Proposition B.3.
Suppose Assumptions 26 hold. Further assume that the function $x\mapsto\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}(x)$ defined in Lemma 2.7 is non-constant. Let $\mathsf{H}=(\mathsf{B}^{\star},\mathsf{D},\mathsf{D}\Xi,\mathsf{L},\mathsf{E}_{b},\mathsf{E},\mathsf{P}_{0})$ be as defined in Proposition B.2. Set $\mathsf{X}_{1}=F(\mathsf{P}_{0},\mathsf{B}^{\star})$, set $\Delta_{1}=\mathbb{E}[\mathsf{X}_{1}^{2}]\in\mathbb{R}^{1\times 1}$, and define iteratively $\mathsf{S}_{t},\mathsf{Y}_{t},\mathsf{X}_{t+1},\Delta_{t+1}$ for $t=1,2,3,\ldots$ such that
$(\mathsf{S}_{1},\ldots,\mathsf{S}_{t})\sim N(\rm{0},\Delta_{t}),\qquad(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{t})\sim N(\rm{0},\kappa_{*}\Delta_{t})$
are Gaussian vectors independent of each other and of $\mathsf{H}$, and
$\mathsf{X}_{t+1}=F(\mathsf{Y}_{t}+\mathsf{E},\mathsf{B}^{\star}),\qquad\Delta_{t+1}=\mathbb{E}\left[\left(\mathsf{X}_{1},\ldots,\mathsf{X}_{t+1}\right)\left(\mathsf{X}_{1},\ldots,\mathsf{X}_{t+1}\right)^{\top}\right]\in\mathbb{R}^{(t+1)\times(t+1)}.$
Then for each $t\geq 1$, $\Delta_{t}\succ 0$ strictly, $\tau_{**}=\mathbb{E}\mathsf{X}_{t}^{2}$, and $\kappa_{*}\tau_{**}=\mathbb{E}\mathsf{Y}_{t}^{2}$.
Furthermore, let $\mathbf{X}_{t}=\left({\mathbf{x}}^{1},\ldots,\mathbf{x}^{t}\right)\in\mathbb{R}^{p\times t}$, $\mathbf{S}_{t}=\left(\mathbf{s}^{1},\ldots,\mathbf{s}^{t}\right)\in\mathbb{R}^{p\times t}$, and $\mathbf{Y}_{t}=\left({\mathbf{y}}^{1},\ldots,\mathbf{y}^{t}\right)\in\mathbb{R}^{p\times t}$ collect the iterates of (64), starting from the initialization ${\mathbf{x}}^{1}=F(\mathbf{q}^{0},\bm{\beta}^{\star})$. Then for any fixed $t\geq 1$, almost surely as $p,n\rightarrow\infty$,
$\left(\mathbf{H},\mathbf{X}_{t},\mathbf{S}_{t},\mathbf{Y}_{t}\right)\stackrel{{\scriptstyle W_{2}}}{{\to}}\left(\mathsf{H},\mathsf{X}_{1},\ldots,\mathsf{X}_{t},\mathsf{S}_{1},\ldots,\mathsf{S}_{t},\mathsf{Y}_{1},\ldots,\mathsf{Y}_{t}\right).$
Noting that each matrix $\Delta_{t}$ is the upper-left submatrix of $\Delta_{t+1}$, let us denote the entries of these matrices as $\Delta_{t}=(\delta_{rs})_{r,s=1}^{t}$. We also denote $\delta_{*}:=\tau_{**}$ and $\sigma_{*}^{2}:=\kappa_{*}\tau_{**}$.
Remark B.4.
In case where $\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}(x)$ is constant in $x$ (e.g. ridge penalty), the iterates converges in one iteration and the above result holds for $t\leq 1$.
Proof of the following Corollary is deferred to Section B.3.1.
Corollary B.5.
Under Assumptions 26, almost surely as $p,n\to\infty$
$\left(\hat{\mathbf{x}}_{1t},\mathbf{r}_{1t},\bm{\beta}^{\star}\right)\stackrel{{\scriptstyle W_{2}}}{{\to}}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right),\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star},\mathsf{B}^{\star}\right).$ (65)
Furthermore, almost surely as $p,n\to\infty$,
$\frac{1}{p}\left\|\mathbf{X}\mathbf{r}_{2t}-\mathbf{y}\right\|^{2}\rightarrow\tau_{**}\mathbb{E}\mathsf{D}^{2}+\delta$ (66)
$\frac{1}{p}\norm{\mathbf{y}-\mathbf{X}\hat{\mathbf{x}}_{2t}}^{2}\to\tau_{**}\cdot\mathbb{E}\frac{\mathsf{D}^{2}(\eta_{*}-\gamma_{*})^{2}}{(\mathsf{D}^{2}+\eta_{*}-\gamma_{*})^{2}}+\frac{n-p}{p}+\mathbb{E}\quantity(\frac{\eta_{*}-\gamma_{*}}{\mathsf{D}^{2}+\eta_{*}-\gamma_{*}})^{2}.$
B.1.2 Cauchy convergence of VAMP iterates
The following Proposition is analogous to [34, Proposition 2.3] and [59, Lemma B.2.] in the context of rotationally invariant spin glass and Bayesian linear regression. However, it requires observing a simple but crucial property of the R-transform (i.e. $-zR^{\prime}(z)/R(z)<1$ for all $z$ on the domain) and its interplay with the non-expansiveness of the proximal map. We defer the proof to Section B.3.1.
Proposition B.6.
Under Assumptions 26,
$\lim_{\min(s,t)\rightarrow\infty}\delta_{st}=\delta_{*}$
where $\delta_{st}=\mathbb{E}\mathsf{X}_{s}\mathsf{X}_{t}$.
We can then obtain the convergence of vector iterates for the oracle VAMP algorithm. We defer the proof to Section B.3.1.
Corollary B.7.
Under Assumptions 26, for $j=1,2$,
$\begin{gathered}\lim_{(s,t)\rightarrow\infty}\left(\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\mathbf{x}^{t}-\mathbf{x}^{s}\right\|^{2}\right)=\lim_{(s,t)\rightarrow\infty}\left(\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\mathbf{y}^{t}-\mathbf{y}^{s}\right\|^{2}\right)\\ \qquad\qquad\qquad=\lim_{(s,t)\rightarrow\infty}\left(\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\mathbf{r}_{jt}-\mathbf{r}_{js}\right\|^{2}\right)=\lim_{(s,t)\rightarrow\infty}\left(\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\hat{\mathbf{x}}_{jt}-\hat{\mathbf{x}}_{js}\right\|^{2}\right)=0\end{gathered}$ (67)
where the inner limits exist almost surely for each fixed $t$ and $s$.
B.1.3 Characterize limits of empirical distribution
Recall definition of $\mathbf{r}_{*},\mathbf{r}_{**}$ from (34). The following is a direct consequence of the Cauchy convergence of the VAMP iterates and the strong convexity in the penalized loss function. We defer the proof to Section B.3.2.
Proposition B.8.
Under Assumptions 26 with $c_{0}>0$, for $j=1,2$,
$\lim_{t\rightarrow\infty}\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\hat{\bm{\beta}}-\hat{\mathbf{x}}_{jt}\right\|_{2}^{2}=\lim_{t\rightarrow\infty}\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\mathbf{r}_{jt}-\mathbf{r}_{j*}\right\|_{2}^{2}=0.$
where the inner limits exist almost surely for each fixed $t$.
Combining Proposition B.8 and Corollary B.5 yields the proof of Theorem 5.1.
Proof of Theorem 5.1.
We prove (35) first. Fix function $\psi:\mathbb{R}^{3}\mapsto\mathbb{R}$ satisfying, for some constant $C>0$, the pseudo-Lipschitz condition
$\left|\psi(\mathbf{v})-\psi\left(\mathbf{v}^{\prime}\right)\right|\leq C\left(1+\|\mathbf{v}\|_{2}+\left\|\mathbf{v}^{\prime}\right\|_{2}\right)\left\|\mathbf{v}-\mathbf{v}^{\prime}\right\|_{2}.$
For any fixed $t$, we have
$\left|\frac{1}{p}\sum_{i=1}^{p}\psi\left(\hat{x}_{1t,i},r_{1t,i},{\beta}^{\star}_{i}\right)-\frac{1}{p}\sum_{i=1}^{p}\psi\left(\hat{{\beta}}_{i},r_{*,i},{\beta}^{\star}_{i}\right)\right|$
$\leq\frac{C}{p}\sum_{i=1}^{p}\left(\left|\hat{x}_{1t,i}-\hat{{\beta}}_{i}\right|^{2}+\left|r_{1t,i}-r_{*,i}\right|^{2}\right)^{\frac{1}{2}}$
×$\qquad\qquad\times\left(1+\sqrt{\hat{x}_{1t,i}^{2}+r_{1t,i}^{2}+\beta_{i}^{\star 2}}+\sqrt{\hat{{\beta}}_{i}^{2}+r_{*,i}^{2}+\beta_{i}^{\star 2}}\right)$
$\stackrel{{\scriptstyle(\star)}}{{\leq}}C\left(\frac{1}{p}\sum_{i=1}^{p}\left|\hat{x}_{1t,i}-\hat{{\beta}}_{i}\right|^{2}+\left|r_{1t,i}-r_{*,i}\right|^{2}\right)^{\frac{1}{2}}$
×$\qquad\qquad\times\left(\frac{1}{p}\sum_{i=1}^{p}\left(1+\sqrt{\hat{x}_{1t,i}^{2}+r_{1t,i}^{2}+\beta_{i}^{\star 2}}+\sqrt{\hat{{\beta}}_{i}^{2}+r_{*,i}^{2}+\beta_{i}^{\star 2}}\right)^{2}\right)^{\frac{1}{2}}$
$\leq C\left(\frac{1}{p}\left\|\hat{\mathbf{x}}_{1t}-\hat{\bm{\beta}}\right\|_{2}^{2}+\frac{1}{p}\left\|\hat{\mathbf{x}}_{1t}-\hat{\bm{\beta}}\right\|_{2}^{2}\right)^{\frac{1}{2}}$
×$\qquad\qquad\times\left(3+\frac{3}{p}\left(\left\|\hat{\mathbf{x}}_{1t}\right\|_{2}^{2}+3\left\|\mathbf{r}_{1t}\right\|_{2}^{2}+2\left\|\bm{\beta}^{\star}\right\|_{2}^{2}+2\left\|\mathbf{r}_{1t}-\mathbf{r}_{*}\right\|_{2}^{2}\right)\right)^{\frac{1}{2}}$
where ()($\star$)( ⋆ ) is by Cauchy-Schwarz inequality. This, along with Proposition B.8, Assumption 3, Corollary B.5 implies that
$\lim_{t\rightarrow\infty}\lim_{p\rightarrow\infty}\left|\frac{1}{p}\sum_{i=1}^{p}\psi\left(\hat{x}_{1t,i},r_{1t,i},{\beta}^{\star}_{i}\right)-\frac{1}{p}\sum_{i=1}^{p}\psi\left(\hat{{\beta}}_{i},r_{*,i},{\beta}^{\star}_{i}\right)\right|=0$ (68)
Using Corollary B.5 and Proposition A.1, we have that
$\lim_{p\rightarrow\infty}\left|\frac{1}{p}\sum_{i=1}^{p}\psi\left(\hat{x}_{1t,i},r_{1t,i},{\beta}^{\star}_{i}\right)-\mathbb{E}\psi\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right),\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star},\mathsf{B}^{\star}\right)\right|=0$ (69)
By triangle inequality, we also have
[…]
Taking $p$ and then $t$ to infinity on both sides of the above, by (68) and (69),
$\lim_{p\rightarrow\infty}\left|\frac{1}{p}\sum_{i=1}^{p}\psi\left(\hat{{\beta}}_{i},r_{*,i},{\beta}^{\star}_{i}\right)-\mathbb{E}\psi\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right),\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star},\mathsf{B}^{\star}\right)\right|=0$
where we used the fact that lhs does not depend on $t$. An application of Proposition A.1 with $\mathfrak{p}=2,k=3$ completes the proof for (35).
To see first result in (36), note that
$\bigg{|}\frac{1}{p}\left\|\mathbf{X}\mathbf{r}_{**}-\mathbf{y}\right\|^{2}$ $-\frac{1}{p}\left\|\mathbf{X}\mathbf{r}_{2t}-\mathbf{y}\right\|^{2}\bigg{|}=\bigg{|}\frac{1}{p}\left\langle\mathbf{X}\mathbf{r}_{**}-2\mathbf{y}+\mathbf{X}\mathbf{r}_{2t},\mathbf{X}\mathbf{r}_{**}-\mathbf{X}\mathbf{r}_{2t}\right\rangle\bigg{|}$
$\leq\frac{1}{p}\left\|\mathbf{X}\mathbf{r}_{**}-2\mathbf{y}+\mathbf{X}\mathbf{r}_{2t}\right\|_{2}\left\|\mathbf{X}\mathbf{r}_{**}-\mathbf{X}\mathbf{r}_{2t}\right\|_{2}$
$\leq\frac{1}{p}\left(\|\mathbf{X}\|_{\mathrm{op}}\left(\left\|\mathbf{r}_{**}-\bm{\beta}^{\star}\right\|_{2}+\left\|\mathbf{r}_{2t}-\bm{\beta}^{\star}\right\|\right)+2\|\bm{\varepsilon}\|_{2}\right)\|\mathbf{X}\|_{\mathrm{op}}\left\|\mathbf{r}_{**}-\mathbf{r}_{2t}\right\|_{2}.$
Using this inequality and $\|\mathbf{X}\|_{\mathrm{op}}=\max_{i\in[p]}\left|d_{i}\right|\rightarrow\sqrt{d_{+}}$ (cf. Assumption 2), we obtain that almost surely
$\lim_{t\rightarrow\infty}\limsup_{p\rightarrow\infty}\left|\frac{1}{p}\left\|\mathbf{X}\mathbf{r}_{**}-\mathbf{y}\right\|^{2}-\frac{1}{p}\left\|\mathbf{X}\mathbf{r}_{2t}-\mathbf{y}\right\|^{2}\right|=0.$ (70)
From triangle inequality, we have
$\bigg{|}\frac{1}{p}\left\|\mathbf{X}\mathbf{r}_{**}-\mathbf{y}\right\|^{2}$ $-\left(\tau_{**}\mathbb{E}\mathsf{D}^{2}+\delta\right)\bigg{|}$
$\leq\left|\frac{1}{p}\left\|\mathbf{X}\mathbf{r}_{2t}-\mathbf{y}\right\|^{2}-\left(\tau_{**}\mathbb{E}\mathsf{D}^{2}+\delta\right)\right|+\left|\frac{1}{p}\left\|\mathbf{X}\mathbf{r}_{**}-\mathbf{y}\right\|^{2}-\frac{1}{p}\left\|\mathbf{X}\mathbf{r}_{2t}-\mathbf{y}\right\|^{2}\right|.$
Apply limit operation $\lim_{t\rightarrow\infty}\limsup_{p\rightarrow\infty}$ on both sides. Using (70), (66) and the fact that the LHS does not depend on $t$, we have that almost surely
$\limsup_{p\rightarrow\infty}\left|\frac{1}{p}\left\|\mathbf{X}\mathbf{r}_{**}-\mathbf{y}\right\|^{2}-\left(\tau_{**}\mathbb{E}\mathsf{D}^{2}+\delta\right)\right|=0.$
The proof of the second result in (36) is analgous using Corollary B.5. This completes the proof. ∎
B.2 Prove result B: Consistent estimation
We prove existence and uniqueness of the solution to the adjustment equation (23) in Section B.2.1, show that the adjustment equation converges to a population limit in Section B.2.2, and prove Theorem 3.1 in Section B.2.3. To streamline the presentation, proofs of intermediate claims are collected in Section B.4.
B.2.1 Properties of the adjustment equation
Recall definition of function $g_{p}:(0,+\infty)\mapsto\mathbb{R}$ from (22). We outline in Lemma B.9 the conditions under which it is well-defined, strictly increasing and the equation
$g_{p}(\gamma)=1$ (71)
admits a unique solution on $(0,+\infty)$. The proof is deferred to Section B.4.1.
Lemma B.9.
Fix $p\geq 1$. Assume that $h^{\prime\prime}(\hat{\beta_{j}})\geq 0$ for all $j\in[p]$. We then have the following statements:
  • (a)
    If $d_{i}\neq 0$ for all $i$, the function $\gamma\mapsto g_{p}(\gamma)$ is well-defined. If for some $i\in[p],d_{i}=0$, the function $\gamma\mapsto g_{p}(\gamma)$ is well-defined if and only if $\norm{h^{\prime\prime}(\hat{\bm{\beta}})}_{0}>0$.
  • (b)
    Given that $g_{p}$ is well-defined, it is strictly increasing if there exists some $j\in[p]$ such that $h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq+\infty$, or else $g_{p}(\gamma)=1,\forall\gamma\in(0,+\infty)$.
  • (c)
    Given that $\left\|h^{\prime\prime}(\hat{\bm{\beta}})\right\|_{0}=p$ or for all $i,d_{i}\neq 0$, by which $g_{p}$ is well-defined from (a), (71) has a unique solution if and only if there exists some $j\in[p]$ such that $h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq+\infty$.
  • (d)
    Given that $\left\|h^{\prime\prime}(\hat{\bm{\beta}})\right\|_{0}<p$ and for some $i$, $d_{i}=0$, $g_{p}$ is well-defined and (71) has a unique solution on $(0,+\infty)$ if and only if $\norm{d}_{0}+\norm{h^{\prime\prime}(\hat{\bm{\beta}})}_{0}>p$.
The following assumption is made to simplify the conditions outlined in Lemma B.9.
Assumption 9.
Fix $p\geq 1$ and suppose that Assumption 4 holds. If $\left\|h^{\prime\prime}(\hat{\bm{\beta}})\right\|_{0}=p$ or that $\mathbf{X}^{\top}\mathbf{X}$ is non-singular, we require only that there exists some $i\in[p]$ such that $h^{\prime\prime}(\hat{{\beta}}_{i})\neq+\infty$. Otherwise, we require in addition that $\norm{d}_{0}+\norm{h^{\prime\prime}(\hat{\bm{\beta}})}_{0}>p$.
The following is a direct consequence of Lemma B.9 which in turn has Proposition 3.1 as a special case.
Proposition B.10.
Fix $p\geq 1$ and suppose that Assumption 4 holds. Then, Assumption 9 holds if and only if the function $\gamma\mapsto g_{p}(\gamma)$ is well-defined for any $\gamma>0$, strictly increasing, and the equation (71) admits a unique solution contained in $(0,+\infty).$
B.2.2 Population limit of the adjustment equation
From now on, we use notation for the following random variable
$\mathsf{U}:=h^{\prime\prime}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)\right).$
Define $g_{\infty}:(0,+\infty)\mapsto\mathbb{R}$ by
$g_{\infty}(\gamma)=\mathbb{E}\frac{1}{\left(\mathsf{D}^{2}-\gamma\right)\mathbb{E}\frac{1}{\gamma+\mathsf{U}}+1}.$
which is well-defined under Assumption 4, 6 as shown in Lemma B.11 below. We defer its proof to Section B.4.2.
Lemma B.11.
Under Assumption 4, 6, $g_{\infty}$ is well-defined on and strictly increasing on $(0,+\infty)$. The equation $g_{\infty}(\gamma)=1$ admits a unique solution $\gamma_{*}$ on $(0,+\infty)$.
Remark B.12.
We emphasize that the proof of Lemma B.11 does not require (20) admits a unique solution, only that a solution exists.
We can show that the LHS of the sample adjustment equation converges to the LHS of the population adjustment equation. We defer its proof to Section B.4.2.
Proposition B.13.
Under Assumption 26, almost surely for all sufficiently large $p$, $g_{p}$ is well-defined on $(0,+\infty)$ and for any $\gamma>0$, almost surely
$\lim_{p\to\infty}g_{p}(\gamma)=g_{\infty}(\gamma).$ (72)
B.2.3 Consistent estimation of fixed points
We are now ready to prove Theorem 5.2 which shows that the quantities defined in (26) indeed converges to their population counterparts.
Proof of Theorem 5.2.
We first show that $\lim_{p\rightarrow\infty}\widehat{\mathsf{adj}}\left(p\right)\rightarrow\gamma_{*}$ almost surely. Fix any $0<\epsilon<\gamma_{*}$. Note that almost surely
$\lim_{p\rightarrow\infty}g_{p}\left(\gamma_{*}-\epsilon\right)=g_{\infty}\left(\gamma_{*}-\epsilon\right)<g_{\infty}\left(\gamma_{*}\right)=1,$
$\lim_{p\rightarrow\infty}g_{p}\left(\gamma_{*}+\epsilon\right)=g_{\infty}\left(\gamma_{*}+\epsilon\right)>g_{\infty}\left(\gamma_{*}\right)=1$
as a direct consequence of Proposition B.13 and that $g_{\infty}$ is strictly increasing (cf. Lemma B.11). It follows that almost surely for all $m$ sufficiently large
$g_{p}\left(\gamma_{*}-\epsilon\right)<1,\quad g_{p}\left(\gamma_{*}+\epsilon\right)>1.$ (73)
By the assumptions and Lemma B.11, $g_{p}$ is continuous, strictly increasing and (71) admits a unique solution $\widehat{\mathsf{adj}}\left(p\right)$ for all $p$. (73) thus implies that almost surely for all $p$ sufficiently large $\left|\widehat{\mathsf{adj}}\left(p\right)-\gamma_{*}\right|<\epsilon$. This completes the proof for $\lim_{p\rightarrow\infty}\widehat{\mathsf{adj}}\left(p\right)\rightarrow\gamma_{*}$. The consistency of $\hat{\eta}_{*}$ immediately follows. To show $\hat{\bm{\beta}}^{u}(p)\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{B}^{\star}+\sqrt{\tau_{*}}\mathsf{Z}$ almost surely as $p\to\infty$, note that $\hat{\bm{\beta}}^{u}(p)-\mathbf{r}_{*}\stackrel{{\scriptstyle W_{2}}}{{\to}}0$ by consistency of $\widehat{\mathsf{adj}}$ and $\limsup_{p\to\infty}p^{-1}\left\|\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})\right\|_{2}^{2}<+\infty$, and the claims follow from (35) and an application of Proposition A.4. A similar argument shows that $\frac{1}{p}\norm{\hat{\mathbf{{r}}}_{**}-\mathbf{r}_{**}}^{2}\rightarrow 0$ almost surely as $p\to\infty$. The consistency statements for $\hat{\sigma}^{2},\hat{\tau}_{**},\hat{\tau}_{*}$ follow from results above, (36), (20d) and Assumption 7. ∎
B.3 Supporting proofs for result A
B.3.1 Oracle VAMP proofs
Proof of Proposition B.1.
By Assumption 2, and Assumption 6, (20c),
$\lim_{p\rightarrow\infty}\frac{1}{p}\operatorname{Tr}(\bm{\Lambda})=\mathbb{E}\left(\eta_{*}-\gamma_{*}\right)\left(\frac{\eta_{*}}{\gamma_{*}\left(\mathsf{D}^{2}+\left(\eta_{*}-\gamma_{*}\right)\right)}-\frac{1}{\gamma_{*}}\right)=0.$
The limiting values of $\kappa_{*}:=\lim_{p\rightarrow\infty}\frac{1}{p}\operatorname{Tr}\left(\bm{\Lambda}^{2}\right)$ and $b_{*}:=\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\mathbf{e}_{b}\right\|^{2}$ is found analogously under Assumption 2. The identity $\tau_{*}=b_{*}+\kappa_{*}\tau_{**}$ is obtained by rewriting (20d) using definitions of $b_{*},\kappa_{*}$. Using (20a), we have that
$\mathbb{E}F^{\prime}\left(\sqrt{\tau_{*}}\mathsf{Z},\mathsf{B}^{\star}\right)=\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\left(\mathbb{E}\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}\left(\mathsf{B}^{\star}+\sqrt{\tau_{*}}\mathsf{Z}\right)-\frac{\gamma_{*}}{\eta_{*}}\right)=0.$
The Lipschitz continuity of $(q,x)\mapsto F(q,x)$ on $\mathbb{R}$ follows from Proposition A.7, (b). To show $\mathbb{E}F\left(\sqrt{\tau_{*}}\mathsf{Z},\mathsf{B}^{\star}\right)^{2}=\tau_{**}$, note that
$\mathbb{E}F\left(\sqrt{\tau_{*}}\mathsf{Z},\mathsf{B}^{\star}\right)^{2}=\mathbb{E}\left(\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)-\mathsf{B}^{\star}\right)-\frac{\gamma_{*}}{\eta_{*}-\gamma_{*}}\sqrt{\tau_{*}}\mathsf{Z}\right)^{2}$
$=\left(\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\right)^{2}\mathbb{E}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)-\mathsf{B}^{\star}\right)^{2}+\left(\frac{\gamma_{*}}{\eta_{*}-\gamma_{*}}\right)^{2}\tau_{*}$
$\quad-2\frac{\gamma_{*}}{\eta_{*}-\gamma_{*}}\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\mathbb{E}\left(\sqrt{\tau_{*}}\mathsf{Z}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)-\mathsf{B}^{\star}\right)\right)$
=$\stackrel{{\scriptstyle(a)}}{{=}}\left(\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\right)^{2}\mathbb{E}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)-\mathsf{B}^{\star}\right)^{2}$
$\qquad\qquad+\left(\frac{\gamma_{*}}{\eta_{*}-\gamma_{*}}\right)^{2}\tau_{*}-2\frac{\gamma_{*}}{\eta_{*}-\gamma_{*}}\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\frac{\gamma_{*}}{\eta_{*}}\tau_{*}$
$=\left(\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\right)^{2}\mathbb{E}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)-\mathsf{B}^{\star}\right)^{2}-\left(\frac{\gamma_{*}}{\eta_{*}-\gamma_{*}}\right)^{2}\tau_{*}$
=$\stackrel{{\scriptstyle(b)}}{{=}}\tau_{**}$
where in $(a)$ we used Stein’s lemma and (20a) for the following
$\mathbb{E}\left(\mathsf{Z}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)-\mathsf{B}^{\star}\right)\right)=\mathbb{E}\left(\operatorname{Prox}_{\gamma^{-1}h}^{\prime}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)\right)=\frac{\gamma_{*}}{\eta_{*}}$
and in (b) we used (20c). We remark that although the function $x\mapsto\operatorname{Prox}_{\gamma_{*}^{-1}h}(x)$ may not be differentiable on a finite set of points, Stein’s lemma can still be applied (cf. [79, Lemma 1]). ∎
Proof of Proposition B.2.
Note that $\bm{\xi}=\mathbf{Q}\bm{\varepsilon}\sim N(\rm{0},\mathbf{I}_{n})$. Then $\mathbf{D}^{\top}\bm{\xi}\in\mathbb{R}^{n}$ may be written as the entrywise product of $\mathbf{D}^{\top}\bm{1}_{n\times 1}\in\mathbb{R}^{p}$ and a vector $\bar{\bm{\xi}}\sim N(\rm{0},\mathbf{I}_{p})$, both when $p\geq n$ and when $n\leq p$. The almost-sure convergence $H\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{H}$ is then a straightforward consequence of Propositions A.2, A.3, and A.6, where all random variables of $\mathsf{H}$ have finite moments of all orders under Assumptions 2 and 3. The identities $\kappa_{*}=\mathbb{E}\mathsf{L}^{2}$ and $b_{*}=\mathbb{E}\mathsf{E}_{b}^{2}$ follows from definitions of $\kappa_{*},b_{*}$ in Proposition B.1. ∎
Proof of Proposition B.3.
We have $\delta_{11}=\mathbb{E}\mathsf{X}_{1}^{2}=\delta_{*}$ by the last identity of (62c). Supposing that $\delta_{tt}=\mathbb{E}\mathsf{X}_{t}^{2}=\delta_{*}$, we have by definition $\mathbb{E}\mathsf{Y}_{t}^{2}=\kappa_{*}\delta_{tt}=\sigma_{*}^{2}=\delta_{*}\kappa_{*}$. Since $\mathsf{Y}_{t}$ is independent of $\mathsf{E}$, we have $\mathsf{Y}_{t}+\mathsf{E}\sim N(0,\sigma_{*}^{2}+b_{*})$ where this variance is $\sigma_{*}^{2}+b_{*}=\tau_{*}$ by last identity of (62b). Then $\mathbb{E}\mathsf{X}_{t+1}^{2}=\delta_{*}$ by the last identity of (62c), so $\mathbb{E}\mathsf{X}_{t}^{2}=\delta_{*}$ and $\mathbb{E}\mathsf{Y}_{t}^{2}=\sigma_{*}^{2}$ for all $t\geq 1$.
Noting that $\Delta_{t}$ is the upper-left submatrix of $\Delta_{t+1}$, let us denote
$\Delta_{t+1}=\begin{pmatrix}\Delta_{t}&\delta_{t}\\ \delta_{t}^{\top}&\delta_{*}\end{pmatrix}$
We now show by induction on $t$ the following three statements:
  1. 1.
    $\Delta_{t}\succ 0$ strictly.
  2. 2.
    We have
    $\mathsf{Y}_{t}=\sum_{k=1}^{t-1}\mathsf{Y}_{k}\left(\Delta_{t-1}^{-1}\delta_{t-1}\right)_{k}+\mathsf{U}_{t},\quad\mathsf{S}_{t}=\sum_{k=1}^{t-1}\mathsf{S}_{k}\left(\Delta_{t-1}^{-1}\delta_{t-1}\right)_{k}+\mathsf{U}^{\prime}_{t}$ (74)
    where $\mathsf{U}_{t},\mathsf{U}_{t}^{\prime}$ are Gaussian variables with strictly positive variance, independent of $\mathsf{H}$, $\left(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{t-1}\right)$, and $\left(\mathsf{S}_{1},\ldots,\mathsf{S}_{t-1}\right)$.
  3. 3.
    $\left(\mathbf{H},\mathbf{X}_{t+1},\mathbf{S}_{t},\mathbf{Y}_{t}\right)\stackrel{{\scriptstyle W_{2}}}{{\to}}\left(\mathsf{H},\mathsf{X}_{1},\ldots,\mathsf{X}_{t+1},\mathsf{S}_{1},\ldots,\mathsf{S}_{t},\mathsf{Y}_{1},\ldots,\mathsf{Y}_{t}\right)$.
We take as base case $t=0$, where the first two statements are vacuous, and the third statement requires $(\mathbf{H},{\mathbf{x}}^{1})\stackrel{{\scriptstyle W_{2}}}{{\to}}(\mathsf{H},\mathsf{X}_{1})$ almost surely as $p\to\infty$. Recall that ${\mathbf{x}}^{1}=F(p^{0},\bm{\beta}^{\star})$, and that $F(p,\beta)$ is Lipschitz by Proposition Proposition B.1. Then this third statement follows from Propositions B.2 and A.3.
Supposing that these statements hold for some $t\geq 0$, we now show that they hold for $t+1$. To show the first statement $\Delta_{t+1}\succ 0$, note that for $t=0$ this follows from $\Delta_{1}=\delta_{*}>0$ by Assumption 6. For $t\geq 1$, given that $\Delta_{t}\succ 0$, $\Delta_{t+1}$ is singular if and only if there exist constants $\alpha_{1},\ldots,\alpha_{t}\in\mathbb{R}$ such that
$\mathsf{X}_{t+1}=F\left(\mathsf{Y}_{t}+\mathsf{E},\mathsf{B}^{\star}\right)=\sum_{r=1}^{t}\alpha_{r}\mathsf{X}_{r}$
with probability 1. From the induction hypothesis, $\mathsf{Y}_{t}=\sum_{k=1}^{t-1}\mathsf{Y}_{k}\left(\Delta_{r}^{-1}\delta_{r}\right)_{k}+\mathsf{U}_{t}$ where $\mathsf{U}_{t}$ is independent of $\mathsf{H},\mathsf{Y}_{1},\ldots,\mathsf{Y}_{t-1}$ and hence also of $\mathsf{E},\mathsf{B}^{\star},\mathsf{X}_{1},...,\mathsf{X}_{t}$. We now show that for any realized values $(e_{0},x_{0},w_{0})$ of
$\left(\mathsf{E}+\sum_{k=1}^{t-1}\mathsf{Y}_{k}\left(\Delta_{r}^{-1}\delta_{r}\right)_{k},\quad\mathsf{B}^{\star},\quad\sum_{r=1}^{t}\alpha_{r}\mathsf{X}_{r}\right),$
we have that $\mathbb{P}\left(F\left(\mathsf{U}_{t}+e_{0},x_{0}\right)\neq w_{0}\right)>0$. This would imply that $\Delta_{t+1}\succ 0$. Suppose to the contrary, we then have that
$\mathbb{P}\left(\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\operatorname{Prox}_{\gamma_{\gamma_{*}^{-1}}}\left(\mathsf{U}_{t}+e_{0}+x_{0}\right)-\frac{\gamma_{*}}{\eta_{*}-\gamma_{*}}\mathsf{U}_{t}=w_{0}+\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}x_{0}+\frac{\gamma_{*}}{\eta_{*}-\gamma_{*}}e_{0}\right)=1.$
Since $\mathsf{U}_{t}$ is Gaussian with strictly positive variance, the above implies that the function
$u\mapsto\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(u+e_{0}+x_{0}\right)-\frac{\gamma_{*}}{\eta_{*}-\gamma_{*}}u$
is constant almost everywhere. This in turn is equivalent to that $\operatorname{Prox}_{\gamma_{*}^{-1}h}(u)=C+\frac{\gamma_{*}}{\eta_{*}}u$ almost everywhere for some constant $C\in\mathbb{R}$ by a change of variable. Noting that $u\mapsto\operatorname{Prox}_{\gamma_{*}^{-1}h}(u)$ is continuous, we thus have that $\operatorname{Prox}_{\gamma_{*}^{-1}h}(u)=C+\frac{\gamma_{*}}{\eta_{*}}u$ for all $u\in\mathbb{R}$. This implies that $\operatorname{Prox}_{\gamma_{*}^{-1}h}(u)$ is continuously differentiable and has constant derivative $\frac{\gamma_{*}}{\eta_{*}}$, which contradicts to the assumption that $x\mapsto\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}(x)$ is non-constant. We thus have proved the first inductive statement that $\Delta_{t+1}\succ 0$.
To study the empirical limit of $s_{t+1}$, let $\mathbf{U}=\left(\mathbf{e}_{b},\mathbf{S}_{t},\bm{\Lambda}\mathbf{S}_{t}\right)$ and $\mathbf{V}=\left(\mathbf{e},\mathbf{X}_{t},\mathbf{Y}_{t}\right)$. (For $t=0$, this is simply $\mathbf{U}=\mathbf{e}_{b}$ and $\mathbf{V}=\mathbf{e}$.) By the induction hypothesis, the independence of $(\mathsf{S}_{1},\ldots,\mathsf{S}_{t})$ with $(\mathsf{E}_{b},\mathsf{L})$, and the identities $\mathbb{E}\mathsf{E}_{b}^{2}=b_{*}$ and $\mathbb{E}\mathsf{L}=0$ and $\mathbb{E}\mathsf{L}^{2}=\kappa_{*}$, almost surely as $p\to\infty$,
$\frac{1}{p}\left(\mathbf{e}_{b},\mathbf{S}_{t},\bm{\Lambda}\mathbf{S}_{t}\right)^{\top}\left(\mathbf{e}_{b},\mathbf{S}_{t},\bm{\Lambda}\mathbf{S}_{t}\right)\rightarrow\left(\begin{array}[]{ccc}b_{*}&0&0\\ 0&\Delta_{t}&0\\ 0&0&\kappa_{*}\Delta_{t}\end{array}\right)\succ 0$
So almost surely for sufficiently large $p$, conditional on $(\mathbf{H},\mathbf{X}_{t+1},\mathbf{S}_{t},\mathbf{Y}_{t})$, the law of $\mathbf{s}^{t+1}$ is given by its law conditioned on $\mathbf{U}=\mathbf{O}V$, which is (see [33, Lemma B.2])
$\mathbf{s}^{t+1}\big{|}_{\mathbf{U}=\mathbf{O}\mathbf{V}}=\mathbf{O}\mathbf{x}^{t+1}\big{|}_{\mathbf{U}=\mathbf{O}\mathbf{V}}\stackrel{{\scriptstyle L}}{{=}}\mathbf{U}\left(\mathbf{U}^{\top}\mathbf{U}\right)^{-1}\mathbf{V}^{\top}\mathbf{x}^{t+1}+\bm{\Pi}_{\mathbf{U}^{\perp}}\tilde{\mathbf{O}}\bm{\Pi}_{\mathbf{V}^{\perp}}^{\top}\mathbf{x}^{t+1}$ (75)
where $\tilde{\mathbf{O}}\sim\operatorname{Haar}(\mathbb{O}(p-(2t+1)))$ and $\bm{\Pi}_{\mathbf{U}^{\perp}},\bm{\Pi}_{\mathbf{V}^{\perp}}\in\mathbb{R}^{p\times(p-(2t+1))}$ are matrices with orthonormal columns spanning the orthogonal complements of the column spans of $\mathbf{U},\mathbf{V}$ respectively. We may replace $\mathbf{s}^{t+1}$ by the right side of (75) without affecting the joint law of $\left(\mathbf{H},\mathbf{X}_{t+1},\mathbf{S}_{t},\mathbf{Y}_{t},\mathbf{s}^{t+1}\right)$.
For $t=0$, we have $\mathbb{E}\mathsf{X}_{1}\mathsf{E}=0$ since $\mathsf{X}_{1}$ is independent of $\mathsf{E}$. For $t\geq 1$, by the definition of $\mathsf{X}_{t+1}$, the condition $\mathbb{E}F^{\prime}(\mathsf{P},\mathsf{B}^{\star})=0$ from (62c), and Stein’s lemma, we have $\mathbb{E}\mathsf{X}_{t+1}\mathsf{E}=0$ and $\mathbb{E}\mathsf{X}_{t+1}\mathsf{Y}_{r}=0$ for each $r=1,\ldots,t$. Then by the induction hypothesis, almost surely as $p\rightarrow\infty$,
$\left(p^{-1}\mathbf{U}^{\top}\mathbf{U}\right)^{-1}\rightarrow\left(\begin{array}[]{ccc}b_{*}&0&0\\ 0&\Delta_{t}&0\\ 0&0&\kappa_{*}\Delta_{t}\end{array}\right)^{-1},\quad p^{-1}\mathbf{V}^{\top}\mathbf{x}^{t+1}\rightarrow\left(\begin{array}[]{c}0\\ \delta_{t}\\ 0\end{array}\right).$
Then by (75) and Propositions A.4 and A.6, it follows that
$\left(\mathbf{H},\mathbf{X}_{t+1},\mathbf{S}_{t},\mathbf{Y}_{t},\mathbf{s}^{t+1}\right)$
$\qquad\qquad\stackrel{{\scriptstyle W_{2}}}{{\to}}\left(\mathsf{H},\mathsf{X}_{1},\ldots,\mathsf{X}_{t+1},\mathsf{S}_{1},\ldots,\mathsf{S}_{t},\mathsf{Y}_{1},\ldots\mathsf{Y}_{t},\sum_{r=1}^{t}\mathsf{S}_{r}\left(\Delta_{t}^{-1}\delta_{t}\right)_{r}+\mathsf{U}^{\prime}_{t+1}\right)$
where $\mathsf{U}^{\prime}_{t+1}$ is the Gaussian limit of the second term on the right side of (75) and is independent of $\mathsf{H},\mathsf{X}_{1},\ldots,\mathsf{X}_{t+1},\mathsf{S}_{1},\ldots,\mathsf{S}_{t},\mathsf{Y}_{1},\ldots\mathsf{Y}_{t}$. We can thus set $\mathsf{S}_{t+1}:=\sum_{r=1}^{t}\mathsf{S}_{r}\left(\Delta_{t}^{-1}\delta_{t}\right)_{r}+\mathsf{U}^{\prime}_{t+1}$. Then $(\mathsf{S}_{1},\ldots,\mathsf{S}_{t+1})$ is multivariate Gaussian and remains independent of $\mathsf{H}$ and $(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{t})$. Since $p^{-1}\|\mathbf{s}^{t+1}\|^{2}=p^{-1}\|\mathbf{x}^{t+1}\|^{2}\rightarrow\delta_{*}$ almost surely as $p\rightarrow\infty$ by the induction hypothesis, we have $\mathbb{E}\mathsf{S}_{t+1}^{2}=\delta_{*}$. From the form of $\mathsf{S}_{t+1}$, we may check also $\mathbb{E}\mathsf{S}_{t+1}(\mathsf{S}_{1},\ldots,\mathsf{S}_{t})=\delta_{t}$, so $(\mathsf{S}_{1},\ldots,\mathsf{S}_{t+1})$ has covariance $\Delta_{t+1}$ as desired. Furthermore $\sum_{r=1}^{t}\mathsf{~{}S}_{r}\left(\Delta_{t}^{-1}\delta_{t}\right)_{r}\sim N\left(0,\delta_{t}^{\top}\Delta_{t}^{-1}\delta_{t}\right)$. From $\Delta_{t+1}\succ 0$ and the Schur complement formula, $\delta_{*}-\delta_{t}^{\top}\Delta_{t}^{-1}\delta_{t}>0$ strictly. Then $\mathsf{U}^{\prime}_{t+1}$ has strictly positive variance, since the variance of $\sum_{r=1}^{t}\mathsf{S}_{r}\left(\Delta_{t}^{-1}\delta_{t}\right)_{r}$ is less than the variance of $\mathsf{S}_{t+1}$. This proves the second equation in (74) for $t+1$.
Now, we study the empirical limit of $\mathbf{y}^{t+1}$. Let $\mathbf{U}=\left(\mathbf{e},\mathbf{X}_{t+1},\mathbf{Y}_{t}\right)$, $\mathbf{V}=\left(\mathbf{e}_{b},\mathbf{S}_{t+1},\bm{\Lambda}\mathbf{S}_{t}\right)$. Similarly by the induction hypothesis and the empirical convergence of $(\mathbf{H},\mathbf{S}_{t+1})$ already shown, almost surely as $p\rightarrow\infty$,
$\frac{1}{p}\left(\mathbf{e}_{b},\mathbf{S}_{t+1},\bm{\Lambda}\mathbf{S}_{t}\right)^{\top}\left(\mathbf{e}_{b},\mathbf{S}_{t+1},\bm{\Lambda}\mathbf{S}_{t}\right)\rightarrow\left(\begin{array}[]{ccc}b_{*}&0&0\\ 0&\Delta_{t+1}&0\\ 0&0&\kappa_{*}\Delta_{t}\end{array}\right)\succ 0.$
Then the law of $\mathbf{y}^{t+1}$ conditional on $(\mathbf{H},\mathbf{X}_{t+1},\mathbf{S}_{t+1},\mathbf{Y}_{t})$ is given by its law conditioned on $\mathbf{U}=\mathbf{O}^{\top}\mathbf{V}$, which is
$\mathbf{y}^{t+1}\big{|}_{\mathbf{U}=\mathbf{O}^{\top}\mathbf{V}}=\mathbf{O}^{\top}\bm{\Lambda}\mathbf{s}^{t+1}\big{|}_{\mathbf{U}=\mathbf{O}^{\top}\mathbf{V}}\stackrel{{\scriptstyle L}}{{=}}\mathbf{U}\left(\mathbf{V}^{\top}\mathbf{V}\right)^{-1}\mathbf{V}^{\top}\bm{\Lambda}\mathbf{s}^{t+1}+\bm{\Pi}_{\mathbf{U}^{\perp}}\tilde{\mathbf{O}}\bm{\Pi}_{\mathbf{V}^{\perp}}^{\top}\bm{\Lambda}\mathbf{s}^{t+1}$ (76)
where $\tilde{\mathbf{O}}\sim\operatorname{Haar}(\mathbb{O}(p-(2t+2)))$. From the convergence of $(\mathbf{H},\mathbf{S}_{t+1})$ already shown, almost surely as $p\rightarrow\infty$,
$\left(n^{-1}\mathbf{V}^{\top}\mathbf{V}\right)^{-1}\rightarrow\left(\begin{array}[]{ccc}b_{*}&0&0\\ 0&\Delta_{t+1}&0\\ 0&0&\kappa_{*}\Delta_{t}\end{array}\right)^{-1},\quad n^{-1}\mathbf{V}^{\top}\bm{\Lambda}\mathbf{s}^{t+1}\rightarrow\left(\begin{array}[]{c}0\\ 0\\ \kappa_{*}\delta_{t}\end{array}\right).$
Then by (76) and Propositions A.4 and A.6,
$\left(\mathbf{H},\mathbf{X}_{t+1},\mathbf{S}_{t+1},\mathbf{Y}_{t},\mathbf{y}^{t+1}\right)$
$\qquad\qquad\stackrel{{\scriptstyle W_{2}}}{{\to}}\left(\mathsf{H},\mathsf{X}_{1},\ldots,\mathsf{X}_{t+1},\mathsf{~{}S}_{1},\ldots,\mathsf{S}_{t+1},\mathsf{Y}_{1},\ldots\mathsf{Y}_{t},\sum_{r=1}^{t}\mathsf{Y}_{r}\left(\Delta_{t}^{-1}\delta_{t}\right)_{r}+\mathsf{U}_{t+1}\right)$
where $\mathsf{U}_{t+1}$ is the limit of the second term on the right side of (76), which is Gaussian and independent of $\mathsf{H},\mathsf{S}_{1},\ldots,\mathsf{S}_{t+1},\mathsf{Y}_{1},\ldots\mathsf{Y}_{t}$. Setting $\mathsf{Y}_{t+1}:=\sum_{r=1}^{t}\mathsf{Y}_{r}\left(\Delta_{t}^{-1}\delta_{t}\right)_{r}+\mathsf{U}_{t+1}$, it follows that $(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{t+1})$ remains independent of $\mathsf{H}$ and $(\mathsf{S}_{1},\ldots,\mathsf{S}_{t+1})$. We may check that $\mathbb{E}\mathsf{Y}_{t+1}(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{t})=\kappa_{*}\delta_{t}$, and we have also $n^{-1}\|\mathbf{y}^{t+1}\|^{2}=n^{-1}\|\bm{\Lambda}\mathbf{s}^{t+1}\|^{2}\rightarrow\kappa_{*}\delta_{*}$ so $\mathbb{E}\mathsf{Y}_{t+1}^{2}=\kappa_{*}\delta_{*}$. From $\Delta_{t+1}\succ 0$ and the Schur complement formula, note that $\sum_{r=1}^{t}\mathsf{Y}_{r}\left(\Delta_{t}^{-1}\delta_{t}\right)_{r}$ has variance $\kappa_{*}\delta_{t}^{\top}\Delta_{t}^{-1}\delta_{t}$ which is strictly smaller than $\kappa_{*}\delta_{*}$, so $\mathsf{U}_{t+1}$ has strictly positive variance. This proves the first equation in (74) for $t+1$, and completes the proof of this second inductive statement.
Finally, recall $\mathbf{x}^{t+2}=F\left(\mathbf{y}^{t+1}+\mathbf{e},\bm{\beta}^{\star}\right)$ where $F$ is Lipschitz. Then by Proposition A.3, almost surely
$\left(\mathbf{H},\mathbf{X}_{t+2},\mathbf{S}_{t+1},\mathbf{Y}_{t+1}\right)\stackrel{{\scriptstyle W_{2}}}{{\to}}\left(\mathsf{H},\mathsf{X}_{1},\ldots,\mathsf{X}_{t+2},\mathsf{~{}S}_{1},\ldots,\mathsf{S}_{t+1},\mathsf{Y}_{1},\ldots,\mathsf{Y}_{t+1}\right)$
where $\mathsf{X}_{t+2}=F\left(\mathsf{Y}_{t+1}+\mathsf{E},\mathsf{B}^{\star}\right)$, showing the third inductive statement and completing the induction. ∎
Proof of Corollary B.5.
(65) is a direct consequence of Proposition B.3, (63), Proposition A.3, Proposition A.4, (59a) and the fact that proximal map is 1-Lipschitz. To see the first result in (66), note that
$\mathbf{X}\mathbf{r}_{2t}-\mathbf{y}=\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}\left(\mathbf{r}_{2t}-\bm{\beta}^{\star}\right)-\bm{\varepsilon}=\mathbf{Q}^{\top}\mathbf{D}\mathbf{s}^{t}-\bm{\varepsilon}$
and thus almost surely
$\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\mathbf{X}\mathbf{r}_{2t}-\mathbf{y}\right\|^{2}=\lim_{p\rightarrow\infty}\frac{1}{p}(\mathbf{s}^{t})^{\top}\mathbf{D}^{\top}\mathbf{D}\mathbf{s}^{t}+\frac{1}{p}\|\bm{\varepsilon}\|_{2}^{2}-\frac{2}{p}(\mathbf{s}^{t})^{\top}\mathbf{D}^{\top}\mathbf{Q}\bm{\varepsilon}=\tau_{**}\mathbb{E}\mathsf{D}^{2}+\delta.$
To see the second result in (66), we first note the identities
$\hat{\mathbf{x}}_{2t}-\bm{\beta}^{\star}=\mathbf{O}^{\top}\left(\mathbf{D}^{\top}\mathbf{D}+\left(\eta_{*}-\gamma_{*}\right)\cdot\mathbf{I}_{p}\right)^{-1}\left[\mathbf{D}^{\top}\mathbf{Q}\bm{\varepsilon}+\left(\eta_{*}-\gamma_{*}\right)\cdot\mathbf{s}^{t}\right]$ (77)
We also have that
$\frac{1}{p}\norm{\mathbf{y}-\mathbf{X}\hat{\mathbf{x}}_{2t}}^{2}=\frac{1}{p}\norm{\bm{\varepsilon}}^{2}$ $+\frac{1}{p}\quantity(\hat{\mathbf{x}}_{2t}-\bm{\beta}^{\star})\mathbf{O}^{\top}\mathbf{D}^{\top}\mathbf{D}\mathbf{O}(\hat{\mathbf{x}}_{2t}-\bm{\beta}^{\star})$ (78)
$-2\bm{\varepsilon}^{\top}\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}(\hat{\mathbf{x}}_{2t}-\bm{\beta}^{\star})$
Using (77), (78) above, Assumption 2,3, and Proposition B.3, we obtain that almost surely as $p\to\infty$
$\frac{1}{p}\norm{\mathbf{y}-\mathbf{X}\hat{\mathbf{x}}_{2t}}^{2}\to\tau_{**}\cdot\mathbb{E}\frac{\mathsf{D}^{2}(\eta_{*}-\gamma_{*})^{2}}{(\mathsf{D}^{2}+\eta_{*}-\gamma_{*})^{2}}+\frac{n-p}{p}+\mathbb{E}\quantity(\frac{\eta_{*}-\gamma_{*}}{\mathsf{D}^{2}+\eta_{*}-\gamma_{*}})^{2}$ (79)
as required. ∎
Proof of Proposition B.6.
Recall that $\delta_{tt}=\delta_{*}$ for all $t\geq 1$ from Theorem B.3. Then $\delta_{st}=\mathbb{E}\left[\mathsf{X}_{s}\mathsf{X}_{t}\right]\leq\sqrt{\mathbb{E}\left[\mathsf{X}_{s}^{2}\right]\mathbb{E}\left[\mathsf{X}_{t}^{2}\right]}=\delta_{*}$ for all $s,t\geq 1$. For $s=1$ and any $t\geq 2$, observe also that
$\begin{gathered}\delta_{1t}=\mathbb{E}\mathsf{X}_{1}\mathsf{X}_{t}=\mathbb{E}\left[F\left(\mathsf{P}_{0},\mathsf{B}^{\star}\right)F\left(\mathsf{Y}_{t-1}+\mathsf{E},\mathsf{B}^{\star}\right)\right]=\mathbb{E}\left[\mathbb{E}\left[F\left(\mathsf{P}_{0},\mathsf{B}^{\star}\right)F\left(\mathsf{Y}_{t-1}+\mathsf{E},\mathsf{B}^{\star}\right)\mid\mathsf{B}^{\star}\right]\right]\\ =\mathbb{E}[\mathbb{E}\left[F\left(\mathsf{P}_{0},\mathsf{B}^{\star}\right)\mid\mathsf{B}^{\star}\right]^{2}]\geq 0\end{gathered}$ (80)
where the last equality holds because $\mathsf{P}_{0}$, $\mathsf{Y}_{t-1}+\mathsf{E}$, and $\mathsf{B}^{\star}$ are independent, with $\mathsf{P}_{0}$ and $\mathsf{Y}_{t-1}+\mathsf{E}$ equal in law (by the identity $\sigma_{*}^{2}+b_{*}=\tau_{*}$). Consider now the map $\delta_{st}\mapsto\delta_{s+1,t+1}$. Recalling that $\mathbb{E}\mathsf{Y}_{t}^{2}=\sigma_{*}^{2}$ and $\mathbb{E}\mathsf{Y}_{s}\mathsf{Y}_{t}=\kappa_{*}\delta_{st}$, we may represent
$\left(\mathsf{Y}_{s}+\mathsf{E},\mathsf{Y}_{t}+\mathsf{E}\right)\stackrel{{\scriptstyle L}}{{=}}\left(\sqrt{\kappa_{*}\delta_{st}+b_{*}}\mathsf{G}+\sqrt{\sigma_{*}^{2}-\kappa_{*}\delta_{st}}\mathsf{G}^{\prime},\sqrt{\kappa_{*}\delta_{st}+b_{*}}\mathsf{G}+\sqrt{\sigma_{*}^{2}-\kappa_{*}\delta_{st}}\mathsf{G}^{\prime\prime}\right)$
where $\mathsf{G},\mathsf{G}^{\prime},\mathsf{G}^{\prime\prime}$ are jointly independent standard Gaussian variables. Denote
$\mathsf{P}_{\delta}^{\prime}:=\sqrt{\kappa_{*}\delta+b_{*}}\cdot\mathsf{G}+\sqrt{\sigma_{*}^{2}-\kappa_{*}\delta}\cdot\mathsf{G}^{\prime},\quad\mathsf{P}_{\delta}^{\prime\prime}:=\sqrt{\kappa_{*}\delta+b_{*}}\cdot\mathsf{G}+\sqrt{\sigma_{*}^{2}-\kappa_{*}\delta}\cdot\mathsf{G}^{\prime\prime}$
and define $g:\left[0,\delta_{*}\right]\to\mathbb{R}$ by $g(\delta):=\mathbb{E}\left[F\left(P_{\delta}^{\prime},\mathsf{B}^{\star}\right)F\left(P_{\delta}^{\prime\prime},\mathsf{B}^{\star}\right)\right]$. Then $\delta_{s+1,t+1}=g(\delta_{st})$.
We claim that for any $\delta\in[0,\delta_{*}]$, we have $g(\delta)\geq 0$, $g^{\prime}(\delta)\geq 0$, and $g^{\prime\prime}(\delta)\geq 0$. The first bound $g(\delta)\geq 0$ follows from
$g(\delta)=\mathbb{E}\Big{[}\mathbb{E}[F\left(\mathsf{P}_{\delta}^{\prime},\mathsf{B}^{\star}\right)F\left(\mathsf{P}_{\delta}^{\prime\prime},\mathsf{B}^{\star}\right)\mid\mathsf{B}^{\star},\mathsf{G}]\Big{]}=\mathbb{E}\left[\mathbb{E}\left[F\left(\mathsf{P}_{\delta}^{\prime},\mathsf{B}^{\star}\right)\mid\mathsf{B}^{\star},\mathsf{G}\right]^{2}\right]\geq 0,$
because $\mathsf{P}_{\delta}^{\prime},\mathsf{P}_{\delta}^{\prime\prime}$ are independent and equal in law conditional on $\mathsf{G},\mathsf{B}^{\star}$. Differentiating in $\delta$ and applying Gaussian integration by parts,
$g^{\prime}(\delta)=2\mathbb{E}\left[F^{\prime}\left(\mathsf{P}_{\delta}^{\prime},\mathsf{B}^{\star}\right)F\left(\mathsf{P}_{\delta}^{\prime\prime},\mathsf{B}^{\star}\right)\left(\frac{\kappa_{*}}{2\sqrt{\kappa_{*}\delta+b_{*}}}\cdot\mathsf{G}-\frac{\kappa_{*}}{2\sqrt{\sigma_{*}^{2}-\kappa_{*}\delta}}\cdot\mathsf{G}^{\prime}\right)\right]$
$=\frac{\kappa_{*}}{\sqrt{\kappa_{*}\delta+b_{*}}}\mathbb{E}\left[F^{\prime}\left(\mathsf{P}_{\delta}^{\prime},\mathsf{B}^{\star}\right)F\left(\mathsf{P}_{\delta}^{\prime\prime},\mathsf{B}^{\star}\right)\mathsf{G}\right]-\frac{\kappa_{*}}{\sqrt{\sigma_{*}^{2}-\kappa_{*}\delta}}\mathbb{E}\left[F^{\prime}\left(\mathsf{P}_{\delta}^{\prime},\mathsf{B}^{\star}\right)F\left(\mathsf{P}_{\delta}^{\prime\prime},\mathsf{B}^{\star}\right)\mathsf{G}^{\prime}\right]$
$=\kappa_{*}\mathbb{E}\left[F^{\prime\prime}\left(\mathsf{P}_{\delta}^{\prime},\mathsf{B}^{\star}\right)F\left(\mathsf{P}_{\delta}^{\prime\prime},\mathsf{B}^{\star}\right)+F^{\prime}\left(\mathsf{P}_{\delta}^{\prime},\mathsf{B}^{\star}\right)F^{\prime}\left(\mathsf{P}_{\delta}^{\prime\prime},\mathsf{B}^{\star}\right)\right]-\kappa_{*}\mathbb{E}\left[F^{\prime\prime}\left(\mathsf{P}_{\delta}^{\prime},\mathsf{B}^{\star}\right)F\left(\mathsf{P}_{\delta}^{\prime\prime},\mathsf{B}^{\star}\right)\right]$
$=\kappa_{*}\mathbb{E}\left[F^{\prime}\left(\mathsf{P}_{\delta}^{\prime},\mathsf{B}^{\star}\right)F^{\prime}\left(\mathsf{P}_{\delta}^{\prime\prime},\mathsf{B}^{\star}\right)\right].$
Then $g^{\prime}(\delta)=\kappa_{*}\mathbb{E}\left[\mathbb{E}[F^{\prime}(\mathsf{P}_{\delta}^{\prime},\mathsf{B}^{\star})\mid\mathsf{G},\mathsf{B}^{\star}]^{2}\right]\geq 0$, and a similar argument shows $g^{\prime\prime}(\delta)\geq 0$. Observe that at $\delta=\delta_{*}$, we have $\mathsf{P}_{\delta_{*}}^{\prime}=\mathsf{P}_{\delta_{*}}^{\prime\prime}=\sqrt{\sigma_{*}^{2}+b_{*}}\cdot\mathsf{G}=\sqrt{\tau_{*}}\mathsf{G}$ which is equal in law to $\mathsf{P}\sim N(\rm{0},\tau_{*})$. Then $g(\delta_{*})=\mathbb{E}[F(\mathsf{P},\mathsf{B}^{\star})^{2}]=\delta_{*}$ by Proposition B.1. So $g:[0,\delta_{*}]\to[0,\delta_{*}]$ is a non-negative, increasing, convex function with a fixed point at $\delta_{*}$. We claim that
$g^{\prime}(\delta_{*})<1$ (81)
This then implies that $\delta_{*}$ is the unique fixed point of $g(\cdot)$ over $[0,\delta_{*}]$, and $\lim_{t\to\infty}g^{(t)}(\delta)=\delta_{*}$ for any $\delta\in[0,\delta_{*}]$. Observe from (80) that $\delta_{1t}=\delta_{12}$ for all $t\geq 2$, so $\delta_{t,t+s}=g^{(t-1)}(\delta_{1,1+s})=g^{(t-1)}(\delta_{12})$ for any $s\geq 1$. Then $\lim_{\min(s,t)\to\infty}\delta_{st}=\delta_{*}$ follows.
It remains to show (81). Using Proposition B.1,
$g^{\prime}\left(\delta_{*}\right)=\kappa_{*}\mathbb{E}[$ $\left.F^{\prime}\left(\mathsf{P}_{\delta_{*}}^{\prime},\mathsf{B}^{\star}\right)^{2}\right]=\kappa_{*}\mathbb{E}\left[\left(\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}\left(\mathsf{P}_{\delta_{*}}^{\prime}+\mathsf{B}^{\star}\right)-\frac{\gamma_{*}}{\eta_{*}-\gamma_{*}}\right)^{2}\right]$ (82)
$=\left(\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\right)^{2}\kappa_{*}\mathbb{E}\left[\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}\left(\mathsf{P}_{\delta_{*}}^{\prime}+\mathsf{B}^{\star}\right)-\frac{\gamma_{*}}{\eta_{*}}\right)^{2}\right]$
$=\left(\frac{\eta_{*}}{\gamma_{*}}\right)^{2}\left(\mathbb{E}\frac{\eta_{*}^{2}}{\left(\mathsf{D}^{2}+\eta_{*}-\gamma_{*}\right)^{2}}-1\right)\mathbb{E}\left[\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}\left(\mathsf{P}_{\delta_{*}}^{\prime}+\mathsf{B}^{\star}\right)\right)^{2}-\left(\frac{\gamma_{*}}{\eta_{*}}\right)^{2}\right].$
Using Lemma A.8 (c), we obtain that
$R^{\prime}\left(\eta_{*}^{-1}\right)=-\left(\mathbb{E}\frac{1}{\left(\mathsf{D}^{2}+\eta_{*}-\gamma_{*}\right)^{2}}\right)^{-1}+\eta_{*}^{2}\implies\frac{\eta_{*}^{2}}{\eta_{*}^{2}-R^{\prime}\left(\eta_{*}^{-1}\right)}=\mathbb{E}\frac{\eta_{*}^{2}}{\left(\mathsf{D}^{2}+\eta_{*}-\gamma_{*}\right)^{2}}.$ (83)
Note also that by Jensen’s inequality and (20c) that
$\mathbb{E}\frac{\eta_{*}^{2}}{\left(\mathsf{D}^{2}+\eta_{*}-\gamma_{*}\right)^{2}}-1\geq 0$ (84)
By Lemma A.10 and (16), we have
$\mathbb{E}\left[\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}\left(\mathsf{P}_{\delta_{*}}^{\prime}+\mathsf{B}^{\star}\right)\right)^{2}\right]<\mathbb{E}\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}\left(\mathsf{P}_{\delta_{*}}^{\prime}+\mathsf{B}^{\star}\right)=\frac{\gamma_{*}}{\eta_{*}}.$
This implies that
$0\leq\mathbb{E}\left[\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}\left(\mathsf{P}_{\delta_{*}}^{\prime}+\mathsf{B}^{\star}\right)\right)^{2}-\left(\frac{\gamma_{*}}{\eta_{*}}\right)^{2}\right]<\frac{\gamma_{*}}{\eta_{*}}-\left(\frac{\gamma_{*}}{\eta_{*}}\right)^{2}.$ (85)
Combining (82),(83),(84) and (85) above, we obtain that
$g^{\prime}\left(\delta_{*}\right)<\left(\frac{R^{\prime}\left(\eta_{*}^{-1}\right)}{\eta_{*}^{2}-R^{\prime}\left(\eta_{*}^{-1}\right)}\right)\left(\frac{\eta_{*}}{\gamma_{*}}-1\right).$
To show the rhs is less than 1, we observe that
$\left(\frac{R^{\prime}\left(\eta_{*}^{-1}\right)}{\eta_{*}^{2}-R^{\prime}\left(\eta_{*}^{-1}\right)}\right)\left(\frac{\eta_{*}}{\gamma_{*}}-1\right)<1\Leftrightarrow\frac{R^{\prime}\left(\eta_{*}^{-1}\right)}{\eta_{*}^{2}-R^{\prime}\left(\eta_{*}^{-1}\right)}<\frac{\eta_{*}\gamma_{*}}{\eta_{*}^{2}-\eta_{*}\gamma_{*}}\stackrel{{\scriptstyle(i)}}{{\Leftrightarrow}}R^{\prime}\left(\eta_{*}^{-1}\right)$ (86)
$<\eta_{*}\gamma_{*}\stackrel{{\scriptstyle(ii)}}{{\Leftrightarrow}}-\frac{\eta_{*}^{-1}R^{\prime}\left(\eta_{*}^{-1}\right)}{R\left(\eta_{*}^{-1}\right)}<1$
where in $(i)$ we used that $x\mapsto\frac{x}{\eta_{*}^{2}-x}$ is strictly increasing and in $(ii)$ we used (20c). Finally, we conclude the proof by noting that the rhs of (86) holds true by Lemma A.8, (d). ∎
Proof of Corollary B.7.
Note that
$\lim_{(s,t)\rightarrow\infty}\left(\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\mathbf{x}^{t}-\mathbf{x}^{s}\right\|^{2}\right)=\lim_{(s,t)\rightarrow\infty}\left(\delta_{ss}+\delta_{tt}-2\delta_{st}\right)=0$
$\lim_{(s,t)\rightarrow\infty}\left(\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\mathbf{y}^{t}-\mathbf{y}^{s}\right\|^{2}\right)=\lim_{(s,t)\rightarrow\infty}\kappa_{*}\left(\delta_{ss}+\delta_{tt}-2\delta_{st}\right)=0$
using Proposition B.6. The convergence of iterates $\mathbf{r}_{1t},\mathbf{r}_{2t}$ follows from $\mathbf{r}_{2t}=\mathbf{x}^{t}+\bm{\beta}^{\star},\mathbf{r}_{1t}=\mathbf{y}^{t}+\bm{\beta}^{\star}+\mathbf{e}$. The convergence of $\hat{\mathbf{x}}_{1t},\hat{\mathbf{x}}_{2t}$ follows from the fact they can be expressed as Lipschitz function applied to iterates $\mathbf{r}_{1,t-1}$ and $\mathbf{r}_{2t}$, i.e. (59a) and (59c). ∎
B.3.2 Track regularized estimator using VAMP iterates
Proof of Proposition B.8.
We assume that $c_{0}>0$. Proof for the case where $d_{-}>0$ is completely analogous. From strong convexity of the penalty function, almost surely, for all sufficiently large $p$,
$\mathcal{L}\left(\hat{\mathbf{x}}_{1t}\right)\geq\mathcal{L}(\hat{\bm{\beta}})\geq\mathcal{L}\left(\hat{\mathbf{x}}_{1t}\right)+\left\langle\mathcal{L}^{\prime}\left(\hat{\mathbf{x}}_{1t}\right),\hat{\bm{\beta}}-\hat{\mathbf{x}}_{1t}\right\rangle+\frac{1}{2}c_{0}\left\|\hat{\bm{\beta}}-\hat{\mathbf{x}}_{1t}\right\|_{2}^{2}$ (87)
holds for any $\mathcal{L}^{\prime}\left(\hat{\mathbf{x}}_{1t}\right)\in\mathbf{X}^{\top}\left(\mathbf{X}\hat{\mathbf{x}}_{1t}-\mathbf{y}\right)+\partial h\left(\hat{\mathbf{x}}_{1t}\right)$ where $\partial h$ denotes sub-gradients of $h$. So we can take $\mathcal{L}^{\prime}\left(\hat{\mathbf{x}}_{1t}\right)=\mathbf{X}^{\top}\left(\mathbf{X}\hat{\mathbf{x}}_{1t}-\mathbf{y}\right)+\gamma_{*}\left(\mathbf{r}_{1,t-1}-\hat{\mathbf{x}}_{1t}\right)$ because $\hat{\mathbf{x}}_{1t}=\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\mathbf{r}_{1,t-1}\right)\Leftrightarrow\mathbf{r}_{1,t-1}-\hat{\mathbf{x}}_{1t}\in\gamma_{*}^{-1}\partial h\left(\hat{\mathbf{x}}_{1t}\right).$ By Cauchy-Schwartz inequality, we have that
$\left\|\hat{\bm{\beta}}-\hat{\mathbf{x}}_{1t}\right\|_{2}\leq\frac{2}{c_{0}}\left\|\mathcal{L}^{\prime}\left(\hat{\mathbf{x}}_{1t}\right)\right\|_{2}$ (88)
Now note that
$\mathcal{L}^{\prime}\left(\hat{\mathbf{x}}_{1t}\right)$ $=\left(\mathbf{X}^{\top}\mathbf{X}-\gamma_{*}I\right)\hat{\mathbf{x}}_{1t}-\mathbf{X}^{\top}\mathbf{y}+\gamma_{*}\mathbf{r}_{1,t-1}$
=$\stackrel{{\scriptstyle(a)}}{{=}}\left(1-\frac{\gamma_{*}}{\eta_{*}}\right)\left(\mathbf{X}^{\top}\mathbf{X}+\gamma_{*}\mathbf{I}_{p}\right)\left(\mathbf{r}_{2t}-\mathbf{r}_{2,t-1}\right)+\left(\mathbf{X}^{\top}\mathbf{X}+\left(\eta_{*}-\gamma_{*}\right)\mathbf{I}_{p}\right)\hat{\mathbf{x}}_{2,t-1}$
$\qquad\qquad-\mathbf{X}^{\top}\mathbf{y}-\left(\eta_{*}-\gamma_{*}\right)\mathbf{r}_{2,t-1}$
=$\stackrel{{\scriptstyle(b)}}{{=}}\left(1-\frac{\gamma_{*}}{\eta_{*}}\right)\left(\mathbf{X}^{\top}\mathbf{X}+\gamma_{*}\mathbf{I}_{p}\right)\left(\mathbf{r}_{2t}-\mathbf{r}_{2,t-1}\right)$
where we used in $(a)$
$\hat{\mathbf{x}}_{1t}=\left(1-\frac{\gamma_{*}}{\eta_{*}}\right)\left(\mathbf{r}_{2t}-\mathbf{r}_{2,t-1}\right)+\hat{\mathbf{x}}_{2,t-1}$ (89)
which follows from (59d),(59b) and in $(b)$,
$\left(\mathbf{X}^{\top}\mathbf{X}+\left(\eta_{*}-\gamma_{*}\right)\mathbf{I}_{p}\right)\hat{\mathbf{x}}_{2,t-1}=\mathbf{X}^{\top}\mathbf{y}+\left(\eta_{*}-\gamma_{*}\right)\mathbf{r}_{2,t-1}$
which follows from (59c). We thus have that almost surely
$\lim_{t\rightarrow\infty}\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\mathcal{L}^{\prime}\left(\hat{\mathbf{x}}_{1t}\right)\right\|_{2}^{2}\leq\lim_{t\rightarrow\infty}\lim_{p\rightarrow\infty}\left(1-\frac{\gamma_{*}}{\eta_{*}}\right)\left\|\mathbf{X}^{\top}\mathbf{X}+\gamma_{*}\mathbf{I}_{p}\right\|_{\mathrm{op}}^{2}\cdot\frac{1}{p}\left\|\mathbf{r}_{2t}-\mathbf{r}_{2,t-1}\right\|_{2}^{2}=0$
which along with (88) implies that
$\lim_{t\rightarrow\infty}\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\hat{\bm{\beta}}-\hat{\mathbf{x}}_{1t}\right\|_{2}^{2}=0$ (90)
By (89), we also have that
$\lim_{t\rightarrow\infty}\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\hat{\bm{\beta}}-\hat{\mathbf{x}}_{2t}\right\|_{2}^{2}=0.$ (91)
Rearranging (59a)—(59d), we have
$\mathbf{r}_{2t}=\hat{\mathbf{x}}_{2t}+\frac{1}{\left(\eta_{*}-\gamma_{*}\right)}\mathbf{X}^{\top}\left(\mathbf{X}\hat{\mathbf{x}}_{2t}-\mathbf{y}\right),\quad\mathbf{r}_{1t}=\hat{\mathbf{x}}_{2t}+\frac{1}{\gamma_{*}}\mathbf{X}^{\top}\left(\mathbf{y}-\mathbf{X}\hat{\mathbf{x}}_{2t}\right)$
which along with (90), (91) implies that
$\lim_{t\rightarrow\infty}\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\mathbf{r}_{1t}-\mathbf{r}_{*}\right\|_{2}^{2}=\lim_{t\rightarrow\infty}\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\mathbf{r}_{2t}-\mathbf{r}_{*}\right\|_{2}^{2}=0.$
B.4 Supporting proofs for result B
B.4.1 Properties of sample adjustment equation
Proof of Lemma B.9.
We can write $g_{p}(\gamma)$ as
[…]
Let us first consider the case where $d_{i}\neq 0$ for all $i$. In this case, only the first sum remain and the denominators of the summands are
$\frac{d_{i}^{2}}{p}\left(\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq+\infty,0}\frac{1}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}+\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=0}\frac{1}{\gamma}\right)$
$\qquad+1-\frac{1}{p}\left(\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq+\infty,0}\frac{\gamma}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}+\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=0}1\right)$
Observe that
$1-\frac{1}{p}\left(\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)_{\neq+\infty,0}}\frac{\gamma}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}+\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=0}1\right)\geq 0$
and
$\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq+\infty,0}\frac{1}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}+\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=0}\frac{1}{\gamma}=0$
$\qquad\qquad\Leftrightarrow\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq+\infty,0}\frac{\gamma}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}+\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=0}1=0.$
These two observations and the assumption that $d_{i}\neq 0$ for all $i$ implies that for all $i\in[p]$, $g_{p}$ is well-defined on $(0,+\infty)$. For the case where $d_{i}=0$ for some $i$, all the denominators in (92) are non-zero (and thus $g_{p}$ is well defined on $(0,+\infty)$) if
$01-\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq 0,+\infty}\frac{\gamma}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}-\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=0}1>$ 0 (93)
which is equivalent to $\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq 0}1>\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq 0,+\infty}\frac{\gamma}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}$. The condition (93) is also necessary when $\exists i\in[p],d_{i}>0$. Meanwhile, we have that
$\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq 0}1\stackrel{{\scriptstyle(a)}}{{\geq}}\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq 0,+\infty}1\stackrel{{\scriptstyle(b)}}{{\geq}}\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq 0,+\infty}\frac{\gamma}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}.$
Therefore, (93) holds if and only if at least one of $(a),(b)$ is strict. Note that $(a)$ is strict if and only if $\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=+\infty}1>0$ and $(b)$ is strict if and only if
$\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq 0,+\infty}\left(1-\frac{\gamma}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}\right)>0\Leftrightarrow\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq 0,+\infty}1>0.$
Note that $\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq 0}1>0$ if and only if $\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq 0,+\infty}1>0$ or $\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=+\infty}1>0$. This shows that (93) holds if and only if there exists some $i\in[p]$ such that $h^{\prime\prime}\left(\hat{{\beta}}_{i}\right)\neq 0$. The latter statement holds if $\|d\|_{0}+\left\|h^{\prime\prime}(\hat{\bm{\beta}})\right\|_{0}>p$.
From now on, suppose that $g_{p}$ is well-defined. It follows from (93) that it is differentiable. Taking derivative of (93) yields
$g^{\prime}_{p}(\gamma)=$ $\frac{1}{p}\sum_{i:d_{i}\neq 0}\frac{\frac{1}{p}\left(\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq+\infty,0}\frac{h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)+d_{i}^{2}}{\left(\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\right)^{2}}+\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=0}\frac{d_{i}^{2}}{\gamma^{2}}\right)}{\left(\frac{1}{p}\left(\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq+\infty,0}\frac{d_{i}^{2}-\gamma}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}+\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=0}\frac{d_{i}^{2}-\gamma}{\gamma}\right)+1\right)^{2}}$ (94)
$+\frac{1}{p}\sum_{i:d_{i}=0}\frac{\frac{1}{p}\left(\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq 0,+\infty}\frac{h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}{\left(\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\right)^{2}}\right)}{\left(\frac{1}{p}\left(\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq 0,+\infty}\frac{-\gamma}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}-\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=0}1\right)+1\right)^{2}}>0$
We claim that given $\gamma\mapsto g(\gamma)$ is well-defined, $g_{p}^{\prime}(\gamma)>0,\forall\gamma\in(0,+\infty)$ if and only if for some $j,\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq+\infty}1>0$. Note that if $\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq 0,+\infty}1>0$, then
$\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq 0,+\infty}\frac{h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}{\left(\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\right)^{2}}>0$
and the above will be positive. Also note that if $\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=0}1>0$, then the assumption $D\neq$ 0 implies that there exists some $i\in[p]$ such that $\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=0}\frac{d_{i}^{2}}{\gamma^{2}}>0$ and the above will be positive. Note that $\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq+\infty}1>0$ if and only if $\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq 0,+\infty}1>0$ or $\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=0}1>0$. Therefore, the positivity of the above follows from the assumption that there exists some $j\in[p]$ such that $h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq+\infty$. Conversely, if $h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=+\infty,\forall j,g_{p}(\gamma)=$ $1,\forall\gamma\in(0,+\infty).$
Note that if $\left\|h^{\prime\prime}(\hat{\bm{\beta}})\right\|_{0}<p$ and for all $i,d_{i}\neq 0$, $\lim_{\gamma\rightarrow 0}g_{p}(\gamma)=0$; if $\left\|h^{\prime\prime}(\hat{\bm{\beta}})\right\|_{0}=0$ and for some $i,d_{i}=0$, $g_{p}$ is not well-defined per discussion above; if $0<\left\|h^{\prime\prime}(\hat{\bm{\beta}})\right\|_{0}<p$ and for some $i,d_{i}=0$,
$\lim_{\gamma\rightarrow 0}g_{p}(\gamma)=\frac{p-\|d\|_{0}}{\left\|h^{\prime\prime}(\hat{\bm{\beta}})\right\|_{0}}<1$
given that $\|d\|_{0}+\left\|h^{\prime\prime}(\hat{\bm{\beta}})\right\|_{0}>p$; if $\left\|h^{\prime\prime}(\hat{\bm{\beta}})\right\|_{0}=p,$
$\lim_{\gamma\rightarrow 0}g_{p}(\gamma)=\frac{1}{p}\left(\sum_{i:d_{i}\neq 0}\frac{1}{\frac{1}{p}\left(\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq+\infty,0}h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\right)+1}+\sum_{i:d_{i}=0}1\right)<1$
since $\mathbf{D}\neq 0$. We also have that
$\lim_{\gamma\rightarrow+\infty}g_{p}(\gamma)=\frac{1}{1-\left(\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq+\infty}1\right)}\in(1,+\infty]$
if for some $i$, $h^{\prime\prime}\left(\hat{{\beta}}_{i}\right)\neq+\infty$. The proof is complete after combining these facts. ∎
B.4.2 Population limit of the adjustment equation
Proof of Lemma B.11.
We can write $g_{\infty}(\gamma)$ as
$g_{\infty}(\gamma)=\mathbb{E}\frac{\mathbb{I}\left(\mathsf{D}^{2}>0\right)}{\left(\mathsf{D}^{2}-\gamma\right)\mathbb{E}\frac{\mathbb{I}(\mathsf{U}\neq+\infty,0)}{\gamma+\mathsf{U}}+\left(\mathsf{D}^{2}-\gamma\right)\frac{1}{\gamma}\mathbb{P}(\mathsf{U}=0)+1}$ (95)
$\qquad\qquad\qquad+\frac{\mathbb{P}\left(\mathsf{D}^{2}=0\right)}{\mathbb{E}\frac{-\gamma\mathbb{I}(\mathsf{U}\neq+\infty,0)}{\gamma+\mathsf{U}}-\mathbb{P}(\mathsf{U}=0)+1}$
Note that the denominators of both terms in (95) are non-zero (and thus $g_{\infty}$ is well-defined) if
$01-\mathbb{E}\frac{\gamma\mathbb{I}(\mathsf{U}\neq+\infty,0)}{\gamma+\mathsf{U}}-\mathbb{P}(\mathsf{U}=0)>$ 0 (96)
which is equivalent to $\mathbb{P}(\mathsf{U}\neq 0)>\mathbb{E}\frac{\gamma\mathbb{I}(\mathsf{U}\neq+\infty,0)}{\gamma+\mathsf{U}}$. Meanwhile we have that
$\mathbb{P}(\mathsf{U}\neq 0)\stackrel{{\scriptstyle(a)}}{{\geq}}\mathbb{P}(\mathsf{U}\neq 0,+\infty)\stackrel{{\scriptstyle(b)}}{{\geq}}\mathbb{E}\frac{\gamma\mathbb{I}(\mathsf{U}\neq+\infty,0)}{\gamma+\mathsf{U}}$
Therefore, (96) holds if at least one of $(a),(b)$ is strict. Note that $(a)$ is strict if and only if $\mathbb{P}(\mathsf{U}=+\infty)>0$ and $(b)$ is strict if and only if
$\mathbb{E}\mathbb{I}(\mathsf{U}\neq+\infty,0)\left(1-\frac{\gamma}{\gamma+\mathsf{U}}\right)>0\Leftrightarrow\mathbb{P}(\mathsf{U}\neq 0,+\infty)>0.$
Note that $\mathbb{P}(\mathsf{U}\neq 0)>0$ if and only if $\mathbb{P}(\mathsf{U}\neq 0,+\infty)>0$ or $\mathbb{P}(\mathsf{U}=+\infty)>0$. This shows that (96) holds and thus $g_{\infty}$ is well-defined since $\mathbb{P}(\mathsf{U}\neq 0)>0$ by Lemma A.10.
It follows from (20a), (20c) and (16) that $\gamma_{*}$ is a solution of the equation $g_{\infty}(\gamma)=1$ We prove that $\gamma_{*}$ is a unique solution by showing $g_{\infty}$ is strictly increasing. Applying [90, Proposition A.2.1], we obtain that $g_{\infty}$ is differentiable and can be differentiated inside the expectation as follows
$g_{\infty}^{\prime}(\gamma)=$ $\mathbb{E}\frac{\mathbb{I}\left(\mathsf{D}^{2}>0\right)\left(\mathbb{E}\frac{\mathsf{U}\mathbb{I}(\mathsf{U}\neq+\infty,0)}{(\gamma+\mathsf{U})^{2}}+\mathsf{D}^{2}\mathbb{E}\frac{\mathbb{I}(\mathsf{U}\neq+\infty,0)}{(\gamma+\mathsf{U})^{2}}+\left(\mathsf{D}^{2}\frac{1}{\gamma^{2}}\right)\mathbb{P}(\mathsf{U}=0)\right)}{\left(\left(\mathsf{D}^{2}-\gamma\right)\mathbb{E}\frac{\mathbb{I}(\mathsf{U}\neq+\infty,0)}{\gamma+\mathsf{U}}+\left(\mathsf{D}^{2}-\gamma\right)\frac{1}{\gamma}\mathbb{P}(\mathsf{U}=0)+1\right)^{2}}$
$+\mathbb{E}\frac{\mathbb{I}\left(\mathsf{D}^{2}=0\right)\left(\mathbb{E}\frac{\mathsf{U}\mathbb{I}(\mathsf{U}\neq+\infty,0)}{(\gamma+\mathsf{U})^{2}}\right)}{\left(\mathbb{E}\frac{-\gamma\mathbb{I}(\mathsf{U}\neq+\infty,0)}{\gamma+\mathsf{U}}-\mathbb{P}(\mathsf{U}=0)+1\right)^{2}}$
To prove $g_{\infty}^{\prime}(\gamma)>0,\forall\gamma\in(0,+\infty)$, note that if $\mathbb{P}(\mathsf{U}\neq+\infty,0)>0$, then $\mathbb{E}\frac{\mathsf{U}\mathbb{I}(\mathsf{U}\neq+\infty,0)}{(\gamma+\mathsf{U})^{2}}>0$ and the above will be positive. Also note that if $\mathbb{P}(\mathsf{U}=0)>0$, then $\mathbb{I}\left(\mathsf{D}^{2}>0\right)\left(\mathsf{D}^{2}\frac{1}{\gamma^{2}}\right)\mathbb{P}(\mathsf{U}=0)>$ 0 with positive probability and the above will be positive. Note that $\mathbb{P}(\mathsf{U}\neq+\infty)>0$ if and only if $\mathbb{P}(\mathsf{U}\neq 0$ and $\mathsf{U}\neq+\infty)>0$ or $\mathbb{P}(\mathsf{U}=0)>0$. Therefore, the positivity of $g_{\infty}^{\prime}(\gamma)$ follows from $\mathbb{P}(\mathsf{U}\neq+\infty)>0$ which holds by Lemma A.10. The proof is now complete. ∎
Proof of Proposition B.13.
We first note that
$\hat{\bm{\beta}}=\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\mathbf{r}_{*}\right).$ (97)
This follows from $\mathbf{r}_{*}\in\hat{\bm{\beta}}+\frac{1}{\gamma_{*}}\partial h(\hat{\bm{\beta}})$ and the equivalence relation $\mathbf{r}_{*}\in\hat{\bm{\beta}}+\frac{1}{\gamma_{*}}\partial h(\hat{\bm{\beta}})\Leftrightarrow\hat{\bm{\beta}}=\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\mathbf{r}_{*}\right)$. The former is a consequence of the KKT condition $\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})\in\partial h(\hat{\bm{\beta}})$ and the latter follows from Proposition A.7, $(a)$. Also note that for any $\gamma>0$,
$\mathbb{P}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\in\left\{x\in\mathbb{R}:\frac{1}{\gamma+h^{\prime\prime}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}(x)\right)}\text{ is continuous at }x\right\}\right)=1$ (98)
which follows from that $x\mapsto\frac{1}{\gamma+h^{\prime\prime}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}(x)\right)}$ has only finitely many discontinuities (cf. Lemma 2.7) and that $\tau_{*}>0$. Then, almost surely,
$\lim_{p\rightarrow\infty}\frac{1}{p}\sum_{i=1}^{p}\frac{1}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{i}\right)}$ =$\stackrel{{\scriptstyle(a)}}{{=}}\lim_{p\rightarrow\infty}\frac{1}{p}\sum_{j=1}^{p}\frac{1}{\gamma+h^{\prime\prime}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(r_{*,j}\right)\right)}$
=$\stackrel{{\scriptstyle(b)}}{{=}}\mathbb{E}\frac{1}{\gamma+h^{\prime\prime}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)\right)}$
where $(a)$ follows from (97) and $(b)$ follows from Theorem 5.1, Proposition A.5 and (98). This, along with Assumption 2 and Proposition A.4, implies that almost surely
$\left(\operatorname{diag}\left(\mathbf{D}^{\top}\mathbf{D}\right)-\gamma\right)\left(\frac{1}{p}\sum_{j=1}^{p}\frac{1}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}\right)\stackrel{{\scriptstyle W_{2}}}{{\rightarrow}}\left(\mathsf{D}^{2}-\gamma\right)\mathbb{E}\frac{1}{\gamma+\mathsf{U}}.$ (99)
Almost sure convergence (72) follows from (99), Proposition A.5 and the fact from Lemma B.11 that $01+\left(\mathsf{D}^{2}-\gamma\right)\mathbb{E}\frac{1}{\gamma+\mathsf{U}}>$ 0 almost surely. ∎
B.5 Finite or single coordinate inference under exchangeability
Proof of Corollary 3.8.
We only show that $\frac{{r}_{*,i}-{\beta}^{\star}_{i}}{\sqrt{\tau_{*}}}\Rightarrow N(0,1)$ for $\mathbf{r}_{*}$ defined in (34). (27) then follows from consistency of $\hat{\tau}_{*}$ and $\widehat{\mathsf{adj}}$ (cf. Theorem 5.2) and the Slutsky’s theorem. Let $\mathbf{U}\in\mathbb{R}^{p\times p}$ denote a permutation operator drawn uniformly at random independent of $\bm{\beta}^{\star},\mathbf{X},\bm{\varepsilon}$. We have that
$\left(\mathbf{X}\mathbf{U}^{\top},\mathbf{U}\bm{\beta}^{\star},\bm{\varepsilon}\right)\stackrel{{\scriptstyle L}}{{=}}\left(\mathbf{X},\bm{\beta}^{\star},\bm{\varepsilon}\right)$
where we use =$\stackrel{{\scriptstyle L}}{{=}}$ to denote equality in law. Note that
$\hat{\bm{\beta}}=\underset{\bm{\beta}}{\operatorname{argmin}}\frac{1}{2}\left\|\mathbf{X}\mathbf{U}^{\top}\mathbf{U}\left(\bm{\beta}^{\star}-\bm{\beta}\right)+\bm{\varepsilon}\right\|^{2}+h\left(\mathbf{U}^{\top}\mathbf{U}\bm{\beta}\right)$
$\qquad=\mathbf{U}^{\top}\underset{\mathbf{U}\beta}{\operatorname{argmin}}\frac{1}{2}\left\|\mathbf{X}\mathbf{U}^{\top}\left(\mathbf{U}\bm{\beta}^{\star}-\mathbf{U}\bm{\beta}\right)+\bm{\varepsilon}\right\|^{2}+h(\mathbf{U}\bm{\beta})$
where $h$ applies entry-wise to its argument. The above then implies
$\left(\mathbf{U}\hat{\bm{\beta}},\mathbf{X}\mathbf{U}^{\top},\mathbf{U}\bm{\beta}^{\star},\bm{\varepsilon}\right)$ (100)
$\qquad=\left(\underset{\bm{\beta}}{\operatorname{argmin}}\frac{1}{2}\left\|\mathbf{X}\mathbf{U}^{\top}\left(\mathbf{U}\bm{\beta}^{\star}-\bm{\beta}\right)+\bm{\varepsilon}\right\|^{2}+h(\bm{\beta}),\mathbf{X}\mathbf{U}^{\top},\mathbf{U}\bm{\beta}^{\star},\bm{\varepsilon}\right)$
[…]
$\qquad=\left(\hat{\bm{\beta}},\mathbf{X},\bm{\beta}^{\star},\bm{\varepsilon}\right)$
Below we prove the Corollary for =$\mathcal{L}=\{i,k\},i\neq k$. The general case is analogous. For standard basis $\mathbf{e}_{i},\mathbf{e}_{k}$, and any constant $c_{1},c_{2}\in\mathbb{R}$,
$\mathbb{P}$ $\left(\frac{\mathbf{e}_{i}^{\top}\mathbf{r}_{*}-\mathbf{e}_{i}^{\top}\bm{\beta}^{\star}}{\sqrt{\tau_{*}}}<c_{1},\frac{\mathbf{e}_{k}^{\top}\mathbf{r}_{*}-\mathbf{e}_{k}^{\top}\bm{\beta}^{\star}}{\sqrt{\tau_{*}}}<c_{2}\right)$
=$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{P}\left(\frac{\mathbf{e}_{i}^{\top}\mathbf{U}\mathbf{r}_{*}-\mathbf{e}_{i}^{\top}\mathbf{U}\bm{\beta}^{\star}}{\sqrt{\tau_{*}}}<c_{1},\frac{\mathbf{e}_{k}^{\top}\mathbf{U}\mathbf{r}_{*}-\mathbf{e}_{k}^{\top}\mathbf{U}\bm{\beta}^{\star}}{\sqrt{\tau_{*}}}<c_{2}\right)$
=$\stackrel{{\scriptstyle(b)}}{{=}}\mathbb{E}\left(\mathbb{P}\left(\frac{\mathbf{e}_{i}^{\top}\mathbf{U}\mathbf{r}_{*}-\mathbf{e}_{i}^{\top}\mathbf{U}\bm{\beta}^{\star}}{\sqrt{\tau_{*}}}<c_{1},\frac{\mathbf{e}_{k}^{\top}\mathbf{U}\mathbf{r}_{*}-\mathbf{e}_{k}^{\top}\mathbf{U}\bm{\beta}^{\star}}{\sqrt{\tau_{*}}}<c_{2}\mid\mathcal{F}\left(\bm{\beta}^{\star},\bm{\varepsilon},\mathbf{X}\right)\right)\right)$
=$\stackrel{{\scriptstyle(c)}}{{=}}\mathbb{E}\frac{1}{p(p-1)}\sum_{j_{1}\neq j_{2}\in[p]}\mathbb{I}\left(\frac{1}{\sqrt{\tau_{*}}}\left(r_{*,j_{1}}-\beta_{j_{1}}^{\star}\right)<c_{1}\right)\mathbb{I}\left(\frac{1}{\sqrt{\tau_{*}}}\left(r_{*,j_{2}}-\beta_{j_{2}}^{\star}\right)<c_{2}\right)$
where in $(a)$ we used (100) above, in $(b)$ we used $\mathcal{F}\left(\bm{\beta}^{\star},\bm{\varepsilon},\mathbf{X}\right)$ to denote sigma-field generated by $\bm{\beta}^{\star},\bm{\varepsilon},\mathbf{X}$ and in $(c)$ we used that $\mathbf{U}$ is a permutation operator drawn uniformly at random.
Note that almost surely as $p\to\infty$,
$\bigg{|}\frac{1}{p(p-1)}\sum_{j_{1}\neq j_{2}\in[p]}\mathbb{I}\left(\frac{1}{\sqrt{\tau_{*}}}\left(r_{*,j_{1}}-\beta_{j_{1}}^{\star}\right)<c_{1}\right)\mathbb{I}\left(\frac{1}{\sqrt{\tau_{*}}}\left(r_{*,j_{2}}-\beta_{j_{2}}^{\star}\right)<c_{2}\right)-$
$\qquad\frac{1}{p^{2}}\sum_{j=1}^{p}\mathbb{I}\left(\frac{1}{\sqrt{\tau_{*}}}\left(r_{*,j}-\beta_{j}^{\star}\right)<c_{1}\right)\sum_{j=1}^{p}\mathbb{I}\left(\frac{1}{\sqrt{\tau_{*}}}\left(r_{*,j}-\beta_{j}^{\star}\right)<c_{2}\right)\bigg{|}\to 0.$
Note also that for $\iota=1,2$ almost surely
$\lim_{p\to\infty}\frac{1}{p}\sum_{j=1}^{p}\mathbb{I}\left(\frac{1}{\sqrt{\tau_{*}}}\left(r_{*,j}-\beta_{j}^{\star}\right)<c_{\iota}\right)=\mathbb{P}(\mathsf{Z}<c_{\iota})$
where $\mathsf{Z}\sim N(0,1)$. Here, we used Theorem 5.2 and Proposition A.5. Using dominated convergence theorem, we conclude that
$\mathbb{P}\left(\frac{\mathbf{e}_{i}^{\top}\mathbf{r}_{*}-\mathbf{e}_{i}^{\top}\bm{\beta}^{\star}}{\sqrt{\tau_{*}}}<c_{1},\frac{\mathbf{e}_{k}^{\top}\mathbf{r}_{*}-\mathbf{e}_{k}^{\top}\bm{\beta}^{\star}}{\sqrt{\tau_{*}}}<c_{2}\right)\rightarrow\mathbb{P}(\mathsf{Z}<c_{1})\mathbb{P}(\mathsf{Z}<c_{2})$
as required. ∎
B.6 Hypothesis testing and confidence intervals
Proof of Corollary 3.9.
To see $(a)$, We have that almost surely
$\begin{gathered}\lim_{p\to\infty}\frac{\frac{1}{p}\sum_{j=1}^{p}\mathbb{I}\left(P_{j}\leq\alpha,\beta_{j}^{\star}=0\right)}{\frac{1}{p}\sum_{j=1}^{p}\mathbb{I}\left(\beta_{j}^{\star}=0\right)}=\lim_{p\to\infty}\frac{\frac{1}{p}\sum_{j=1}^{p}\mathbb{I}\left(\left|\frac{\hat{r}_{*,j}-\beta_{j}^{\star}}{\sqrt{\hat{\tau}_{*}}}\right|\geq\Phi^{-1}\left(1-\frac{\alpha}{2}\right),\absolutevalue{\beta_{j}^{\star}}\leq\frac{\mu_{0}}{2}\right)}{\frac{1}{p}\sum_{j=1}^{p}\mathbb{I}\left(\absolutevalue{\beta_{j}^{\star}}\leq\frac{\mu_{0}}{2}\right)}\\ =\frac{\mathbb{P}\left(|\mathsf{Z}|\geq\Phi^{-1}\left(1-\frac{\alpha}{2}\right),\absolutevalue{\mathsf{B}^{\star}}\leq\frac{\mu_{0}}{2}\right)}{\mathbb{P}\left(\absolutevalue{\mathsf{B}^{\star}}\leq\frac{\mu_{0}}{2}\right)}=\mathbb{P}\left(|\mathsf{Z}|\geq\Phi^{-1}\left(1-\frac{\alpha}{2}\right)\right)=\alpha\end{gathered}$
by Theorem 5.2 and Proposition A.5. Using exchangeability of columns of $\mathbf{X}$
$\mathbb{E}\frac{\frac{1}{p}\sum_{j=1}^{p}\mathbb{I}\left(P_{j}\leq\alpha,\beta_{j}^{\star}=0\right)}{\frac{1}{p}\sum_{j=1}^{p}\mathbb{I}\left(\beta_{j}^{\star}=0\right)}=\frac{\mathbb{P}\left(T_{i}=1\right)\frac{1}{p}\sum_{j=1}^{p}\mathbb{I}\left(\beta_{j}^{\star}=0\right)}{\frac{1}{p}\sum_{j=1}^{p}\mathbb{I}\left(\beta_{j}^{\star}=0\right)}=\mathbb{P}\left(T_{i}=1\right)$
The the coordinate-wise result follows from an application of the dominated convergence theorem.
To see $(b)$, note that by Theorem 5.2 and Proposition A.5, almost surely
$\lim_{p\to\infty}\frac{1}{p}\sum_{i=1}^{p}\mathbb{I}\left({\beta}^{\star}_{i}\in\mathsf{CI}_{i}\right)=\lim_{p\to\infty}\frac{1}{p}\sum_{i=1}^{p}\mathbb{I}\left(a<\frac{{\beta}^{\star}_{i}-\hat{{\beta}}^{u}_{i}}{\sqrt{\hat{\tau}_{*}}}<b\right)=\mathbb{P}(a<\mathsf{Z}<b)=1-\alpha.$
Remark B.14 (Asymptotic limit of TPR).
Note that we can further calculate the exact asymptotic limit of the TPR as follows. Under the assumption of Corollary 3.9 (a), we have that almost surely
$\lim_{p\to\infty}\mathsf{TPR}(p)$ $=\lim_{p\to\infty}\frac{\sum_{j=1}^{p}\mathbb{I}\left(P_{j}\leq\alpha,\left|\beta_{j}^{\star}\right|\geq\mu_{0}\right)}{\sum_{j=1}^{p}\mathbb{I}\left(\absolutevalue{\beta_{j}^{\star}}\geq\mu_{0}\right)}$
$=\lim_{p\to\infty}\frac{\frac{1}{p}\sum_{j=1}^{p}\mathbb{I}\left(\left|\frac{r_{*,j}}{\sqrt{\hat{\tau}_{*}}}\right|\geq\Phi^{-1}\left(1-\frac{\alpha}{2}\right),\left|{\beta_{j}^{\star}}\right|\geq\mu_{0}\right)}{\frac{1}{p}\sum_{j=1}^{p}\mathbb{I}\left(\absolutevalue{\beta_{j}^{\star}}\geq\mu_{0}\right)}$
$=\frac{\mathbb{P}\left(\left|\frac{1}{\sqrt{\tau_{*}}}\mathsf{B}^{\star}+\mathsf{Z}\right|\geq\Phi^{-1}\left(1-\frac{\alpha}{2}\right),\left|\mathsf{B}^{\star}\right|\geq\mu_{0}\right)}{\mathbb{P}\left(\left|\mathsf{B}^{\star}\right|\geq\mu_{0}\right)}$
where we used in the second line Theorem 5.2 and Proposition A.5.
Appendix C PCR-Spectrum-Aware Debiasing
C.1 Pseudo-code for PCR-Spectrum-Aware Debiasing
Algorithm 1 below summarizes the PCR-Spectrum-Aware Debiasing procedure in algorithmic format.
Algorithm 1 PCR-Spectrum-Aware Debiasing
1:Response and design $(\mathbf{y},\mathbf{X})$, a penalty function $h$ and an index set of PCs $\mathcal{J}\subset\mathcal{N}$ (see (LABEL:defJJ)).
2:Conduct eigen-decomposition: $\mathbf{X}^{\top}\mathbf{X}=\mathbf{O}^{\top}\mathbf{D}^{\top}\mathbf{D}\mathbf{O}$ and let $\mathbf{O}_{\mathcal{J}},\mathbf{O}_{\bar{\mathcal{J}}}$ be PCs indexed by $\mathcal{J}$ and $\bar{\mathcal{J}}=\mathcal{N}\setminus\mathcal{J}$ respectively.
3:Compute alignment PCR estimator
$\hat{\bm{\beta}}_{\mathsf{al}}\leftarrow\mathbf{O}_{\mathcal{J}}^{\top}\left(\bm{W}_{\mathcal{J}}^{\top}\bm{W}_{\mathcal{J}}\right)^{-1}\bm{W}_{\mathcal{J}}^{\top}\mathbf{y}$
where $\bm{W}_{\mathcal{J}}:=\mathbf{X}\mathbf{O}_{\mathcal{J}}^{\top}.$
4:Construct new data
$\mathbf{y}_{\mathsf{new}}\leftarrow\quantity(\mathbf{D}_{\bar{\mathcal{J}}}^{\top}\mathbf{D}_{\bar{\mathcal{J}}})^{1/2}\left(\bm{W}_{\bar{\mathcal{J}}}^{\top}\bm{W}_{\bar{\mathcal{J}}}\right)^{-1}\bm{W}_{\bar{\mathcal{J}}}^{\top}\mathbf{y},\quad\mathbf{X}_{\mathsf{new}}\leftarrow\quantity(\mathbf{D}_{\bar{\mathcal{J}}}^{\top}\mathbf{D}_{\bar{\mathcal{J}}})^{1/2}\mathbf{O}_{\bar{\mathcal{J}}}$
where $\bm{W}_{\bar{\mathcal{J}}}=\mathbf{X}\mathbf{O}_{\bar{\mathcal{J}}}^{\top}$ and $\mathbf{D}_{\bar{\mathcal{J}}}$ consists of columns of $\mathbf{D}$ indexed by $\bar{\mathcal{J}}$.
5:Find minimizer $\hat{\bm{\beta}}$ of $\mathcal{L}(\bm{\cdot}\;;\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}})$ for $\mathcal{L}$ defined in (1)
6:Compute the eigenvalues $(d_{i}^{2})_{i=1}^{p}$ of $\mathbf{X}_{\mathsf{new}}^{\top}\mathbf{X}_{\mathsf{new}}$
7:Find solution $\widehat{\mathsf{adj}}(\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}},h)$ of (23) and compute complement PCR estimator
$\hat{\bm{\beta}}_{\mathsf{co}}\leftarrow\hat{\bm{\beta}}+\widehat{\mathsf{adj}}^{-1}\mathbf{X}_{\mathsf{new}}^{\top}(\mathbf{y}_{\mathsf{new}}-\mathbf{X}_{\mathsf{new}}\hat{\bm{\beta}})$ (101)
and $\hat{\tau}_{*}(\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}},h)$ from (26)
8:PCR-Spectrum-Awaure estimator
$\hat{\bm{\beta}}^{u}_{\mathsf{pcr}}\leftarrow\hat{\bm{\beta}}_{\mathsf{al}}+\hat{\bm{\beta}}_{\mathsf{co}}$
and the associated variance estimator $\hat{\tau}_{*}\leftarrow\hat{\tau}_{*}(\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}},h)$.
C.2 Asymptotic normality
Proof of Theorem 4.1.
(a) Alignment PCR. Let $\mathbf{D}_{\mathcal{J}}\in\mathbb{R}^{n\times J}$ consist of columns of $\mathbf{D}$ indexed by $\mathcal{J},\mathbf{O}_{\mathcal{J}}\in\mathbb{R}^{J\times p}$ consist of rows of $\mathbf{O}$ indexed by $\mathcal{J}$, and $\mathbf{P}_{\mathcal{J}}=\mathbf{O}_{\mathcal{J}}^{\top}\mathbf{O}_{\mathcal{J}}$. Note that
$\hat{\bm{\beta}}_{\mathsf{al}}(\mathcal{J})$ $=\mathbf{O}_{\mathcal{J}}^{\top}\hat{\bm{\theta}}_{\mathsf{pcr}}(\mathcal{J})$ (102)
$=\mathbf{O}_{\mathcal{J}}^{\top}\left(\mathbf{W}_{\mathcal{J}}^{\top}\mathbf{W}_{\mathcal{J}}\right)^{-1}\mathbf{W}_{\mathcal{J}}^{\top}\mathbf{y}$
$=\mathbf{O}_{\mathcal{J}}^{\top}\left(\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{D}_{\mathcal{J}}\right)^{-1}\mathbf{D}_{\mathcal{J}}^{\top}\left(\mathbf{D}\mathbf{O}\bm{\beta}^{\star}+\mathbf{Q}\bm{\varepsilon}\right)$
$=\mathbf{O}_{\mathcal{J}}^{\top}\mathbf{O}_{\mathcal{J}}\bm{\beta}^{\star}+\mathbf{O}_{\mathcal{J}}^{\top}\left(\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{D}_{\mathcal{J}}\right)^{-1}\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{Q}\bm{\varepsilon}$
$=\bm{\beta}^{\star}_{\mathsf{al}}+\mathbf{O}_{\mathcal{J}}^{\top}\mathbf{O}_{\mathcal{J}}\bm{\zeta}^{\star}+\mathbf{O}_{\mathcal{J}}^{\top}\left(\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{D}_{\mathcal{J}}\right)^{-1}\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{Q}\bm{\varepsilon}$
where we used that
$\mathbf{W}_{\mathcal{J}}=\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}\mathbf{O}_{\mathcal{J}}^{\top}=\mathbf{Q}^{\top}\mathbf{D}_{\mathcal{J}},\quad\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\bm{\varepsilon}=\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}\bm{\beta}^{\star}+\bm{\varepsilon}$
in the penultimate equality and (30) in the last equality.
Using rotational invariance of $\mathbf{O}$, we have
$\mathbb{E}\quantity[\quantity(\frac{1}{p}\left\|\mathbf{P}_{\mathcal{J}}\bm{\zeta}^{\star}\right\|_{2}^{2})^{2}\mid\bm{\zeta}^{\star}]$ $=\frac{1}{p^{2}}\mathbb{E}\quantity[\left\|\mathbf{O}_{\mathcal{J}}\bm{\zeta}^{\star}\right\|_{2}^{4}\mid\bm{\zeta}^{\star}]$ (103)
$=\quantity(\frac{\left\|\bm{\zeta}^{\star}\right\|_{2}^{2}}{p})^{2}\mathbb{E}\quantity[\sum_{i=1}^{J}O_{1i}^{2}]^{2}=O\quantity(\frac{1}{p^{2}})$
where we used that $J$ is finite not growing with $p$ and basic moment property of entries of $\mathbf{O}$ (see e.g. [65, Proposition 2.5]). It follows from a straightforward application of Markov inequality and Borel-Cantelli lemma that almost surely
$\lim_{p\to\infty}\frac{1}{p}\left\|\mathbf{P}_{\mathcal{J}}\bm{\zeta}^{\star}\right\|_{2}^{2}=0.$ (104)
Meanwhile, using $\mathbf{Q}\bm{\varepsilon}\stackrel{{\scriptstyle L}}{{=}}\bm{\varepsilon}$, we have
$\frac{1}{p}\left\|\mathbf{O}_{\mathcal{J}}^{\top}\left(\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{D}_{\mathcal{J}}\right)^{-1}\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{Q}\bm{\varepsilon}\right\|_{2}^{2}\stackrel{{\scriptstyle L}}{{=}}\frac{1}{p}\bm{\varepsilon}^{\top}\mathbf{D}_{\mathcal{J}}\left(\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{D}_{\mathcal{J}}\right)^{-2}\mathbf{D}_{\mathcal{J}}^{\top}\bm{\varepsilon}=\frac{1}{p}\sum_{i\in\mathcal{J}}\frac{{\varepsilon}_{i}^{2}}{d_{i}^{2}}.$ (105)
Using $J$ is finite, we obtain that almost surely,
$\lim_{p\to\infty}\frac{1}{p}\left\|\mathbf{O}_{\mathcal{J}}^{\top}\left(\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{D}_{\mathcal{J}}\right)^{-1}\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{Q}\bm{\varepsilon}\right\|_{2}^{2}=0.$ (106)
The result then follows from (104) and (106).
(b) Complement PCR. Similarly to (102), we have that
$\hat{\bm{\theta}}_{\mathsf{pcr}}(\bar{\mathcal{J}})=\mathbf{O}_{\bar{\mathcal{J}}}\bm{\zeta}^{\star}+\left(\mathbf{D}_{\bar{\mathcal{J}}}^{\top}\mathbf{D}_{\bar{\mathcal{J}}}\right)^{-1}\mathbf{D}_{\bar{\mathcal{J}}}^{\top}\mathbf{Q}\bm{\varepsilon}.$ (107)
It follows that
$\mathbf{y}_{\mathsf{new}}=\left(\mathbf{D}_{\bar{\mathcal{J}}}^{\top}\mathbf{D}_{\bar{\mathcal{J}}}\right)^{\frac{1}{2}}\hat{\bm{\theta}}_{\mathsf{pcr}}(\bar{\mathcal{J}})\in\mathbb{R}^{N-J},\qquad\mathbf{X}_{\mathsf{new}}=\left(\mathbf{D}_{\bar{\mathcal{J}}}^{\top}\mathbf{D}_{\bar{\mathcal{J}}}\right)^{\frac{1}{2}}\mathbf{O}_{\bar{\mathcal{J}}}\in\mathbb{R}^{(N-J)\times p}$
defined in (32) satisfy the following relation:
$\mathbf{y}_{\mathsf{new}}=\mathbf{X}_{\mathsf{new}}\bm{\zeta}^{\star}+\epsilon_{\mathsf{new}}$ (108)
for
$\epsilon_{\mathsf{new}}=\left(\mathbf{D}_{\bar{\mathcal{J}}}^{\top}\mathbf{D}_{\bar{\mathcal{J}}}\right)^{-\frac{1}{2}}\mathbf{D}_{\bar{\mathcal{J}}}^{\top}\mathbf{Q}\bm{\varepsilon}\sim N\left(\bm{0},\sigma^{2}\mathbf{I}_{N-J}\right).$
Note that the new design matrix $\mathbf{X}_{\mathsf{new}}$ admits singular value decomposition
$\mathbf{X}_{\mathsf{new}}=\mathbf{Q}_{\mathsf{new}}^{\top}\mathbf{D}_{\mathsf{new}}\mathbf{O}$
where
$\mathbf{Q}_{\mathsf{new}}=\mathbf{I}_{N-J},\qquad\mathbf{D}_{\mathsf{new}}=\left[\left(\mathbf{D}_{\bar{\mathcal{J}}}^{\top}\mathbf{D}_{\bar{\mathcal{J}}}\right)^{\frac{1}{2}},\bm{0}_{(N-J)\times(p+J-N)}\right]\in\mathbb{R}^{(N-J)\times p}.$
Note that since $J$ is finite not growing with $n,p$,
$\mathbf{D}_{\mathsf{new}}^{\top}\bm{1}_{(N-J)\times 1}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{D}.$
The above, along with the assumption we made in Theorem 4.1, reduces the new regression problem defined by (108) to the same one considered in Section 3. Since $\hat{\bm{\beta}}_{\mathsf{co}}$ is Spectrum-Aware debiased estimator with respect to the new regression problem, the result follows from Theorem 3.1. The consistency of $\hat{\sigma}^{2}$ follows a similar reasoning.
(c) Debiased PCR. By the definition of $\hat{\bm{\beta}}^{u}_{\mathsf{pcr}}$, we have that
$\hat{\tau}_{*}^{-1/2}\quantity(\hat{\bm{\beta}}^{u}_{\mathsf{pcr}}-\bm{\beta}^{\star})=\hat{\tau}_{*}^{-1/2}\quantity(\hat{\bm{\beta}}_{\mathsf{al}}-\bm{\beta}^{\star}_{\mathsf{al}})+\hat{\tau}_{*}^{-1/2}\quantity(\hat{\bm{\beta}}_{\mathsf{co}}-\bm{\zeta}^{\star}).$
The result then follows from (a), (b) above and Proposition A.4. ∎
C.3 Finite or single coordinate inference
Corollary C.1.
Suppose Assumptions 48 hold. If $\quantity({\zeta}^{\star}_{j})_{j=1}^{p}$ are exchangeable as in Definition 3.7, then for any fixed, finite index set $\mathcal{I}\subset[p],$ we have that almost surely as $p\to\infty$,
$\hat{\bm{\beta}}_{\mathsf{al},\mathcal{I}}(\mathcal{J})\to\bm{\beta}^{\star}_{\mathsf{al},\mathcal{I}},\;\;\hat{\tau}_{*}^{-1/2}\quantity(\hat{\bm{\beta}}_{\mathsf{co},\mathcal{I}}(\bar{\mathcal{J}})-\bm{\zeta}^{\star}_{\mathcal{I}})\Rightarrow N(\bm{0},\mathbf{I}_{|\mathcal{I}|})$ (109)
$\hat{\tau}_{*}^{-1/2}\quantity(\hat{\bm{\beta}}^{u}_{\mathsf{pcr},\mathcal{I}}-\bm{\beta}^{\star}_{\mathcal{I}})\Rightarrow N\quantity(\bm{0},\mathbf{I}_{|\mathcal{I}|}).$
Proof of Corollary C.1.
To see the first result in (109), recall from (102), we have that
$\hat{\bm{\beta}}_{\mathsf{al}}(\mathcal{J})=\bm{\beta}^{\star}_{\mathsf{al}}+\mathbf{O}_{\mathcal{J}}^{\top}\mathbf{O}_{\mathcal{J}}\bm{\zeta}^{\star}+\mathbf{O}_{\mathcal{J}}^{\top}\left(\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{D}_{\mathcal{J}}\right)^{-1}\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{Q}\bm{\varepsilon}$ (110)
Note that when $\bm{\zeta}^{\star}$ is exchangeable, we have that for any fixed $i\in[p]$
$\mathbb{E}\left[\left(\left(\mathbf{O}_{\mathcal{J}}^{\top}\mathbf{O}_{\mathcal{J}}\bm{\zeta}^{\star}\right)_{i}\right)^{2}\right]$ $=\mathbb{E}\quantity[\quantity(\mathbf{e}_{i}^{\top}\mathbf{U}\mathbf{O}_{\mathcal{J}}^{\top}\mathbf{O}_{\mathcal{J}}\mathbf{U}^{\top}\mathbf{U}\bm{\zeta}^{\star})^{2}]$ (111)
$=\mathbb{E}\left[\left(\frac{1}{p}\left\|\mathbf{P}_{\mathcal{J}}\bm{\zeta}^{\star}\right\|_{2}^{2}\right)^{2}\right]=O\left(\frac{1}{p^{2}}\right)$
where we used that for a permutation matrix $\mathbf{U}\in\mathbb{R}^{p\times p}$ drawn uniformly, $(\mathbf{O}_{\mathcal{J}}\mathbf{U}^{\top},\mathbf{U}\bm{\zeta}^{\star})\stackrel{{\scriptstyle L}}{{=}}(\mathbf{O}_{\mathcal{J}},\bm{\zeta}^{\star})$ and (103). And by rotational invariance of $\mathbf{O}$,
$\mathbb{E}\left(\mathbf{O}_{\mathcal{J}}^{\top}\left(\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{D}_{\mathcal{J}}\right)^{-1}\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{Q}\bm{\varepsilon}\right)_{i}^{2}=\mathbb{E}\frac{1}{p}\left\|\mathbf{O}_{\mathcal{J}}^{\top}\left(\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{D}_{\mathcal{J}}\right)^{-1}\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{Q}\bm{\varepsilon}\right\|_{2}^{2}=O\quantity(\frac{1}{p^{2}})$ (112)
where we used (105) at the last equality. The first result in (109) then follows from Markov inequality and Borel-Cantelli lemma. The second result in (109) can be proved similarly to Corollary 3.8. The third result in (109) follows from the first two results and an application of the Slutsky’s theorem. ∎
C.4 Alignment test
Proof of Corollary 4.3.
Similarly to (102), we have that
$\hat{\bm{\theta}}_{\mathsf{pcr}}(\mathcal{J})-\bm{\upsilon}^{\star}=\mathbf{O}_{\mathcal{J}}\bm{\zeta}^{\star}+\left(\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{D}_{\mathcal{J}}\right)^{-1}\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{Q}\bm{\varepsilon}$ (113)
Now note that by basic properties of Haar measure on orthogonal groups [65], as $p\rightarrow\infty$,
$\mathbf{O}_{\mathcal{J}}\bm{\zeta}^{\star}\Rightarrow N\left(\bm{0},\mathbb{E}\left(\mathsf{C}^{\star}\right)^{2}\cdot\mathbf{I}_{J}\right)$
where we used the assumption that $\bm{\zeta}^{\star}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{C}^{\star}$, and that
$\left(\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{D}_{\mathcal{J}}\right)^{-1}\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{Q}\bm{\varepsilon}\sim N\left(\bm{0},\sigma^{2}\cdot\left(\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{D}_{\mathcal{J}}\right)^{-1}\right).$
By independence of $\mathbf{O}$ and $\bm{\varepsilon}$, we have that
$\mathbf{O}_{\mathcal{J}}\bm{\zeta}^{\star}+\left(\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{D}_{\mathcal{J}}\right)^{-1}\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{Q}\bm{\varepsilon}\Rightarrow N\left(\bm{0},\mathbb{E}\left(\mathsf{C}^{\star}\right)^{2}\cdot\mathbf{I}_{J}+\sigma^{2}\cdot\left(\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{D}_{\mathcal{J}}\right)^{-1}\right)$
Desired result then follows from the fact that $\hat{\omega}$ consistently estimates $\mathbb{E}\left(\mathsf{C}^{\star}\right)^{2}$ and $\hat{\sigma}^{2}$ consistently estimate $\sigma^{2}$. That is, almost surely
$\hat{\omega}=p^{-1}\norm{\hat{\bm{\beta}}_{\mathsf{co}}}^{2}-\hat{\tau}_{*}\to\mathbb{E}(\mathsf{C}^{\star})^{2},\quad\hat{\sigma}^{2}\to\sigma^{2}$
as $p\to\infty$. The former follows from the fact that almost surely $\hat{\bm{\beta}}_{\mathsf{co}}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{C}^{\star}+\sqrt{\tau_{*}}\mathsf{Z}$ for $\mathsf{Z}$ independent of $\mathsf{C}^{\star}$. ∎
Appendix D Conjectures for ellipsoidal models
We conjecture that debiasing is possible in a more general settings than considered in this paper. Namely, one would like to consider the design matrix $\mathbf{X}=\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}\bm{\Sigma}^{1/2}$ where $\bm{\Sigma}\in\mathbb{R}^{p\times p}$ is non-singular, $\mathbf{Q}\in\mathbb{R}^{n\times n},\mathbf{O}\in\mathbb{R}^{p\times p}$ are orthogonal matrices and $\mathbf{D}\in\mathbb{R}^{n\times p}$ is diagonal matrix. We assume that $\bm{\Sigma}\in\mathbb{R}^{p\times p}$ is observed and $\mathbf{O}$ is drawn uniformly from the orthogonal group $\mathbb{O}(p)$ independent of $\bm{\varepsilon},\mathbf{D},\mathbf{Q}$. We refer to this class of random design matrices as ellipsoidal invariant designs. The special case where $\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}$ is an isotropic Gaussian matrix is studied extensively in prior literature [12, 18, 54, 53, 11]. Similarly to the anisotropic Gaussian design case, the challenge in applying such a model arises when $\bm{\Sigma}$ is unknown. We discuss this in Remark D.2 at the end.
Furthermore, one would like to consider the case where the convex penalty function $\vec{h}:\mathbb{R}^{p}\mapsto\mathbb{R}$ is non-separable (e.g. SLOPE, group-Lasso) and $\hat{\bm{\beta}}\in\underset{\mathbf{b}\in\mathbb{R}^{p}}{\arg\min}\frac{1}{2}\|\mathbf{y}-\mathbf{X}\mathbf{b}\|^{2}+\vec{h}\left(\mathbf{b}\right).$ where $\vec{h}$ is assumed to be proper and closed. To illustrate, we give debiasing formulas under the case $\sigma^{2}=1$:
$\hat{\bm{\beta}}^{u}=\hat{\bm{\beta}}+\frac{1}{\widehat{\mathsf{adj}}}\bm{\Sigma}^{-1}\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})$ (114)
where $\widehat{\mathsf{adj}}$ is solution of the following equation
$\frac{1}{p}{{\sum}}_{i=1}^{p}\frac{1}{\frac{d_{i}^{2}-\widehat{\mathsf{adj}}}{p}\operatorname{Tr}\left(\left(\widehat{\mathsf{adj}}\cdot\mathbf{I}_{p}+\bm{\Sigma}^{-1}\left(\nabla^{2}\vec{h}(\hat{\bm{\beta}})\right)\right)^{-1}\right)+1}=1.$ (115)
Here, we assumed that $\vec{h}$ is twice-differentiable or that it admits a twice-differentiable extension as in Lemma 2.7. Notice that the equation (115) becomes (23) if one let $\bm{\Sigma}=\mathbf{I}_{p}$ and $(\vec{h}(x))_{i}=h(x_{i}),\forall i\in p$ for some $h:\mathbb{R}\mapsto\mathbb{R}$. Analogous to (26), we define
$\hat{\eta}_{*}(p):=\left(\frac{1}{p}\operatorname{Tr}\quantity(\widehat{\mathsf{adj}}\cdot\mathbf{I}_{p}+\bm{\Sigma}^{-1}\nabla^{2}\vec{h}(\hat{\bm{\beta}}))\right)^{-1}$ (116)
$\hat{\mathbf{{r}}}_{**}(p):=\hat{\bm{\beta}}+\frac{1}{\hat{\eta}_{*}-\widehat{\mathsf{adj}}}\bm{\Sigma}^{-1}\mathbf{X}^{\top}(\mathbf{X}\hat{\bm{\beta}}-\mathbf{y}),\quad\hat{\tau}_{**}(p):=\frac{\frac{1}{p}\left\|\mathbf{X}\hat{\mathbf{{r}}}_{**}-\mathbf{y}\right\|^{2}-\frac{n}{p}}{\frac{1}{p}\sum_{i=1}^{p}d_{i}^{2}}$
$\hat{\tau}_{*}(p):=\left(\frac{\hat{\eta}_{*}}{\widehat{\mathsf{adj}}}\right)^{2}\frac{1}{p}{{\sum}}_{i=1}^{p}\frac{d_{i}^{2}}{\left(d_{i}^{2}+\hat{\eta}_{*}-\widehat{\mathsf{adj}}\right)^{2}}$
$\qquad\qquad+\left(\frac{\hat{\eta}_{*}-\widehat{\mathsf{adj}}}{\widehat{\mathsf{adj}}}\right)^{2}\left(\frac{1}{p}{{\sum}}_{i=1}^{p}\left(\frac{\hat{\eta}_{*}}{d_{i}^{2}+\hat{\eta}_{*}-\widehat{\mathsf{adj}}}\right)^{2}-1\right)\hat{\tau}_{**}$
One can then make the following conjecture on the distribution of $\hat{\bm{\beta}}^{u}$.
Conjecture D.1.
Under suitable conditions, there is a unique solution $\widehat{\mathsf{adj}}$ of (115) and
$\hat{\tau}_{*}^{-1/2}(\hat{\bm{\beta}}^{u}-\bm{\beta}^{\star})=\bm{\Sigma}^{1/2}\mathbf{z}+O\quantity(p^{-1/2})$
where $\mathbf{z}\sim N(\bm{0},\mathbf{I}_{p})$ and $O\quantity(p^{-1/2})$ denotes a vector $\mathbf{v}\in\mathbb{R}^{p}$ satisfying $\frac{1}{p}\norm{\mathbf{v}}^{2}\to 0$ almost surely as $p\to\infty$.
The derivation of the above is by considering a change of variable $\tilde{\bm{\beta}}=\bm{\Sigma}^{1/2}\hat{\bm{\beta}}$ whereby $\tilde{\bm{\beta}}\in\underset{\mathbf{b}\in\mathbb{R}^{p}}{\arg\min}\frac{1}{2}\|\mathbf{y}-\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}\mathbf{b}\|^{2}+h\left(\bm{\Sigma}^{-\frac{1}{2}}\mathbf{b}\right)$ and using the iterates of the VAMP algorithm (for non-separable penalties [37, Algorithm 1]) to track $\tilde{\bm{\beta}}$. One can then obtain (116) and Conjecture D.1 from the state evolution of the VAMP algorithm [37, Eq. (19), Theorem 1]. If Conjecture D.1 holds, it will be straightforward to develop inference procedure for $\bm{\beta}^{\star}$. A main gap to prove Conjecture D.1 in our opinion is to establish an analogue of Proposition B.8, i.e. the non-separable VAMP iterates indeed tracks $\hat{\bm{\beta}}$. We leave the proof of Conjecture D.1 as an open problem.
Remark D.2.
When $\bm{\Sigma}$ is unknown, we require access to a large unlabeled dataset $\mathbf{X}_{0}\in\mathbb{R}^{n_{0}\times p}$, with $n_{0}\gg p$, similar to the anisotropic Gaussian case. Let $\mathbf{X}_{0}=\mathbf{Z}_{0}\bm{\Sigma}^{1/2}$ where $\mathbf{Z}_{0}=\mathbf{Q}_{0}^{\top}\mathbf{D}_{0}\mathbf{O}_{0},\mathbf{Q}_{0}\in\mathbb{R}^{n\times n},\mathbf{O}_{0}\in\mathbb{R}^{p\times p},\mathbf{D}_{0}\in\mathbb{R}^{n\times p}$. We require the modeling assumption that $\mathbf{D}_{0}^{\top}\mathbf{D}_{0}\approx\mathbf{I}_{p}$. This assumes without loss of generality that as $n\to\infty$ and $p$ is fixed, the spectrum of $\mathbf{Z}_{0}^{\top}\mathbf{Z}_{0}$ converges to a point mass.
Under this assumption, we can estimate $\bm{\Sigma}$ using the standard sample covariance estimator $\hat{\bm{\Sigma}}=\mathbf{X}_{0}^{\top}\mathbf{X}_{0}$. The ellipsoidal-invariant assumption then requires that $\mathbf{X}\hat{\bm{\Sigma}}^{-1/2}$ can be modeled as a right-rotationally invariant random matrix. Unlike the anisotropic Gaussian assumption, we do not require the spectrum of the sample covariance matrix of $\mathbf{X}\hat{\bm{\Sigma}}^{-1/2}$ to converge to the Marchenko-Pastur law and is expected to lead to more robust debiasing performance.
Appendix E Numerical experiments
E.1 Details of the design matrices
Throughout the paper, we have illustrated our findings using different design matrices. We provide additional details in this section.
Remark E.1 (Notations used in caption).
we use $\mathsf{InverseWishart}\quantity(\bm{\Psi},\nu)$ to denote inverse-Wishart distribution [98] with scale matrix $\bm{\Psi}$ and degrees-of-freedom $\nu$, $\mathsf{Mult}$-$\mathsf{t}(\nu,\bm{\Psi})$ to denote multivariate-t distribution [99] with location $\bm{0}$, scale matrix $\bm{\Psi}$, and degrees-of-freedom $\nu$.
Remark E.2 (Right-rotationally invariant).
All design matrices in Figure 1, 3 satisfies that $\mathbf{X}\stackrel{{\scriptstyle L}}{{=}}\mathbf{X}\mathbf{O}$ for $\mathbf{O}\sim\operatorname{Haar}(\mathbb{O}(p))$ independent of $\mathbf{X}$. It is easy to verify that this is equivalent to right-rotational invariance as defined in Definition 2.1.
Remark E.3 (Comparison between designs in Figure 1 and Figure 3).
The designs featured in Figure 3 can be seen as more challenging variants of the designs in Figure 1, characterized by heightened levels of correlation, heterogeneity, or both.
Specifically, $\mathbf{\Sigma}^{\mathrm{(col)}}$ under $\mathsf{MatrixNormal}$-$\mathsf{B}$ has a higher correlation coefficient (0.9) compared to the correlation coefficient (0.5) in $\mathsf{MatrixNormal}$. This results in a stronger dependence among the rows of the matrix $\mathbf{X}$. Concurrently, the $\mathbf{\Sigma}^{\mathrm{(row)}}$ in $\mathsf{MatrixNormal}$-$\mathsf{B}$ is sampled from an inverse-Wishart distribution with fewer degrees of freedom, leading to a more significant deviation from the identity matrix compared to the MatrixNormal design presented in Figure 1.
In $\mathsf{Spiked}$-$\mathsf{B}$, there are three significantly larger spikes when compared to $\mathsf{Spiked}$ in Figure 1, which contains 50 spikes of smaller magnitudes. Consequently, issues related to alignment and outlier eigenvalues are much more pronounced in the case of $\mathsf{Spiked}$-$\mathsf{B}$.
Design under $\mathsf{LLN}$-$\mathsf{B}$ is product of four independent isotropic Gaussian matrices whereas $\mathsf{LLN}$-$\mathsf{B}$ contains 20th power of the same $\mathbf{X}_{1}$. The latter scenario presents greater challenge for DF or Spectrum-Aware Debiasing, primarily because the exponentiation step leads to the emergence of eigenvalue outliers.
Larger auto-regressive coefficients are used in $\mathsf{VAR}$-$\mathsf{B}$, leading to stronger dependence across rows.
When designs are sampled from $\mathsf{MultiCauchy}$, it is equivalent to scaling each row of an isotropic Gaussian matrix by a Cauchy-distributed scalar. This results in substantial heterogeneity across rows, with some rows exhibiting significantly larger magnitudes compared to others.
Definition E.4 (Simulated Designs Specification).
Below we give more detailed information on simulated designs from Figure 3, top-left experiment. Without loss of generality, all designs below are re-scaled so that average of the eigenvalues of $\mathbf{X}^{\top}\mathbf{X}$ is 1.
  • (i)
    $\mathsf{MatrixNormal}$-$\mathsf{B}$: $\mathbf{X}\sim N(\rm{0},\mathbf{\Sigma}^{\mathrm{(col)}}\otimes\mathbf{\Sigma}^{\mathrm{(row)}})$ where $\mathbf{\Sigma}^{\mathrm{(col)}}_{ij}=0.9^{|i-j|},\forall i,j\in[n]$ and $\bm{\Sigma}^{\mathrm{(row)}}\sim\mathsf{InverseWishart}\quantity(\mathbf{I}_{p},1.002\cdot p)$ (see Remark E.1 for notation);
  • (ii)
    $\mathsf{Spiked}$-$\mathsf{B}$: $\mathbf{X}=\mathbf{V}\mathbf{R}\mathbf{W}^{\top}+n^{-1}N(\rm{0},\mathbf{I}_{n}\otimes\mathbf{I}_{p})$ where $\mathbf{V},\mathbf{W}$ are drawn randomly from Haar matrices of dimensions $n,p$ respectively with 3 columns retained, and $\mathbf{R}=\operatorname{diag}(500,250,50)$;
  • (iii)
    $\mathsf{LNN}$-$\mathsf{B}$: $\mathbf{X}=\mathbf{X}_{1}^{15}\cdot\mathbf{X}_{2}$ where $\mathbf{X}_{1}\in\mathbb{R}^{n\times n},\mathbf{X}_{2}\in\mathbb{R}^{n\times p}$ have iid entries from $N(0,1)$;
  • (iv)
    $\mathsf{VAR}$-$\mathsf{B}$: $\mathbf{X}_{i,\bullet}=\sum_{k=1}^{\tau\vee i}\alpha_{k}\mathbf{X}_{i-k,\bullet}+\bm{\varepsilon}_{i}$ where $\mathbf{X}_{i,\bullet}$ denotes the $i$-th row of $\mathbf{X}$. Here, $\bm{\varepsilon}_{i}\sim N(\mathbf{0},\mathbf{\Sigma})$ with $\mathbf{\Sigma}\sim\mathsf{InverseWishart}(\mathbf{I}_{p},1.1\cdot p)$. We set $\tau=3,\mathbf{\alpha}=\quantity(0.7,0.14,0.07)$, $\mathbf{X}_{1}=0$;
  • (v)
    $\mathsf{MultiCauchy}$: rows of $\mathbf{X}$ are sampled iid from $\mathsf{Mult}$-$\mathsf{t}(1,\mathbf{I}_{p})$ (see Remark E.1 for notation).
Definition E.5 (Real-data Designs Specification).
Below we give more detailed information on real-data designs from Figure 3, bottom-left experiment. Without loss of generality, all designs below are re-scaled so that average of the eigenvalues of $\mathbf{X}^{\top}\mathbf{X}$ is 1.
  • (i)
    $\mathsf{Speech}$: $200\times 400$ with each row being i-vector (see e.g. [51]) of the speech segment of a English speaker. We imported this dataset from the OpenML repository [100] (ID: 40910) and retained only the last 200 rows of the original design matrix. The original dataset is published in [42].
  • (ii)
    $\mathsf{DNA}$: $100\times 180$ entries with each row being one-hot representation of primate splice-junction gene sequences (DNA). We imported this dataset from the OpenML repository [101] (ID: 40670) and retained only the last 100 rows of the original design matrix. The original dataset is published in [70].
  • (iii)
    $\mathsf{SP500}$: $300\times 496$ entries where each column representing a time series of daily stock returns (percentage change) for a company listed in the S&P 500 index. These time series span 300 trading days, ending on January 1, 2023.. We imported this dataset from Yahoo finance API [102];
  • (iv)
    $\mathsf{FaceImage}$: $1348\times 2914$ entries where each row corresponds to a JPEG image of a single face. We imported this dataset from the scikit-learn package, using the handle sklearn.datasets.fetch_lf2_people [103]. The original dataset is published in [48]
  • (v)
    $\mathsf{Crime}$: $50\times 99$ entries where each column corresponds to a socio-economic metric in the UCI communities and crime dataset [75]. Only the last 50 rows of the dataset is retained. We also discarded categorical features: state, county, community, community name, fold from the original dataset.
E.2 QQ plots
Figure 5, 6 and 7 are QQ-plots of Figure 1, 3 top and bottom row experiments respectively.
Figure 5: QQ plots corresponding to Figure 1.
Figure 6: QQ plots corresponding to Figure 3, top-left.
Figure 7: QQ plots corresponding to Figure 3, bottom-left.
E.3 Marginal inference under exchangeability
Figure 8 below shows an illustration for this result focusing on =$\mathcal{I}=\{1\}$. Observe that we once again outperform degrees-of-freedom debiasing.
Figure 8: Histograms of $\frac{\hat{\beta}_{1}-\beta_{1}^{\star}}{\sqrt{\hat{\tau}_{*}}}$ across from 1000 Monte-Carlo trials using DF and Spectrum-Aware Debiasing. The setting is identical to Figure 1 except that here we set $n=100,p=200$ for computational tractability.
E.4 Alignment tests for simulated designs
Figure 9 shows results of hypothesis tests for the alignment coefficients $\upsilon^{\star}_{i},i=1,...,6$ for experiments in Figure 3.
Figure 9: top-row: Same setting as Figure 3, bottom-left experiment. bottom-row: Same setting as Figure 3, bottom-left experiment, except that we increase difficulty of alignment detection by considering $v_{i}^{*}=\sqrt{p},i\in\{2,4,6\}$ as opposed to $v_{i}^{*}=5\cdot\sqrt{p},i\in\{2,4,6\}$. left-column: Benjamini-Hochberg adjusted p-values $\tilde{P}_{i}$ for $H_{i,0}^{\bm{\upsilon}^{\star}},i=1,...,6$. ** indicates rejection under FDR level 0.05 and * rejection under FDR level 0.1. Recall from Section 4.4.6 that rejection of $H_{i,0}^{\bm{\upsilon}^{\star}}$ indicates alignment between $\bm{\beta}^{\star}$ and $\mathbf{o}_{i}$. right-column: True alignment angles between the signal $\bm{\beta}^{\star}$ and $i$-th PC $\mathbf{o}_{i}$ calculated using $\mathrm{argcos}\{\mathbf{o}_{i}^{\top}\bm{\beta}^{\star}/(\norm{\mathbf{o}_{i}}_{2}.\norm{\bm{\beta}^{\star}}_{2})\}$.
E.5 Misspecified setting
Our theory assumes that all relevant covariates are observed. In practice, there may be unobserved features $\mathbf{Z}$ and responses is generated from $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\mathbf{Z}\bm{\theta}^{\star}+\bm{\epsilon}$. We found that our method is in fact relatively robust to such mis-specification. We demonstrate by introducing unobserved features $\mathbf{Z}$ of shape $n\times\frac{p}{2}$ (i.e. half as many as the observed features $\mathbf{X}\in\mathbb{R}^{n\times p}$) drawn from different distributions.
Figure 10: Z from MatrixNormal Same settings and legends as Figure 3, bottom-left experiment, except that the response is generated from a misspecified model $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\mathbf{Z}\bm{\theta}^{\star}+\bm{\epsilon}$. We set ${\theta}^{\star}_{i}\sim 0.2\cdot N(-5,1)+0.3\cdot N(2,1)+0.5\cdot\delta_{0}$. Here, we let $\mathbf{Z}$ be sampled from MatrixNormal as described in Figure 1 with dimension half as many as $\mathbf{X}$.
Figure 11: Z from Spike. Same settings and legends as Figure 3, bottom-left experiment, except that the response is generated from a misspecified model $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\mathbf{Z}\bm{\theta}^{\star}+\bm{\epsilon}$. We set ${\theta}^{\star}_{i}\sim 0.2\cdot N(-5,1)+0.3\cdot N(2,1)+0.5\cdot\delta_{0}$. Here, we let $\mathbf{Z}$ be sampled from Spike as described in Figure 1 with dimension half as many as $\mathbf{X}$.
Figure 12: Z from LNN. Same settings and legends as Figure 3, bottom-left experiment, except that the response is generated from a misspecified model $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\mathbf{Z}\bm{\theta}^{\star}+\bm{\epsilon}$. We set ${\theta}^{\star}_{i}\sim 0.2\cdot N(-5,1)+0.3\cdot N(2,1)+0.5\cdot\delta_{0}$. Here, we let $\mathbf{Z}$ be sampled from LNN as described in Figure 1 with dimension half as many as $\mathbf{X}$.
Figure 13: Z from VAR. Same settings and legends as Figure 3, bottom-left experiment, except that the response is generated from a misspecified model $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\mathbf{Z}\bm{\theta}^{\star}+\bm{\epsilon}$. We set ${\theta}^{\star}_{i}\sim 0.2\cdot N(-5,1)+0.3\cdot N(2,1)+0.5\cdot\delta_{0}$. Here, we let $\mathbf{Z}$ be sampled from VAR as described in Figure 1 with dimension half as many as $\mathbf{X}$.
Figure 14: Z from Mult-t. Same settings and legends as Figure 3, bottom-left experiment, except that the response is generated from a misspecified model $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\mathbf{Z}\bm{\theta}^{\star}+\bm{\epsilon}$. We set ${\theta}^{\star}_{i}\sim 0.2\cdot N(-5,1)+0.3\cdot N(2,1)+0.5\cdot\delta_{0}$. Here, we let $\mathbf{Z}$ be sampled from Mult-t as described in Figure 1 with dimension half as many as $\mathbf{X}$.